A Complete Bibliography of Publications in Bayesian Analysis

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 August 2024 Version 1.24

Title word cross-reference

3 [BVN09, SLB^{+}21]. _{2} [RCLW17]. A [AGG16]. α [GMY21]. D [AGG16]. F [MP18]. G [ZHG^{+}16, BH11, FN22, HBJ14, SKG15, Wan17]. Γ [GD09]. J [HYDE21]. L^{p} [Scr14]. M [LC17]. M [CCY13]. p [FMM18, SF14]. P (X < Y) [RS13, VR11]. ψ [SM19]. R^{∗} [LV22]. t [CF10, FD14b, HSH21a].

-complete [CCY13, LC17]. -D [BVN09]. -Distributions [FD14b]. -Divergences [GMY21]. -metrics [Scr14]. -minimax [GD09]. -Open [LC17]. -Optimal [AGG16]. -Prior [ZHG^{+}16]. -Priors [FN22, HBJ14, SKG15, Wan17]. -Resolved [HYDE21]. -Stick [SM19]. -value [SF14]. -Values [FMM18]. -walk [CF10].

19 [MBB^{+}23].

Aalen [DRRS17]. ABC [GRM^{+}09, Pra17]. Abnormal [BF17]. Abundance [GSWF19]. Accelerated [GW16]. Accelerating [DEGP22, WSD22]. Accelerator [LSZH06]. Accounts [BG13]. Accuracy [CS12]. Accurate [Gop22]. Acquisition [JGP^{+}19, SMBS23]. Activation [GS21]. Activity [ALC24, HAJF23]. Acyclic [CCVP18, DBHG19]. Adaptation [NdVA^{+}20]. Adapting [Pra17]. Adaptive [BS14, BW15, BCJ21, FM18, FT13, GM16, LLW21, LBBJ16, Ma17, MTS^{+}21, PKLM10, SCHT13b, Scr14, SK17, SCKGC21, SOL^{+}12, SKN24, XX20, YSLR14, RW08]. Additive [KK16, KCK^{+}21, VHV20, ZSM07]. addresses [AAFS06]. adjusted [ZSM07]. Adjustment [APD19, SNMS23]. advances [VR11]. Adversarial [PHG23]. Affecting [OBS13]. After [XTMR17]. Against [GDB20, Gag23, GBGTR19]. Age [MBB^{+}23, BC11a]. age-depth [BC11a]. Age-specific [MBB^{+}23]. Ahead [POWK24]. Air [DWM^{+}21, XTMR17]. al [LH10, Ver06, WFR11b]. al. [AB09, BD09, Car06, Che06, CS07, Dun09, Fea11, Fre12, GM13b, Gli09, Gos12, Hen10, HG08, Hoe06, Koo11, LG06, MV06, Plu06, Poo10, QM09, Ran10, Rig10, Rou08, RC07, Sca12, Sch09, SS10, Sta12, Whi10, Woo13, vD10, vdL06]. Albert [Fre12, Gos12]. Algorithm [SKN24, WOJL22, ZG19, CF10, WT06]. Algorithmic [HSF20]. Algorithms [NdVA^{+}20, PMG14]. Allergy [GHO^{+}13]. Alleviating [OMC19]. Allocation [Mad07]. Allocations [BPJ13]. Almost [AZ13]. alpha [PKL^{+}11]. alpha-stable [PKL^{+}11]. Alternative [OM22]. Analysers [MVG20]. Analyses [GCM24, VG23, WG15, BVN09, CZ10, Chr06, CLM07]. Analysis [APS18, ADL12, BHvD17, BG21, BJM^{+}22, Ber06a, BB24, BCLM24, Bra22, CMG14, CFLN18, CCL^{+}09a, DP12, FSG08, GTHB19, Gol06a, GGPM19, HSH21a, HHG08, KSM^{+}06, KSM^{+}18, KFF19, KEMM19, LBB09, LBBJ16, MC07, NJM18, OS24, QPC24, Raj19, RCLW17, RdGvP06, RMHR15, SXR06, SSML20, SLB^{+}21, SHMM23, SCFJ14, TRWFB17, VGB10a, VG23, WG18, Wan17, YHW16, ZJLC10, ZL24, ZWC^{+}16, ZWF^{+}18, Zho18, dTM10, AZ10, AVCGG08, BM06, Dra06, FMV11, FS11, HKLM10a, JKNR09, Kad06, MPK10, OBS13, RH11, vdL11b]. Analyzing [CG10]. Ancestral [XS07]. angle [HKLM10a]. ANOVA [CS16a, KS10a]. Application [ATF23, AFRB14, BSPD23, BGQ21, GS21, HdHG21, HGXS23, MNS^{+}20, NJ21, RSSSSL21, SS08, SW22, WHG^{+}06, XX20, ZWC^{+}16, AVCGG08, BVN09, FMV11, GP10, LN08, LZN08, Tre08]. Applications [BHJ18, BR13, FCP09, GDNJ18, RL14, ZWDJ14, Hof11b]. Applied [MZMK24, RDP16, Bar11]. Approach [Bha07, BGQ20, Cas21, CCVP18, CGS22, CAD^{+}23, DK15, FH17, GDB20, GMB20, GHO^{+}13, GGPM19, HMC09, HSH^{+}21b, HMZ^{+}22, HSF20, LM16, LM21, LC23, MBBRB17, NBCC14, OJP23, PHG23, RMP12, TK12b, VDP15, WPCAV22, HS09, JP08, MS07a, SB11]. Approaches [SC17]. Approximate [BW15, CNR15, DPM16, GMS16, GL17, HRT24, HSH^{+}21b, JGP^{+}19, LNR19, LC22, PKL^{+}11, RCMO22, SCKGC21, WFR11a]. Approximation [AZ13, BJS23, LR16, NDME18, RV14, SK13, RM08]. Approximations [ADP19, HAJF23, JB18, NS18, QNK23, RSV14]. Aquifer [SHG^{+}10].

Arbitrary [HSBvdW17]. Architectures [FMO16]. Arctic [ZC20]. Area [ADL12, Pol17, RSV14, SW22]. Areal [MC07, OMC19]. arguments [TGM09]. Arithmetic [Paj17]. Armed [CBC23]. arrays [Hof11b]. Arsenic [CCL^{+}09a]. Article [APA^{+}13, Ano14a, Ber14, BCT^{+}16, Bur10, CM13, CB14, Cas14, CD15, CLH^{+}16, Cla12, CC15, Das16, DL15, Dob13, Fea11, Fer12, For14, Fre12, GPP16, Gel10, GM13b, Gos12, GL16, Gra16, GMR15, GB12, HP15, Han16, Hof13, KB15, Koo11, Lam06, LH10, Lia12, LC12, Lys16, MYGE16, MGP15, O’H13, PS13, RF16, Rou15, Sca12, Sco14, Sha14a, Siv15, Sta12, Wan13, Was10, WS14, Woo13, WFR11b, Xu14, Zid15, tHM14, AB09, All11, BD09, Ber08, Car06, Che06, CK09, CGM09, CS07, Cra09, Dah07, Dun09, Fra09, Fre11, FS08, Gel06, Gli09, Han11, HP08, Hen10, HG08, Hoe06, Hof11a, Kad08, KN06, Li09, LG06, Mac07, MCG11, MV06, Mil08, Plu06, Poo10, QM09, Ran10, Rig10, Rob07, Rou08, RC07, Sch09, Sen08, SYvD11, SS10]. article [SK08, Ste09, Ver06, Was08, Whi10, vD10, vdL06]. articles [Chr06, Dra06, Fie06a, Kad06, Kas06, Lad06, O’H06, Was06]. Artificial [Per07]. Aspects [Joh13, NB18]. Assess [CHG12]. Assessment [BE13, GHO^{+}13, Joh07, LG17, MS07b, WG15, Rob10, Tre08]. Assessments [PVC20]. Assisted [DM07a]. associate [MT09b]. Associated [Kad16]. Association [CS12]. Associations [LMC20]. Astrophysics [vDCE^{+}06]. Asymmetric [LG12b, RS13, SSML20, SRG13, SR17]. Asymptotic [AZ13, DG13, GTGC16, GC17, Kom15, Spi08]. Asymptotics [GM13a]. Atlantic [TGK^{+}11]. Atrophy [RGC20]. Attraction [WDML22]. Augmentation [TAN^{+}18, WPS23, PS11a, PS11b]. Auto [DBHG19]. Auto-Regressive [DBHG19]. Automated [TdVPAB17]. Autopsies [LMC20]. Autoregression [DGMQ13, HK22, PKL^{+}11, YHW16]. Autoregressive [CVL12, GLC24, KFF19, KCR19, KG09, LBBJ16, Per07, SCFJ14, BC11a]. Auxiliary [OM20, HH06, vdL11a]. Available [SN07]. Average [YVSG18]. averages [MM07]. Averaging [SXR06, YMP13]. avoiding [LZN08]. Away [RRJW20]. axioms [DT09].

B [MBB^{+}23]. B-splines [MBB^{+}23]. Bagged [HM23]. balancing [GP10]. Balls [WG18]. Banded [LL20, LLL23]. Bandwidth [LL20]. BART [CGMS22]. Baseball [QMRM08, JMW09a]. Based [ANRSL16, BS14, BB24, CBC23, DM15a, DL07, HRT24, JGP^{+}19, LLPR06, LTY21, Nee19, NTL19, PQ15, Per07, RMP12, SCHT13b, SN07, SRG13, SR17, SNMS23, US16, VL20, XLH16, XTMR17, BD06a, BAR23, FI09, GP12, Hof06, HHG08, LAE^{+}09, MS07a, PFS10, RW08, Vir11]. Baseline [Han06]. basic [CO08]. Basket [LTY21]. Bayes [ATF23, Ald08, AKO19, BE13, BVN09, CCDT^{+}22, CCVP18, CS16a, DvdBWL24, DG13, EH17, GTGC16, GHO^{+}13, HC17, HdHG21, LC17, LZN08, MF19, TGM09, WOPF11, Was06, Wei12, Woo14]. Bayesian [Fie06a, Fie06b, Kad06, SR17, vdL11a, APS18, AGG16, ADP22, AM07, AZ10, AO06, AVCGG08, ADL12, APRS22, AFRB14, BPSS15, BM06, Ban17, Bar11, BF17, BB10, BP20, BSPD23, BHvD17, BG06, BG21, BF21, BJM^{+}22, Ber06a, BJS23, BGP15, BHJ18, Bha07, BLE16, BB24, BW15, BC11b, BR10, Bra22, BD06a, BG13, BALO06, BS21, BMBV22, CNR15, CKY20, CHG12, CS13, CZ10, CL24a, CL24b, CCDT^{+}22, CS12, CVCB23, CVL12, CLMtH15, CZGV19, CC21, CEMR12, CB21, CBC23, CHIK08, CFH23, CDH16, CCCG16a, sC16, Chr06, Chr09, Chr24, CO08, COIG19, CFLN18, CGS22, CCL^{+}09a, CAD^{+}23, CT11, CAV23, CHMK22, DCKW08, DM15a, DWM^{+}21, DW13, DRH17, DG11, Des13, DLPS20, DGMQ13, DHDC12, DR16, Dra06, DPM16, DT18, DD07, DT09, DD18, EMS13]. Bayesian [FT12, Fie06b, FH17, FD14b, FMV11, FCP09, GDB20, Gel08a, GLM18, GMP21, GLJB23, GTHB19, Gol06a, Gol06b, GD09, GMB20, GMdPV21, GMS16, GL17, GKMvCT14, GABP19, GW16, GC18, GvO17, GRM22, GR23, GS21, GGPM19, GBGTR19, HAJF23, HMC20, HRT24, HOHS24, HJZ12, HSH21a, HYDE21, HK22, HSBvdW17, HKLM10a, HMC09, HH06, HCGS15, HCH06, HSH^{+}21b, HGXS23, HMZ^{+}22, HD12, HSF20, HYY12, Hut07, IW19, JGP^{+}19, JGVM21, JMW09a, JP16, JKNR09, JD08, JYL17, JL19, Joh07, Joh13, JHB22, Kad06, KJP24, KR21, KS10a, KFF19, KD12, KK22, KDV09, KAL12, KSLP12a, KCR19, KEMM19, KS19, KCK^{+}21, KDG21, Kob17, Kom15, KC23, KMB19, KG09, KGGC10, Kyu11, Lad06, LHE^{+}20, LMLM14, LJCB14, LL18, LNR19, LL20, LL23, LYL24, LG17, LM16, LM21, LKOB19, LC22, LML21, LN08, LL10]. Bayesian [LXL10, LG14, LMC20, LBLS22, LMPS17, LW09, LBB09, LN13, LCL^{+}14, LG24, LC23, MJW08, MC07, ML22, MKIM24, MG23, MMN22, MS07a, MBBRB17, Mar24, MZMK24, MMW15, MNS^{+}20, MS07b, MMJ16, MC15, MW15, MNPM20, MRG19, MG20, MM13a, MHSC16, MQ22, Nee19, NBCC14, NJ21, NGT19, NDME18, NTL19, OS09, OJP23, OBS13, OS24, OGPD19, OM20, OM22, POWK24, PK24, PW19, PMS24, dBPSW08, Per07, PKLM10, PKL^{+}11, Poi06, Pol17, PS17, PPG08, PBT^{+}21, PJM^{+}21, PHG23, Pra16a, PW08, Qia18, QMRM08, Rah16, RCLW17, RCMO22, RdGvP06, RL14, RB07, RtH08, RD11, RH11, RMHR15, RC17, RGC20, RS13, RSST17, RDP16, SRA23, San12b, SMBS23, SW22, Sco11, Scr14, SXR06, SK17, Sha21, SY17, SY19, SCKL22, SS11, SSML20, SPG15, SCKGC21, Ski06, SHMM23, SCFJ14]. Bayesian [Spi08, Spi11, SRG13, SKN24, SB11, SG16, SG17, TM17, TRWFB17, TFHP18, TZG10, TK12b, Tre08, TSA20, US16, VR11, VDP15, VGB10a, VG23, VDP19, WMP11, WG18, WT06, Wan12, WB18, WT20, WCO20, WOJL22, WPS23, WSD22, WWACH16, Wen10, WC18, WGBS17, WS20, WG15, WN21, WM23, WFR11a, XLH16, XX20, XCPX22, XLY^{+}13, YS07, YHW16, YZCC16, YN20, YVSG18, YPVG22, YH11, Yin09a, YMP13, YMX23, YSLR14, ZM23, ZLV24, ZSM07, ZJLC10, ZL15, ZC20, ZL24, ZWC^{+}16, ZWF^{+}18, Zho18, ZG19, ZD17, dCJHdC13, dCPB19, dTM10, pD20, vES21, vdBDB23, vdL11a, vdL11b, vdPvdV18]. Bayesians [Kas06]. be [Fie06a, dBPSW08]. become [Fie06b]. Behavior [EMS13]. Behind [CCL^{+}09a]. Behind-the-Scenes [CCL^{+}09a]. Belief [AE17, BE13, WG15, Hoo08]. Beliefs [TGK^{+}11]. Berger [Chr06, DL15, Dra06, Fie06a, Kad06, Kas06, Lad06, MGP15, O’H06, Rou15, Siv15, Was06]. Bernardo [DL15, MGP15, Rou15, Siv15]. Bernoulli [Kad16]. Bernstein [PS15]. Beta [BJP12, CVL12, CLMtH15, TM17]. Beta-Binomial-Logit [TM17]. Beta2 [PPR17]. Between [CI06, FH17, SF14]. Beyond [KEMM19]. Bi [XLY^{+}13]. Bi-Clustering [XLY^{+}13]. Bias [dOAL^{+}22, LZN08]. Biclustering [MQ22]. Big [Qia18]. Bilateral [MC15]. Binary [AFRB14, Chr24, DK15, HH06, HvDH09, RH11, vdL11a]. Binomial [BJS23, Gop22, Kad16, MJW08, Nee19, TM17, ZWF^{+}18, Zho18, TGM09]. Biological [MMN22, RDP16]. Bipartite [GRM22]. birth [DZP^{+}07a]. Bivariate [PMS24, Leo11]. Blackwell [HP08, Mil08]. Block [APL24, BDW^{+}24]. Block-Correlated [BDW^{+}24]. Blocking [TdVPAB17]. Blockmodels [HLC20]. Board [Ano16a, Ano16b, Ano23a, Ano23b, Ano24a, Ano24b]. Bootstrap [VDP19]. Bootstraps [BP20]. Both [Pol17]. Boundaries [JV23]. Boundary [BHvD17, MC07, RSST17]. Bounded [MDO18]. Bounds [MM16]. Brain [CL24b, DD18, GS21, RGC20, SLB^{+}21]. Branching [GMdPV21]. Breaking [BJP12, FLN^{+}16, GLJB23, HZ22, SM19, RD11]. Breast [DD07]. Bronchial [HCH06]. Browne [Gel06, KN06, Lam06]. Buck [HP08, Mil08]. Buffet [CGZ16, HR20, WDML22]. Building [CCL^{+}09a]. buy [Lad06].

calculating [WT06]. Calculation [ZS09]. Calculations [PHG23]. Calderhead [BCT^{+}16, Das16, Lys16, MYGE16]. Calibrating [PVC20]. Calibration [CLMtH15, Gu19, LNR19, MF22, RMP12, BB08a, BALO06, Dra06]. Calibration-Based [RMP12]. Campbell [Das16, Lys16, MYGE16, BCT^{+}16]. Can [dBPSW08]. Cancer [DD07, MZMK24]. Card [BMBV22]. Carlo [AZ10, BM06, BW15, BCJ21, DT18, FT13, HS09, MKIM24, MBC24, ND20, PMG14, PKLM10, Ryd08a, SPD19, TDY18, TdVPAB17, WCKL18, Wei12, YSH18, ZSZ18, ZL24]. Carvalho [Cas14, For14, tHM14]. Case [Ber06a, FCP09]. Categorical [HRW18, PFS10, PW19, JD08]. Categorization [HdHG21]. Cauchy [GLM18, PS12]. Causal [CC21, GRM22, HMC20, NGT19, SNMS23, ZM23, FS11]. Causality [DMF16]. CDFs [SC06]. Celeux [Car06, Che06, MV06, Plu06, vdL06]. Cells [HCH06]. Censored [Cas21, Han06]. censoring [JD08]. Centered [PHOD21]. Central [HZ22, NJ21]. Centroids [OMC19]. certainty [Rob10]. Chain [AQ17, BHS14, CS16b, PMG14, SPD19, TDY18, TdVPAB17, Wei12, ZL24, HS09, PKLM10, Ryd08a]. Chains [MG23, SOL^{+}12, ZWC^{+}16]. Change [KCG15, MM14, PCM19, QPC24]. Change-Point [PCM19, KCG15]. Changepoint [HHT24, Sha21, WFR11a]. Changes [PS20, ZJLC10, ZC20]. Checking [AKO23, CCL^{+}09a, EM06, NSAL^{+}21]. Chi [NJ21]. Chi-Squared [NJ21]. Chief [Car08, Car09]. Chkrebtii [BCT^{+}16, Das16, Lys16, MYGE16]. Choice [BAR23, MMW15, MNS^{+}20, GRM^{+}09]. Claiming [EMS13]. Claims [CGS22]. Class [DGMQ13, GTGC16, LVW20, RSSSSL21, SR16, SN07, SM17, VGE19, Dah09]. Classes [ANRSL16, CCVP18]. Classification [GR23, LMCD19, SN07, LZN08]. Classifiers [LV22]. Classifying [MMN22]. Clean [DWM^{+}21]. Climate [SFZ08a, SOL^{+}12, VHJS08]. Clinical [FCP09, HSC12, HMZ^{+}22, SY17]. Cluster [CMG14, GM16, Mad07, WG18, RW08]. Clustered [GM16, YMP13, dOAL^{+}22]. Clustering [BGQ21, CBC23, DRH17, LAE^{+}09, NBCC14, PHOD21, PQ15, PFS10, SG16, XLY^{+}13, BC11b, CT11, Dah09, FI09, Hof06, Vir11, YH11]. Clusters [RVWG24, MY08, Ngu10]. Co [CH09, CT11]. co-exposure [CT11]. Co-infection [CH09]. Coarsened [GCM24]. Coefficient [SCFJ14]. Coefficients [PB20]. Coherence [Dra06]. Cointegrated [PKLM10, PKL^{+}11]. Colombian [WPCAV22]. Combination [LN13, AZ10]. Combine [RMP12]. Combined [HYY12]. Combining [ADGJ^{+}12a, BP08, MG23, WHG^{+}06]. Combustion [VDF^{+}12]. Commensurate [HSC12]. Comment [AB09, All11, BD09, Ber08, Ber14, Bur10, Car06, CM13, CB14, Cas14, CD15, CGM09, Cla12, CS07, CC15, Dah07, Das16, DL15, Dob13, Dun09, Fea11, Fer12, For14, Fra09, Fre11, Fre12, FS08, GPP16, Gel10, GM13b, Gli09, Gos12, GL16, Gra16, GMR15, GB12, Han11, HP15, Han16, HP08, Hen10, HG08, Hof11a, Hof13, Kad08, KB15, Koo11, Lam06, LH10, Li09, Lia12, LC12, LG06, Lys16, Mac07, MCG11, MYGE16, MGP15, MV06, Mil08, O’H13, PS13, Plu06, Poo10, QM09, Ran10, RF16, Rig10, Rob07, Rou08, Rou15, RC07, Sca12, Sch09, Sco14, Sen08, SYvD11, Sha14a, Siv15, SS10, SK08, Sta12, Ste09, Ver06, Wan13, Was08, Was10, WS14, Whi10, Woo13, WFR11b, Xu14, Zid15, tHM14, vD10, vdL06, Chr06, Dra06]. comment [Fie06a, Gel06, Hoe06, Kad06, KN06, Kas06, Lad06, O’H06, Was06, vdL11a]. Comments [Che06, CK09, Cra09, Dra06]. Communities [LC23]. Community [APL24, SC17, vdPvdV18]. Comparative [SXR06]. Compare [MRB12]. Comparing [BP07, CEMR12, GBGTR19, ZHA23]. Comparison [CS13, CB21, HK18, TAN^{+}18, WM23, XTMR17, BD06a]. comparisons [Spi11]. Complete [LC17, CCY13]. Completely [CAS^{+}19, AM07]. Completion [YMX23]. Complex [Bha07, WG15, ZG19]. Component [ZHG^{+}16]. Components [JN07b, MB12, SG17, KN06]. Compound [ZL15]. Compressing [LN08]. Computation [BW15, BAR23, CNR15, GL18, JGP^{+}19, PKL^{+}11, RCMO22, SSLD23, SCKGC21, Ski06, Wan12, CHIK08]. Computational [Kyu11, VHJS08, WWACH16, Ryd08a]. Computationally [BHW18]. Computations [WSD22]. Computer [Bha07, JV23, LBB09, MF22, SFZ08a, VG23, WHG^{+}06, LW09]. Computing [Wei12]. Concentration [DRRS17, RRJW20, RSM15, RR12, Ros22, SCKL22]. Concentrations [TFHP18]. Concept [HHG08]. Conclusions [TGK^{+}11]. Conditional [GLC24, JS24, KR21, KFF19, WRC11, CCQ11]. Conditionally [KG09, LBBJ16, MTM12]. Conditioning [LML21]. Conditions [RSST17, SKG15]. Configuration [KD12]. Conflict [AE17, EM06, NSAL^{+}21]. Conflicting [Gag23]. Confounding [APD19, HCPH18, HMC20, OMC19]. Conjugate [COIG19, DP12, KSM^{+}06, KSM^{+}18, PSMB20, CHIK08, KN06, Pac06, WMP11]. Connectivity [SHG^{+}10]. Consensus [CAV23, TGM09]. Consider [FJM14]. considering [HHC07, PW08]. Consistency [CKG20, ML22, OK22, SKG15, SRG13, SR17]. Consistent [CL24b, YN20]. Constant [Hut07]. Constrained [ADMO24, CS13, GM13a, LKOB19]. Constraint [SRA23]. Contaminated [CCO24]. Contents [Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13b, Ano13c, Ano13d, Ano13e, Ano14d, Ano14e, Ano16c, Ano16d, Ano23c, Ano23d, Ano24c, Ano24d]. Context [EMS13, NPKC14, PK24, US16]. Context-Dependent [US16]. Context-Specific [NPKC14]. contingent [LKF09]. Continuous [APD19, HMZ^{+}22, HYY12, QSF09, SS08, Sha21, XCPX22, ZWC^{+}16, CF10, HS09]. Continuous-Time [SS08]. Contraction [CGZ16]. Contributed [APA^{+}13, Ano14a, BCT^{+}16, CLH^{+}16]. Control [CCDT^{+}22, GBGTR19, MTS^{+}21, SY19, SOMD23]. Controlled [GMdPV21]. Controlling [CBR23]. Convergence [LV22, NS23, PK24, WT06]. Convolution [GSWF19, ZKRVA18, MPK10]. Conway [KSM^{+}18, BF21, KSM^{+}06, Kad16]. Cooling [MTS^{+}21]. Copula [GL17, SCHT13b]. Copulas [GC18, KS19, Wil18]. Core [FMO16]. Correction [CB21, KEMM19, SR17, dOAL^{+}22, KSM^{+}18]. Correlated [BBGR21, BDW^{+}24, GL22, NGT19, PSC23, PL16, MAL11]. Correlation [CRAL23, GMP21, LHE^{+}20, PBT^{+}21]. Correlations [MF19]. Correspondence [dTM10]. Count [APS18, BHW18, HIS22, XLY^{+}13]. Count-Valued [BHW18]. Counting [Chr24, DRRS17]. County [SW22]. Coupled [OM20]. Covariance [FJM14, HW13, KK22, LHE^{+}20, LL18, LM16, LM21, MP18, XCPX22, YZCC16, Hof11b]. Covariances [LLL23]. Covariate [HD12, MHSC16]. Covariates [BWD20, DCKW08, DLPS20, PQ15, WT20, MT09b]. COVID [MBB^{+}23]. COVID-19 [MBB^{+}23]. Craigmile [BD09, Dun09, Sch09]. credibility [GD09]. Credible [LNR19, Sha14b, WG18, ZB18, DM07b]. Credit [BMBV22]. Criminal [BS21]. Criteria [CFRT06a, US16, FI09]. critical [Rob10]. Criticism [SMW19]. Cross [BH07, HC17]. Cross-Validation [BH07, HC17]. Cumulative [MMW15]. Current [PMS24]. curve [BALO06]. Curves [CDH16, BB08a]. Cyclist [DWM^{+}21]. Cyclone [TGK^{+}11]. Cytometry [GCM24].

D [BVN09, SLB^{+}21]. DAGAR [DBHG19]. Dark [LC22]. Data [APS18, AE17, ADP22, AQ17, AFRB14, BP20, BHW18, BG13, BR13, Cas21, CFRT06a, CB21, Chr24, DCKW08, DRH17, EH17, EM06, GMP21, GL22, GR20, HIS22, Han06, HRW18, HGXS23, KJP24, KK22, Kom15, Kow21, LJCB14, LM16, LM21, LMPS17, LBBJ16, MCW10b, MC15, MTS^{+}21, NSAL^{+}21, OMC19, PS20, PMS24, PS11a, PS11b, PBT^{+}21, Qia18, QMRM08, SW22, SSML20, SG16, SG17, TM17, TAN^{+}18, VHV20, WPS23, WHG^{+}06, XLY^{+}13, YZCC16, ZKRVA18, ZWF^{+}18, ZD17, dOAL^{+}22, dTM10, DGS09, GM09, GP10, Hof11b, HvDH09, JD08, Ngu10, RH11, Spi11, Vir11, vdL11b]. Data-Dependent [TM17]. Datasets [APRS22, ZSM07, BM06, HKLM10a]. Dawid [GMR15, HP15, KB15]. day [PKL^{+}11]. De-Duplication [TSL20]. Deaths [MBB^{+}23]. Decision [DWM^{+}21, IW19, LV22, OM22, XTMR17]. Decision-Theoretic [OM22, XTMR17]. Decisions [HMZ^{+}22, KM14]. Decomposable [FJM14, BC11b]. Decomposition [Hof16]. Decompositions [ZR21]. Deconvolution [HYDE21, vDCE^{+}06]. Decoupling [BCLM24]. Deep [MF22, PS17, WPS23]. Default [DvdBWL24, Gri10, KN06]. Definite [WC14b]. Degrees [VW14]. Delayed [LTY21, LN13, MBC24]. Demographic [BG13]. Densities [CLMtH15, GMY21, Kom15]. Density [BGQ20, GL18, HK22, JLM^{+}17, POWK24, RV14, Scr14, SRG13, SR17, TZG10, WPCAV22, Gri10, RtH08]. Dependence [CB21, FH17, LL23, LM16, WS20, WFR11a]. Dependencies [WRC11]. dependency [PW08]. Dependent [BJQ12, DD07, JLM^{+}17, KCR19, KK16, MHSC16, RS13, SW22, TM17, US16, ZHA23, ALR21]. depth [BC11a]. Deriving [JS24]. Desiderata [Cla10]. Design [AFRB14, DT18, KDG21, LTY21, LN13, Mad07, OM20, OM22, PHG23, RDP16, SY17, WWACH16, ZM23, dG15]. Designing [TDY18]. Designs [AGG16, SY19]. Detecting [PS20, YS07]. Detection [APL24, BF17, BMBV22, HHT24, MM14, SC17, Sha21, SS11, TGK^{+}11, vdPvdV18]. Determinacy [SHMM23]. Determinantal [BGQ20]. Determination [MJW08, YH11]. Deviance [CFRT06a]. Diaconis [JB18]. Diagnosing [TN14]. Diagnosis [CL24b]. Diagnostic [LV22, YS07]. Diagnostics [Per07]. Diagrams [Mar24]. DIC [MRB12]. did [Fie06b]. diet [CT11]. Differences [DD18]. Different [Kom15]. Differential [BKD21, CCCG16a, HCH06, MMJ16, PMG14, WCO20, YSLR14, DGS09]. Differing [JV23]. difficulties [RM08]. Diffusion [WGBS17, SB11]. Diffusion-Driven [WGBS17]. Dimension [HSBvdW17, JV23, TRKS^{+}17]. Dimensional [APD19, Ban17, BHW18, CKG20, GR23, LYL24, LAE^{+}09, MRG19, OK22, RR12, RGC20, SN18, SKG15, SKN24, YN20, GC17, Joh13, LL20, LL23, MT09b, QNK23, Spi08]. Dimensionality [OK22]. Dimensions [AGG16, ML22]. Direct [DL07, AZ10]. Directed [BG06, CCVP18, DBHG19]. Directional [FJS08, KG09]. Dirichlet [ALR21, AJGM22, BJQ12, BGQ21, BJ06, EDF^{+}19, FD14b, GM13a, HRW18, JN07b, KDV09, KCG15, MCMK20, Raj19, SMBL19, SS11, TK09, TRKS^{+}17, XS07, ZWDJ14, ZB18]. Disconnected [BG21]. Discovery [MZMK24]. Discrepancy [DW13, FMM18, OGPD19]. Discrete [HYY12, PWB12, PNNC17, WT20, FS11]. Discretely [SS08]. Discussion [APA^{+}13, Ano14a, BHJ18, BHW18, BCT^{+}16, CDL^{+}19, CLH^{+}16, COIG19, GLJB23, HMC20, HSF20, KK22, LVW20, LML21, MF22, OGPD19, PHOD21, SNMS23, TSL20, WG18, YVSG18, ZR21, vdPSvdV17, Gol06b]. Disease [CL24b, DBHG19, MC07, MBBRB17, MNS^{+}20, VDF^{+}12]. diseases [JKNR09]. Disparities [GMdPV21]. Displacing [OMC19]. Dissonance [SCKL22]. Distance [Pra17, Sal18, She14]. Distortion [ANRSL16]. Distributed [BAR23, HOHS24, Ngu10]. Distribution [BF21, DG11, GGPM19, HSBvdW17, HGXS23, KSM^{+}06, Kad16, KSM^{+}18, LSZH06, PWB12, PSMB20, PPR17, SF14, TRKS^{+}17, VHV20, VW14, WDML22, Wei12, DZP^{+}07a, GSW^{+}06a, Hoo08, LKF09, Tre08]. Distributional [KK16, KCK^{+}21]. Distributions [AJGM22, BGQ21, CMG14, CCZ17, CFLN18, FD14b, GM13a, GLM18, HW13, Kom15, LMLM14, QSF09, RS13, RMP12, RSSSSL21, Scu13a, TFHP18, WOPF11, Wil18, YVSG18, vdL07, AO06, AVCGG08, CF10, FJS08, Gel06, GOO07, GB10, Hoe06, KS10a, MBC24]. Divergence [LCS^{+}14]. Divergences [GMY21]. Diverging [Wan17]. Divisible [Pas23]. DNA [CLM07]. Do [Lad06]. Does [DZP^{+}07a, Fie06a]. dominating [MM07]. Dominici [CS07, RC07]. Dose [HYY12, LTY21, LN13]. Dose-finding [LN13]. Dose-Schedule [LTY21]. Doses [HYY12]. Doubly [GMS16]. Draper [Gel06, KN06, Lam06]. Driven [WGBS17, ALR21]. Dropout [MCMK20]. Drton [Ano14a, CB14, Sha14a]. Drug [LN13]. Drugs [HYY12]. Duplication [TSL20]. Dyk [LG06]. Dynamic [AQ17, Bha07, CW07, CZGV19, CSN^{+}15, FS11, FSG08, GTHB19, GMB20, GW16, HMZ^{+}22, JP08, Kow21, LHE^{+}20, LLW21, LC23, OGPD19, RM21, SC17, WRC11, LW09, RtH08]. Dynamical [SCHT13b]. Dynamics [OBS13, VDF^{+}12].

Early [DD07, SOL^{+}12]. Ecological [GSWF19]. Econometric [BHvD17]. Economics [Poi06]. Edge [CRAL23]. Edge-Weighted [CRAL23]. Edgeworth [Wen10]. Editor [Car08, Car09]. Editor-in-chief [Car09]. Editorial [Ano16a, Ano16b, Ano23a, Ano23b, Ano24a, Ano24b]. Effect [HCPH18, KCK^{+}21, PW19, SM17, VDP15, DZP^{+}07a]. Effective [MKIM24, MTM12]. Effects [BJS23, BLE16, BB24, BKD21, HMC20, HD12, KDV09, MHSC16, SC06, WGBS17, BVN09, CKS07]. Efficiency [DT18]. Efficient [BHW18, JV23, JGP^{+}19, KMB19, LAE^{+}09, MNS^{+}20, Pra16a, SCHT13b, SOL^{+}12, TDY18, TAN^{+}18, TdVPAB17, Wan12]. Elaborate [MW19, WOPF11]. elastic [LL10]. Electromyographic [AFRB14]. Elemental [TFHP18]. Elicitation [ADGJ^{+}12a, DM07a, DL07, GOO07]. Embedded [SN18]. emerging [JKNR09]. Emphasis [LBB09]. Empirical [SK17, SHMM23, XLH16]. Empirically [Ste15]. Employment [SW22]. Emulation [Bha07, Gu19, IW19, JV23, OM20, LW09]. encompassing [AM07]. Endogenous [Kob17]. Energy [vDCE^{+}06]. Enhancements [WWACH16]. enriched [WMP11]. Ensemble [DEGP22, WOJL22].

Ensembles [LBLS22, YSB22]. Entity [Ste15]. Epidemic [AKO19, AKO23, CO08]. Epithelial [HCH06]. equality [DvdBWL24]. Equation [BKD21]. Equations [CCCG16a, DCKW08, WCO20, YSLR14, AZ10, DGS09]. Equilibrium [RRJW20]. Equivalence [CCVP18, SF14]. Equivariant [Hof16]. ERGMs [YSB22]. Ergodic [MM07]. ERK [PW08]. Error [ADL12, CCDT^{+}22, HKLM10a, LM16, MKIM24, SY19, SC06, CG10, RB07]. Errors [HHHL18, HD12, KGGC10, Per07, RC17]. Estimate [MBB^{+}23]. Estimated [SHMM23]. Estimates [BCHJ19, WT06]. Estimating [BB08a, GSWF19, HMC09, KAL12, Kyu11, Leo11, MP18, Paj17, Sal18, WCKL18]. Estimation [BF21, BG13, DW13, FT12, FMO16, GMY21, GMP21, GMdPV21, Gop22, GMS16, GKMvCT14, HCPH18, HHHL18, JLM^{+}17, KK22, Mar24, PK24, Pol17, PBT^{+}21, RV14, SW22, Scr14, SHK07, SOL^{+}12, SG17, TSL20, TDC^{+}22, VGE19, Vie07, VHJS08, WG18, WC18, WS20, YZCC16, YSH18, BB10, BALO06, Chr09, CO08, DEJL11, Gri10, HS09, LKF09, RtH08, Ryd08a, Sco11, vdL11a]. Estimator [BDPW17]. Estimators [GMdPV21, DM07b]. Evaluating [GLJB23, Vie07]. Evaluation [HRT24, POWK24]. Evaluations [JGVM21]. Event [BHS14, CS16b]. Events [sC16]. Evidence [BHvD17, BG21, DMF16, GHO^{+}13, LR16, DEJL11]. Evidentiary [Sha14b]. Evolutionary [BR10, KAL12]. Evolving [LC23]. ex [POWK24]. Exact [DPM16, Hoo08, Hut07, vES21]. examples [JMKW09]. exchangeable [Woo14]. Exclusive [CB21]. Existence [NJ21]. Exogenous [GR20]. exoplanet [FMV11]. expansion [Wen10]. Expansions [NSAL^{+}21]. Expected [FT12, FND15, FNP18, FN22, HRT24]. Expensive [WSD22]. Experiment [Gin07, LSZH06]. Experimental [AGG16, AFRB14, KDG21, MNS^{+}20, PHG23, RDP16, WHG^{+}06, ZM23]. Experiments [AFRB14, DT18, LKOB19, OS24, OM20, OM22, SXR06, WWACH16, WHG^{+}06]. Expert [ADGJ^{+}12a, sC16, DM07a, DL07, PVC20]. Explaining [GSW^{+}06a]. Explanatory [Bic20]. Exploitation [SMBS23]. Exploiting [FMO16]. Exploration [SMBS23, BR10]. Exponential [DP12, RR12, ZL24]. Exposure [CCL^{+}09a, CT11]. Expression [HCH06]. Extended [JHB22, RB07]. Extension [HdHG21]. Extensions [BJQ12]. External [MTS^{+}21]. Extracting [WG15]. Extrapolated [RCMO22]. Extrinsic [LMCD19].

Factor [APRS22, BJM^{+}22, BCLM24, FSG08, GL22, KC23, LM21, ML22, MF19, MVG20, OK22, Wei12, Zho18]. Factorization [ZG19]. Factors [AKO19, BE13, CCDT^{+}22, CS16a, DvdBWL24, HC17, HdHG21]. Failure [DD07]. Families [DP12, RR12]. Family [CS16b, ZWF^{+}18]. Faraday [SHK07]. Fast [BF21, CCZ17, Gop22, GHM^{+}23, KJP24, SLB^{+}21, ZG19, vES21]. favor [TGM09]. Fay [Pol17]. Feature [BPJ13, pD20, LZN08]. Features [Mar24, SG16, BP08, JP08]. Fecundity [KSLP12a]. Feedback [OBS13].

Ferreira [CD15, CC15, Zid15]. Field [CLMtH15, WOPF11]. Fields [FM18, GRM^{+}09]. file [Ano11a, Ano12a, Ano13a]. Filter [DEGP22]. Filtering [SS08, VGE19, pD20]. Filters [YHW16]. Financial [WRC11]. Finding [HYY12, LC22, ZS09, LN13]. Finegold [Ano14a, CB14, Sha14a]. Finite [FT12, FSMWG21, PS15, YSB22]. Fisher [Ald08, FMM18]. Fit [BPSS15, HC17, Vie07, CCQ11]. Fitting [CCL^{+}09a, TN14, ZG19, BD06a]. Fixed [SK13]. Fixed-Form [SK13]. Fleming [ALR21]. Flexibility [CBR23]. Flexible [ALC24, BC11a, KSLP12a, LHE^{+}20, MHSC16, QSF09, VDP15]. Flow [GCM24]. Flows [BG21, TSA20]. Flyer [WHG^{+}06]. fMRI [CSN^{+}15, LBBJ16, SLB^{+}21]. Focus [KEMM19]. Focused [DRH17]. Folding [VGS^{+}21]. Folklore [HdHG21]. food [Tre08]. Forecasting [GW16, HK18, KCR19, OGPD19, PPG08, FS11]. Forest [OBS13]. Form [SK13]. Formation [VGB10a]. Forms [MRG19]. Formula [HK18]. Formulations [TSA20]. found [CT11]. Frailties [HJZ12]. Frame [SF14]. Framework [CNR15, DWM^{+}21, LG24, TN14, TSL20]. Fraud [BMBV22]. Free [Hof16, TDC^{+}22, DEJL11, GRM^{+}09, Pac06, WFR11a]. free-knot [Pac06]. Freedom [VW14]. French [CT11]. Frequency [YHW16]. Frequentist [CEMR12, CB21, Was06]. Full [Des13, CZ10]. Fully [DK15, DT18]. Function [LLPR06, Pra17, RRJW20, LKF09]. Function-Specific [RRJW20]. Functional [BHJ18, CL24b, EH17, GABP19, HZ22, HGXS23, JP16, KJP24, KCR19, Kow21, KC23, LJCB14, MZMK24, SW22, SCFJ14, SG16, SG17, YZCC16, ZD17, KS10a, vdL11b]. Functional-Coefficient [SCFJ14]. Functions [ANRSL16, BPJ13, CDH16, GABP19, Hut07, PQ15, PBT^{+}21, SMBS23, MM07]. Fusion [PW19].

Galactic [SHK07]. Galaxy [VGB10a]. Gamerman [CD15, CC15, Zid15]. Gamma [ALC24, BDW^{+}24, HOHS24, NB18, Qia18, BC11a, CLM07, GB10, Nee19]. Gamma-Distributed [HOHS24]. Gamma-Pareto [ALC24]. Gammas [Han06]. Gaussian [AZ13, BWD20, BCLM24, CBR23, CKY20, FND15, Gu19, GL22, HSH^{+}21b, JGVM21, JB18, KJP24, KS10a, KFF19, LYL24, LG14, LMC20, LMCD19, MF22, MW15, MBB^{+}23, NS18, OJP23, PVC20, QNK23, Raj19, RV14, Scr14, SHK07, TZG10, VHV20, WWACH16, ZKRVA18, ZHA23, ZKS23]. Gaussian-Process [NS18]. Gelfand [Fer12, GB12, Hoe06, LC12, Ver06]. Gelman [Ber08, Kad08, Sen08, Was08]. Gene [HCH06, NJM18, Bar11]. General [GTGC16, HSBvdW17, HSC12, JS24, MZMK24, Ski06, WB18, CLPT10, CF10, WT06]. Generalised [ALC24, Pol17]. Generalized [BLE16, BB24, BH11, Bra22, CL24a, FNP18, FSMWG21, GKMvCT14, HSC12, LG24, TN14, VDP15, WM23, CHIK08, KN06, MPK10, RH11, Yin09a]. Generating [HRW18]. Generation [XLY^{+}13]. Genetic [BPSS15, CS12, XS07]. Genomics [RL14]. genuinely [dBPSW08]. Geographical [OMC19]. Geographically [DM07a, LG24]. Geographies [BR13]. Geometric [PMG14]. Geometry [dCPB19]. Geostatistics [Ban17, dG15]. Gibbs [CCO24, GRM^{+}09, HR20, NS23, SSLD23, ZR21]. Gibbs-Type [CCO24, HR20]. Girolami [BCT^{+}16, Das16, Lys16, MYGE16]. Girsanov [SS08]. Givens [PJM^{+}21]. Global [GR23, HIS22, PS12, PBT^{+}21, ZLV24, ZB18, Ngu10]. Global-Local [GR23, HIS22, ZLV24, ZB18]. Goldstein [Chr06, Dra06, Fie06a, Kad06, Kas06, Lad06, O’H06, Was06]. Goodness [CCQ11, HC17, Vie07]. Goodness-of-fit [CCQ11]. GPU [GW16]. GPU-Accelerated [GW16]. grade [GM09]. Gradient [COIG19, TSA20, ZL24]. Graph [AQ17, BCHJ19, CKY20, CL24b, CS16b, DBHG19, HHT24, LCL^{+}14, ZL24, vdBDB23]. Graphical [BG06, CW07, CC21, CAS^{+}19, FD14b, GW16, KMB19, LMC20, MMJ16, MW15, MG20, NJM18, NTL19, NPKC14, OJP23, Scu13a, Wan12, Wan15]. Graphs [BHS14, CCVP18, CRAL23, WRC11, BC11b]. Group [BDW^{+}24, DD18, GC17, GHM^{+}23, LMPS17, YN20]. Grouping [RL14]. Growing [RCMO22]. Growth [Poi06]. Guaranteed [NS23]. Guide [WSD22]. Guided [PSC23].

Half [PS12]. Half-Cauchy [PS12]. Hamiltonian [BCJ21, MBC24, ND20, ZSZ18]. Hastings [Pra16a]. Hazard [DD07]. Hazards [HJZ12]. heavy [GOO07, Tre08]. heavy-tailed [GOO07]. Held [vdL11a]. Hellinger [She14]. Herriot [Pol17]. Heterogeneity [SM17]. Heterogeneous [APRS22, HMC20, HLC20, PQ15, VHV20, ZD17]. Heteroscedastic [SCFJ14]. Heteroscedasticity [KR21]. Hidden [BG06, FWLH06, HAJF23, KCG15, MNPM20, XS07, Ryd08a]. Hierarchical [APL24, AZ10, BFPT22, BCR20, BGQ21, BHJ18, Bra22, BS21, CKG20, CL24a, CI06, CCL^{+}09a, CAS^{+}19, DD07, EDF^{+}19, Gop22, GCM24, GB17, JMW09a, KFF19, LLPR06, MM16, MTM12, OGPD19, PVC20, RMHR15, RSST17, YS07, YZCC16, YPVG22, YH11, GSW^{+}06a, Gel06, MS07a]. Hierarchy [SN07]. Hierarchy-Based [SN07]. High [APD19, Ban17, BHW18, CKG20, CRAL23, GC17, GR23, Joh13, LL20, LL23, LYL24, LAE^{+}09, ML22, MRG19, OK22, QNK23, RGC20, SN18, SKG15, SKN24, YN20, vDCE^{+}06, LN08, MT09b, Spi08]. High-Dimensional [APD19, Ban17, BHW18, CKG20, GR23, LYL24, MRG19, OK22, SN18, SKG15, SKN24, GC17, LL20, LL23, QNK23, MT09b, Spi08]. High-Energy [vDCE^{+}06]. high-order [LN08]. Higher [RSV14]. Higher-order [RSV14]. histology [JMKW09]. Historical [HSC12, MTS^{+}21]. History [KAL12]. hitting [JMW09a]. HMM [SN18]. Hoff [All11, Fre11]. Hogg [Hen10, SS10]. Holmes [vdL11a]. Homogeneity [HGXS23]. Homogeneous [BGQ21, FLN^{+}16]. Horseshoe [BDPW17, DG13, vdPSvdV17]. hosts [CH09]. HPD [DM07b]. Human [HCH06, KSLP12a, SMBS23]. Hyper [BH11]. Hyper- [BH11]. Hyperplane [CCZ17]. Hyperplane-Truncated [CCZ17]. Hypotheses [CB21, Sal18]. Hypothesis [BE13, GTGC16, HCGS15, KDV09, SY17].

I-II [LTY21]. Ice [ZC20]. Identification [HCH06]. Identifying [MS07a]. Identity [Paj17, Wen10]. Ignorable [MRB12, MCMK20]. II [LTY21, PB20, SY17]. Illustrated [Vie07]. Image [ZLV24, ZJLC10]. Image-on-Scalar [ZLV24]. Images [LG14]. Imaginary [CS13]. Imaging [BHJ18, LJCB14]. immunofluorescence [JMKW09]. Impact [JTC22, SHMM23, TGK^{+}11, CH09]. implications [Pac06]. Implicit [KS19, KDG21]. Implied [CLMtH15]. Importance [BH07, LR16, AZ10]. Improve [ND20]. Improved [FI09, VGS^{+}21]. Improving [DT18, GKSG21, SN07]. Imputation [dTM10, CCQ11]. incidence [CH09]. Income [HGXS23]. Incomplete [GL22, dTM10]. Inconsistency [GvO17]. Inconsistent [Chr09]. Incorporating [HSC12, PKL^{+}11, RL14, SR16]. Incorporation [MTS^{+}21]. Independence [NTL19, NPKC14]. Independent [MTM12, SPD19]. Index [DLPS20, RGC20, WRC11]. Indexed [SW22]. Indian [CGZ16, HR20, WDML22]. Indices [ATF23]. Indirect [RDP16]. Individual [PPG08, VDF^{+}12, CT11]. Individual-Level [VDF^{+}12]. Induced [GLC24, HCH06, ZJLC10]. Inducing [CRAL23]. Inequalities [BE13]. infection [CH09]. Infectious [MNS^{+}20, VDF^{+}12, JKNR09]. Infer [LMC20, BP08]. Inference [BF21, BLE16, CL24a, CS12, CVCB23, CC21, sC16, Chr24, CH09, DR16, DPM16, DD18, Gop22, GMS16, GL17, GHM^{+}23, GB10, GvO17, GRM22, HMC20, HRT24, HOHS24, HSBvdW17, HSH^{+}21b, HD12, JGVM21, JP16, KG09, LG17, MCW10b, MC15, MNPM20, MG20, MM13a, MQ22, Ngu10, NGT19, PSMB20, PJM^{+}21, QMRM08, RS13, RS14a, RDP16, SCHT13b, SPG15, SNMS23, TDC^{+}22, WGBS17, WM23, XS07, ZC20, dCPB19, vES21, AVCGG08, ALR21, BJ06, Fie06b, GP10, HKLM10a, JHB22, PW08, RB07, SB11, VR11, WMP11, WFR11a]. Inferences [AE17, BSPD23, RW08]. Inferring [LSZH06, SFZ08a, ZM23]. Infinite [AGG16, MVG20, PWB12, RR12]. Infinitely [Pas23]. Inflated [Nee19]. Influence [vdL07]. Influential [MS07b]. Influenza [OGPD19]. Information [CFRT06a, sC16, Gag23, Gin07, HSC12, KDG21, RL14, SR16, SMBL19, US16, Vie07]. Informative [CEMR12, CAV23, CHMK22, HBJ14, PHOD21, WS20, Wil18, JD08, She14]. Informed [BHS14]. Inhomogeneous [DHDC12]. INLA [DWM^{+}21, SHMM23]. Instrumentation [vDCE^{+}06]. Insufficient [LML21]. Insurance [CGS22]. Integer [CSN^{+}15, DPM16]. Integer-Valued [DPM16]. Integral [CKS07, CS13]. Integrated [GSWF19]. Integration [APRS22]. Integrative [NJM18]. Intensities [DRRS17]. Intensity [DR16, Sco11]. Inter [PKL^{+}11]. Inter-day [PKL^{+}11]. Intercept [SLAV13]. Interdependence [BGP15]. Interference [OS24]. Intermediate [ND20]. Interpretation [LC17, SLAV13]. Interval [JNBQ13]. intifada [JP08]. Intraclass [MF19]. Intractable [DPM16, FMO16, OM20, RDP16, VGE19, VDP19]. Intrinsic [KFF19, TRWFB17]. Intuitive [FHK^{+}20]. Invariant [DM07b, DP12, HdHG21, SF14]. Inverse [AZ13, BH07, BDW^{+}24, HOHS24, JYL17, MNPM20, Qia18, RSST17, SMBS23, Scr14]. Inverse-Gamma [BDW^{+}24, HOHS24]. Inverse-Gaussian [AZ13, Scr14]. Investigation [BG21]. Investigations [BS21]. irreducible [SB11]. issue [Ano06e, Ano06f, Ano06g, Ano06h, Ano07e, Ano07f, Ano07g, Ano07h, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10e, Ano10f, Ano10g, Ano10h, Ano11f, Ano11g, Ano11h, Ano11i, Ano12f, Ano12g, Ano12h, Ano12i, Ano13f, Ano13g, Ano13h, Ano13i, Ano14f, Ano14g]. Item [BBB06, WC18]. Iterative [ZG19].

JAGS [SHMM23]. Jain [Dah07, Mac07, Rob07]. Jeffreys [LCS^{+}14, RS14a]. Jensen [AB09, Gli09, QM09]. Joining [GPL^{+}19]. Joint [Bra22, CL24b, FHK^{+}20, GR20, QPC24, TRKS^{+}17, VHV20, HvDH09]. Jointly [Gu19]. Judgements [WG15]. Jump [ALC24]. Jump-Activity [ALC24]. Jumps [ADP22].

Kalman [DEGP22]. Kernel [Scr14, SM19, XX20]. Kernels [GCM24, TDY18]. Kim [Sca12, Sta12]. Kinds [Kas06]. kinetic [PW08]. knot [Pac06]. Knots [BS14, Kyu11]. Known [JV23, MB12, AM07]. Kullback [Vie07].

Lag [HK22]. Lands [GSWF19]. Langevin [PSMB20]. Lans [HH11]. Laplace [SR17, LG12b, RV14, SRG13, TGM09, ZB18]. Laplacian [CKY20, LCL^{+}14]. Large [ADP22, APRS22, GL22, KK22, LL18, MCW10b, TAN^{+}18]. Lasso [Wan12]. Lassos [KGGC10, RC17]. Latent [CDL^{+}19, GDNJ18, GL22, HSH^{+}21b, LMC20, LC23, SR16, SMW19, SC17, SM17, SN18, ZL15, vdL11b]. Lattice [YHW16]. Laws [BJP12]. Leading [LCS^{+}14]. Leaks [XTMR17]. Learning [BG06, BWD20, CCVP18, GW16, LL23, MMN22, MW15, NTL19, PNNC17, PS17, SMBS23, Wan15, WPS23, WM23, XJC16, vdBDB23, CLPT10]. Legislation [WSDC13]. Leibler [Vie07]. Lens [MKIM24]. Level [VDF^{+}12]. Leveraging [CL24b]. Lévy [ALC24]. Life [WPCAV22]. Lifetime [Han06]. Likelihood [BF21, BAR23, DEJL11, GSWF19, JGVM21, KEMM19, LML21, OM20, Paj17, PB20, PNNC17, SF14, SHMM23, TDC^{+}22, WN21, XLH16, BD06a, CNR15, GRM^{+}09, KS10b]. likelihood-based [BD06a]. Likelihood-Free [TDC^{+}22, DEJL11, GRM^{+}09]. Likelihoods [DPM16, FMO16, MM16, PSC23, RDP16, VGE19, VDP19, WCKL18]. Limit [HZ22]. Limited [CCY13]. Limiting [EMS13]. line [BP08]. Linear [ATF23, BCLM24, BDW^{+}24, BH11, CL24a, FND15, FNP18, FN22, GDB20, GTHB19, GMB20, GHO^{+}13, GABP19, GW16, GvO17, HCPH18, HSH21a, HSC12, JP16, JB18, NTL19, Qia18, RMP12, RSST17, SK13, SS08, SN18, TN14, TK12b, VL20, Woo14, WN21, XX20, ZR21, CHIK08, KN06, Leo11, Pac06, RH11]. Lineups [BSPD23]. Link [MMW15]. Linkage [GRM22]. LIO [SMBL19]. Local [CKG20, CS16b, GR23, HIS22, LMLM14, LL23, SG16, ZLV24, ZB18, vdL07].

Local-Mass [LMLM14]. Localization [VGS^{+}21]. Locally [FM18, KCR19, MS07b, Ngu10]. Locally-Autoregressive [KCR19]. Location [RS14a, SHMM23]. Location-Scale [RS14a]. Log [FT12, JB18, MM16, NTL19, RMP12, ZKRVA18, FJS08, KS10b]. Log-Gaussian [ZKRVA18]. log-likelihood [KS10b]. Log-Likelihoods [MM16]. Log-Linear [JB18, RMP12]. Log-Normal [FT12]. log-spline [FJS08]. Logarithmic [CVCB23]. Logic [HSF20]. Logistic [GLM18, GP12, HBJ14, PWB12, RV14, SLAV13, TZG10, LN08]. Logit [TM17, vdL11a]. Long [HMC09]. Longitudinal [BJM^{+}22, GMP21, GR20, PS20, HvDH09]. Look [CCL^{+}09a]. Loss [FT12, LLPR06, VL20]. Loss-Based [VL20]. Low [DPM16, SMBL19, YMX23]. Low-Rank [YMX23]. lower [MM07]. Luce [HK18, JHB22]. Lum [Fer12, GB12, LC12]. Lung [XTMR17].

MacEachern [BJQ12]. machines [PS11a, PS11b]. Magnetic [BHJ18, LJCB14]. make [Fie06a]. Mallows [CAV23]. Manifold [PSMB20, PJM^{+}21]. Manifolds [LMCD19]. Manolopoulou [Rig10, Whi10]. many [MY08]. MAP [DM07b, RCLW17]. MAPK [PW08]. MAPK/ERK [PW08]. Mapping [DBHG19, MBBRB17]. Maps [HHG08, BP08]. Marginal [BLE16, BB24, BAR23, DEGP22, NTL19, Paj17, PNNC17, RSV14, SR16, WCKL18, SB11]. Marginally [HW13]. Marked [GDNJ18, TK12a]. Marker [CKY20]. Markov [CLMtH15, CCVP18, FM18, GPL^{+}19, HAJF23, HS09, JP08, KCG15, MG23, PMG14, PNNC17, PKLM10, Ryd08a, SPD19, TK09, TDY18, TdVPAB17, Wei12, XS07, XJC16, ZL24, ZWC^{+}16]. Markovian [MM14]. Mass [GCM24, LMLM14]. Massive [BP20, BM06]. Matching [KD12, ZSZ18]. material [Ano14b, Ano14c]. materials [BVN09]. Matérn [SLB^{+}21]. Matrices [BCHJ19, GMP21, GL22, HW13, LHE^{+}20, LL18, LL20, MP18, WC14b]. Matrix [CW07, MP18, PSMB20, XCPX22, YMX23, ZWDJ14, FI09]. Matrix- [MP18]. Matrix-Variate [CW07, ZWDJ14]. Max [HSH^{+}21b]. Max-and-Smooth [HSH^{+}21b]. Maximal [Raj19]. Maximum [PB20]. Maxwell [BF21, KSM^{+}06, KSM^{+}18, Kad16]. Maxwell-Binomial [Kad16]. mBART [CGMS22]. MCMC [BH07, DEGP22, LV22, LC22, NS18, NdVA^{+}20, PSC23, SCHT13b, SOL^{+}12]. Mean [Paj17, WOPF11, YZCC16]. Mean-Covariance [YZCC16]. Meaningful [WG15]. Means [BP07, FT12, Pol17]. Measure [Gin07]. Measurement [ADL12, HD12, SC06, CG10, RB07]. Measures [CAS^{+}19, FMM18, KK07, LCS^{+}14, Pas23, SHK07]. Measuring [CZ10]. Mechanisms [Pra16a]. Median [BBGR21]. Melding [GPL^{+}19, MG23]. Membership [GLC24, HLC20, GM09]. Memory [HMC09]. Merge [ZSM07]. Merging [JN07b, NS18]. Message [MW19]. Meta [BG21, BB24, OBS13, SHMM23]. Meta-Analysis [BG21, BB24, SHMM23]. Metabolites [HYDE21]. Metals [HCH06]. Method [COIG19, KAL12, Kyu11, MZMK24, NGT19, SN18, WB18, WCKL18, BM06, LZN08, MT09b, Yin09a]. methodology [GD09]. Methods [BP07, BKD21, CEMR12, FJM14, GHM^{+}23, LC22, LML21, Poi06, VL20, VHJS08, WM23, vDCE^{+}06, BD06a, CZ10, GRM^{+}09, JD08, OS09]. metrics [Scr14]. Metropolis [Pra16a]. Metropolized [SKN24]. Microarray [SXR06, CZ10]. Microbiome [SSML20]. micronutrient [DZP^{+}07a]. Minimax [LL18, GD09]. Mis [SNMS23]. Mis-Specification [SNMS23]. Mises [PS15]. Misinformation [Pac06]. Missing [BWD20, CFRT06a, DCKW08, DLPS20, MRB12, OS24, WT20, GP10]. Missingness [BHS14]. Misspecified [DW13, GvO17, RSM15, SRG13, SR17]. Mitra [APA^{+}13, CM13, Hof13, O’H13]. Mixed [BJS23, BKD21, DRH17, HD12, HLC20, JP16, PL16, TN14, WT20, WGBS17, Bar11, KN06, RH11]. Mixed-Effects [HD12, WGBS17]. Mixing [RRJW20]. Mixture [ADMO24, AJGM22, DRH17, GM16, Han06, HRW18, JN07b, LR16, MCW10b, MCMK20, Raj19, SW22, SM17, SMBL19, SM19, TK09, TK12a, XX20, ZKS23, CLM07, Gri10, JMKW09, WT06, YH11]. Mixtures [BGQ20, FN22, FSMWG21, GCM24, GL18, HOHS24, MB12, MVG20, NB18, Nee19, RVWG24, Scr14, SS11, Wan17, YSB22, AVCGG08, BJ06, CLPT10]. Modal [Dah09]. Model [APL24, ADL12, ADMO24, BBGR21, BBG12, BBB06, BF21, BLE16, BB24, Bra22, BAR23, BS21, CS13, CVL12, CMG14, CZGV19, Cas21, CS16b, CCL^{+}09a, CAV23, DCKW08, DM15a, DLPS20, DD07, GM16, GLC24, GC18, HJZ12, Hof06, HM23, HHG08, JN07b, JNBQ13, JGP^{+}19, Joh07, Joh13, KCG15, KMB19, LG17, LM16, LM21, LBBJ16, LG24, MM14, MMW15, MNS^{+}20, MDO18, MCMK20, MNPM20, NS23, OM22, PFS10, Per07, PKLM10, Pol17, Raj19, RW08, Ros22, SFZ08a, SXR06, SMW19, SOL^{+}12, SCFJ14, TM17, TAN^{+}18, VG23, Vir11, VDF^{+}12, WC14b, XCPX22, YZCC16, YMP13, ZSM07, ZG19, vES21, BR10, CKS07, CLM07, CT11, DEJL11, FMV11, FS11, GM09, GRM^{+}09, HvDH09, JHB22, LW09, MPK10, Pac06, RB07, WT06, vdL11a]. Model-Based [JGP^{+}19, Hof06, HHG08, PFS10, RW08]. Model-Fitting [ZG19]. Modeling [BHJ18, CGS22, CAS^{+}19, DK15, DGS09, EDF^{+}19, FD14b, GSWF19, GR20, Han06, HSBvdW17, HRW18, JYL17, LHE^{+}20, LC23, MCW10b, MHSC16, PCM19, PBT^{+}21, RGC20, SM19, TK12a, TRKS^{+}17, TFHP18, VHV20, WRC11, WSDC13, WB18, XS07, XTMR17, YN20, YSB22, ZKRVA18, ZD17, dCJHdC13, AO06, GSW^{+}06a, Hoe06, JMW09a, KS10a]. Modelling [CNR15, DG11, Des13, GB13, GL18, KR21, RdGvP06, Scu13a, ZWC^{+}16, JMKW09, LW09, Pac06]. Models [AKO19, AQ17, AKO23, BPSS15, BCR20, BHvD17, BG06, BJS23, Bha07, BWD20, BDW^{+}24, BKD21, BH11, BHW18, BR13, BPH21, CHG12, CL24a, CW07, CMG14, CC21, CFRT06a, CI06, Chr24, CSN^{+}15, DBHG19, DW13, DRH17, DM07a, DGMQ13, DPM16, DEGP22, FWLH06, FJM14, FND15, FNP18, FN22, GTHB19, GMB20, Gop22, GPL^{+}19, GL17, GKMvCT14, GHM^{+}23, GB17, GW16, GvO17, HAJF23, HMC20, HK18, HSC12, Hof16, HSH^{+}21b, HRW18, HD12, JV23, JP16, JLM^{+}17, JB18, KFF19, KD12, KDV09, KSLP12a, KCK^{+}21, KDG21, Kow21, KC23, KG09, LLW21, LMLM14, LJCB14, LR16, LMC20, LLPR06, LBB09, Ma17, ML22, MG23, MF22, MRB12, MMW15, MW19, MM16, MS07b, MMJ16, MW15, MTM12, MG20, NJM18, NTL19, NPKC14, OJP23, OK22, OM20, PQ16b, PVC20, PKLM10, PKL^{+}11, PL16]. Models [Pra16a, QPC24, QNK23, Rah16, RSM15, RVWG24, RCMO22, RMHR15, RS14a, RDP16, SR16, SM17, Sha21, SN18, SMBL19, SHK07, TN14, TRWFB17, TAN^{+}18, VGE19, VHJS08, VDP19, VDF^{+}12, WRC11, Wan12, Wan15, Wan17, WC18, WGBS17, WG15, Wil18, WN21, XX20, XJC16, YPVG22, ZR21, ZL24, ZWF^{+}18, ZKS23, AZ10, Bar11, BC11a, BD06a, CCQ11, CHIK08, CO08, Dah09, Gel06, Gri10, HS09, HHC07, HH06, KN06, LKF09, LN08, LZN08, MS07a, MAL11, RD11, RH11, Ryd08a, WFR11a, YH11, vdL11a, vdL11b]. Modes [vdL07]. Modularization [LBB09, OBS13]. Modularized [LG24]. moments [Yin09a]. Monitoring [HAJF23]. Monni [CGM09, Fra09, Li09, Ste09]. Monotone [CGMS22, MM07]. Monotonicity [SRA23]. Monte [BCJ21, ND20, TDY18, AZ10, BM06, BW15, BCJ21, DT18, FT13, HS09, MKIM24, MBC24, PMG14, PKLM10, Ryd08a, SPD19, TdVPAB17, WCKL18, Wei12, YSH18, ZSZ18, ZL24]. Monthly [SW22]. Mortgages [PPG08]. Most [NJ21]. Motivated [Ste15]. Movements [PKL^{+}11]. MR2383247 [HG08, Rou08]. MRI [MZMK24]. Müller [APA^{+}13, CM13, Hof13, O’H13]. Multi [CBC23, CAD^{+}23, FWLH06, FMO16, IW19, POWK24, QMRM08]. Multi-Armed [CBC23]. Multi-Core [FMO16]. Multi-Scale [FWLH06]. Multi-Season [QMRM08]. Multi-State [CAD^{+}23]. Multi-Step [IW19]. Multi-Step-Ahead [POWK24]. Multidimensional [CGMS22, MBBRB17]. Multigrid [ZR21]. Multilevel [CGS22, DCKW08, GKSG21, ZR21, BD06a]. Multimodality [KK07]. Multinomial [BR13, BPH21, Wil18, HH06, TGM09, vdL11a]. Multiple [BPSS15, BF17, Bra22, BG13, BR13, GTGC16, GLC24, GBGTR19, JV23, KDV09, KCG15, LG12b, MC07, MZMK24, MF19, PCM19, Sha21, Woo14, WN21, BP08, CCQ11, CH09, HHC07, WFR11a]. Multiple-Shrinkage [BR13]. Multiplex [APL24]. Multiplicative [DR16, DRRS17, vdL07]. Multiplicity [CB21]. Multiregression [CSN^{+}15]. Multiresolution [DD07]. Multiscale [LG14, MBC24]. Multivariate [APS18, ADMO24, BB24, BHW18, CCZ17, CGS22, DHDC12, LLW21, LMPS17, MC07, NGT19, OM20, PCM19, PL16, QPC24, RSSSSL21, SC06, SSML20, TFHP18, VHV20, WPCAV22, Woo14, FS11, GP10, Hof11b]. Musio [GMR15, HP15, KB15]. Mutual [KDG21]. Mutually [CB21].

naive [LZN08]. NCoRM [GL18]. Neal [Dah07, Mac07, Rob07]. Near [BHvD17, SHK07]. Near-Boundary [BHvD17]. Nearest [ZKS23]. Nearest-Neighbor [ZKS23]. Necessary [SKG15]. needlet [Sco11]. Negative [Nee19, ZWF^{+}18, Zho18]. Neighbor [ZKS23]. Neighborhood [CRAL23]. neonatal [DZP^{+}07a]. Nested [CDL^{+}19, CS13, CFH23, Gop22, HHHL18, HRW18, NdVA^{+}20, Ski06]. net [Hoo08, LL10]. Network [AQ17, BG21, CKY20, CL24b, CHMK22, GR23, NJM18, PS20, PNNC17, RCMO22]. Networks [ATF23, APL24, BG21, CSN^{+}15, DD18, HLC20, LC23, Mad07, MMN22, RdGvP06, SC17, YSB22, ZHA23]. Neural [CHMK22]. Neuronal [RdGvP06]. Neutral [CLMtH15, Spi11]. Neutral-data [Spi11]. neutron [HKLM10a]. Next [XLY^{+}13]. NMR [HYDE21]. Noise [PKL^{+}11]. Noised [LG14]. Noisy [JGVM21, LKOB19, RSST17]. Nominal [DRH17]. Non [BJM^{+}22, CBR23, CS13, CKG20, CS16b, Gop22, KJP24, MRB12, MCMK20, NJ21, SRA23, SS08, She14, SN18, Woo14, ZKS23]. Non-Central [NJ21]. Non-exchangeable [Woo14]. Non-Gaussian [CBR23, KJP24, ZKS23]. Non-Ignorable [MRB12, MCMK20]. Non-informative [She14]. Non-Linear [SS08, SN18]. Non-Local [CKG20, CS16b]. Non-Nested [Gop22, CS13]. Non-Parametric [BJM^{+}22, SRA23]. Nonadherence [OS24]. Noncompliance [FMM18]. Nonconjugate [JN07b]. Nonconvex [ZL15]. Nonignorable [WT20]. Noninformative [HW13]. Nonlinear [HD12]. Nonlocal [CL24a, SSML20]. Nonparametric [CDL^{+}19, CZGV19, sC16, Chr24, DK15, DG11, DGMQ13, DHDC12, FH17, GOO07, GBGTR19, HC17, HK22, HCGS15, JYL17, KK22, KEMM19, LMLM14, LKF09, LC23, MM14, MM13a, NBCC14, PMS24, PBT^{+}21, RD11, SPG15, Vie07, XX20, XLY^{+}13, XTMR17, Zho18, dCJHdC13, BALO06, CT11, WMP11, YH11]. Nonparametrics [GLJB23, Tre08]. Nonparanormal [MG20]. Nonstationary [KK22]. Norm [MM16]. Normal [ALC24, BJS23, BP07, CCZ17, FT12, FN22, GCM24, GHM^{+}23, GGPM19, HSBvdW17, HD12, PWB12, Qia18, vES21, GB10, WT06]. normal-gamma [GB10]. Normal-Generalised [ALC24]. Normal-Inverse-Gamma [Qia18]. Normal-Normal [GHM^{+}23]. Normalization [VGS^{+}21]. Normalized [AZ13, CAS^{+}19, Ros22, Scr14]. Note [KSM^{+}18, Car08, Car09]. Novel [ALC24, HSF20]. NRMIs [FLN^{+}16]. Null [CBC23, MCB23]. Number [Kyu11, MB12, VW14, Wan17, BB10, CO08]. Numbers [TGK^{+}11]. Numerical [CCDT^{+}22, Joh13].

Object [GDNJ18]. Objections [Gel08a]. Objective [ADL12, BB10, Ber06a, BBS15a, BLE16, BB24, CCVP18, CFLN18, HSH21a, KFF19, Lad06, LVW20, MC15, VW14, Fie06a, Kad06, Was06]. objectivity [Dra06, Gol06b]. observability [AM07]. Observations [HOHS24, MS07b, FMV11]. Observed [AKO19, AKO23, DR16, JTC22, MNS^{+}20, SS08]. obtained [GD09]. Occam [Bic20]. Occupancy [TRWFB17]. Old [BP07]. Omnibus [SMBL19]. One [GC17, HK18, BM06, BVN09, CKS07]. One-Group [GC17]. one-pass [BM06]. one-sample [BVN09]. one-way [BVN09]. Only [FJM14]. Open [GSWF19, LC17, XS07]. Operations [WSD22]. Opinion [DM07a, DL07]. Opinions [ADGJ^{+}12a]. Optimal [AE17, AGG16, DT18, GMY21, JB18, LL18, LYL24, RDP16, ZM23, dG15, pD20]. Optimality [GC17]. Optimization [IW19, LKOB19, SMBS23]. Optimize [LTY21]. Optimizing [HMZ^{+}22]. Optional [HdHG21]. Oracle [JL19]. order [LN08, RSV14].

Ordered [Kow21]. Orderings [BSPD23]. Orders [ANRSL16]. Ordinal [CBC23, DRH17, MMW15, Rah16, SRA23]. orientations [BVN09]. Orthogonal [GL22]. Other [LCS^{+}14]. our [LC22]. Outcome [CBC23]. Outcomes [LTY21, OS24]. Outlier [SS11]. Outliers [GDB20, MS07a]. Overall [BBS15a]. Overlapping [RVWG24].

Page [GPP16, GL16, RF16]. Paintboxes [BPJ13]. Paired [dTM10]. Pairwise [CBC23]. paleoclimate [BC11a]. Panel [LM16, LM21]. Panels [ADP22]. Parabolic [RSST17]. Parallel [JGVM21, SOL^{+}12]. Parameter [Des13, HS09, HHHL18, HMC09, PS12, SLAV13, SOL^{+}12, TdVPAB17, VHJS08, WC18, YSH18]. parameterization [HHC07]. Parameters [FHK^{+}20, KK16, RC17, Wan17, Gel06, LN08, MAL11, PW08, TGM09]. Parametric [BJM^{+}22, DW13, KEMM19, SRA23, VDP19, QMRM08]. Pareto [ALC24]. Partial [OJP23, XX20, AM07]. Partially [AKO19, AKO23, DR16, MNS^{+}20]. Particle [BKD21, CLPT10, LSZH06, SS08]. Partition [LAE^{+}09, PHOD21, PQ16b, QPC24, Raj19, Dah09, MAL11]. Partitioning [MZMK24, MT09b]. pass [BM06]. Passing [MW19]. pathogens [CH09]. Paths [RC17]. pathway [PW08]. Pathways [CCL^{+}09a, MMJ16]. Patterns [DD07, LG17, WPCAV22, CG10, GSW^{+}06a]. PDEs [RSST17]. Penalization [ZL15]. Penalized [KGGC10, ZB18]. percentiles [DZP^{+}07a]. Perfect [BFPT22, MB12]. Performance [FJM14, JMW09a]. Permeability [ZJLC10]. Permutation [Chr24]. Persistence [Mar24]. Personalised [DWM^{+}21]. Personalized [HMZ^{+}22]. Perspective [PS17, Ryd08a]. perspectives [Hoe06]. Perturbation [SM19, vdL07]. pesticides [CT11]. Phase [AJGM22, LTY21, SY17]. Phase-Type [AJGM22]. Phylogenetic [CGZ16, MKIM24, ZWC^{+}16]. Physical [HAJF23]. Piece [RS14a]. Piecewise [Hut07]. Pitman [ADP19, BFPT22, Scr14]. Pivotal [Joh07]. Plackett [HK18, JHB22]. Plate [WHG^{+}06]. Players [BSPD23]. Point [BGQ20, KD12, LG17, MM14, PCM19, QPC24, WG18, CG10, JMKW09, KCG15]. Poisson [KSM^{+}18, BF21, DHDC12, GDNJ18, KSM^{+}06, TK12a, ZL15]. Polson [Han11, MCG11, SYvD11]. Pólya [Ma17, Nee19]. Polynomial [BPSS15]. Polynomials [XX20]. Pool [RMP12]. Pooling [CVCB23]. Pools [PPG08]. Poorly [CEMR12]. Population [BG13, DvdBWL24, EDF^{+}19, TSL20]. Populations [GM16, GSWF19, LYL24]. portfolio [GP10]. position [BP08]. Positive [WC14b]. Positive-Definite [WC14b]. Possibly [Kad16]. Post [BCHJ19, LLL23, POWK24]. Post-Processed [LLL23]. Post-Processing [BCHJ19]. Posterior [AKO23, BFPT22, CKG20, CCDT^{+}22, CGZ16, CFH23, DRRS17, FMM18, FND15, FNP18, FN22, GHM^{+}23, JB18, KS10b, LG17, ML22, MM16, MCB23, OK22, PSMB20, PK24, PHG23, RSM15, RR12, Ros22, RSV14, SSLD23, SK13, Scu13a, SF14, SHMM23, SKG15, SRG13, SR17, TM17, TGM09, Wan12, Wei12, WG15, vdL07, FI09, GD09, RM08]. Posteriori [Raj19]. Posteriors [BCHJ19, HM23, LLL23, NS23].

Poststratification [GKSG21]. Potts [MNPM20]. Power [BJP12, CI06, FND15, FNP18, FN22]. Power-Expected-Posterior [FND15, FNP18, FN22]. Powerful [NJ21]. Practice [Gol06a]. Pratola [CLH^{+}16, Gra16, Han16]. Pre [LBBJ16]. Pre-surgical [LBBJ16]. Precision [BCHJ19, LL20, HHC07]. Predicting [SHG^{+}10]. Prediction [ADP22, CCY13, EH17, HvDH09, LLW21, POWK24, SW22]. Predictions [PQ15, San12b]. Predictive [AKO23, ALR21, FMM18, GMY21, Kom15, LG17, MCB23, NDME18, YVSG18, Cla10, TGM09]. Predictors [PW19, PHC17]. Preferential [dG15]. pregnancy [HvDH09]. premiums [GD09]. Prepayment [PPG08]. Presence [CGS22, OS24]. Preserving [LMLM14]. Price [PKL^{+}11]. Principal [SG17]. Principles [Gol06a]. Prior [AE17, ADGJ^{+}12a, BPH21, CKY20, CMG14, CZGV19, CBC23, CI06, CFLN18, DG13, DL07, EM06, Gag23, Gel06, GLM18, GLJB23, Gu19, HW13, JTC22, KDV09, LMLM14, MRG19, MTM12, MP18, NSAL^{+}21, PPR17, PS12, RMP12, RSSSSL21, SR16, Scu13a, SN07, VW14, VL20, Wil18, XCPX22, ZLV24, ZHG^{+}16, GOO07, GB10, KN06, KS10a, Pac06, TGM09, WMP11]. Prior-Data [AE17, EM06, NSAL^{+}21]. Priors [APD19, ANRSL16, BS14, BBS15a, BHJ18, Bic20, BH11, CBR23, CDL^{+}19, CCO24, CS13, CKG20, CL24a, CS16b, CAV23, CHMK22, FM18, FND15, FNP18, FN22, FHK^{+}20, FCP09, GKSG21, GTGC16, GC17, GB13, GB17, GR23, HIS22, HBJ14, HSC12, HZ22, JS24, JB18, KFF19, KK16, LVW20, LCS^{+}14, MBB^{+}23, PHOD21, PSMB20, PB20, RM21, RS14a, She14, SMBL19, SSML20, SLB^{+}21, SKG15, Ste15, Wan17, XLH16, ZWDJ14, ZL15, ZB18, CKS07, CHIK08, Gri10, RB07]. Probabilistic [HK18]. Probabilities [Ros22]. Probability [BBGR21, BPJ13, EMS13, KK07, NTL19, DT09, RM08]. Probit [BR13, BPH21, CC21, RVWG24, Bar11, RD11]. Problem [BP07, RSST17]. Problems [BH07, CCY13, GC17, IW19, OMC19, PS15, GB10]. Procedure [GBGTR19]. Procedures [LNR19]. Process [AZ13, ADP19, AJGM22, ALC24, BFPT22, BGQ21, BGQ20, BWD20, CZGV19, DHDC12, GDNJ18, Gu19, HRW18, JN07b, JGVM21, KDV09, KCG15, LG12b, MBB^{+}23, MCMK20, NB18, NS18, PVC20, PL16, Raj19, RV14, RM21, RDP16, Scr14, SMBL19, SHK07, SS11, TK09, TZG10, XS07, ZWDJ14, BC11a, BJ06, JP08, KS10a]. Processed [LLL23]. Processes [BJQ12, BJP12, CBR23, CVL12, CGZ16, CAD^{+}23, DR16, DRRS17, EDF^{+}19, GMdPV21, GL22, HR20, KCR19, LMCD19, MF22, MNS^{+}20, PHOD21, TK12a, TRKS^{+}17, VHV20, WWACH16, ZKRVA18, ZL15, ZKS23, ALR21, JMKW09, MPK10, MM07, RD11, SB11]. Processing [BCHJ19]. Procrustes [KD12]. Produce [BCHJ19]. Product [MAL11, PQ16b, Dah09, Hof11b]. Prognostic [ATF23]. Programming [CSN^{+}15]. Projected [GGPM19, HSBvdW17, MBB^{+}23]. Projection [TZG10]. Pronged [MRB12]. Propensity [SNMS23]. Proper [DM15a]. Properties [AZ13, ALC24, DG13, GTGC16, JL19, Kom15, SFZ08a, WT06]. Proportional [HJZ12]. Proportions [BBG12, MJW08]. Proposal [GvO17, Pra16a, TDY18]. Proposals [PSC23, SPD19]. Propriety [MM16, TM17]. Prostate [MZMK24]. Proton [LSZH06]. Pseudo [DEGP22, PNNC17, SB11]. Pseudo-Likelihood [PNNC17]. Pseudo-Marginal [DEGP22, SB11]. Public [BR13, GSWF19]. Pure [ALC24]. Pure-Jump [ALC24]. purpose [CF10]. Pursuit [HGXS23].

Quadratic [FT12]. Quantification [CCDT^{+}22, CCCG16a, HYDE21, SHMM23, YMX23, vdPSvdV17]. Quantifying [JTC22]. Quantile [BGP15, DL07, GMB20, Kob17, LG12b, Rah16, SRG13, SR17, TK12b, VDP15, WT20, WN21, XLH16, LXL10]. Quantitative [BPSS15, DL07, NTL19]. Quantities [Joh07]. Quasi [CNR15, DT18, Pas23]. Quasi-Infinitely [Pas23]. Quasi-likelihood [CNR15]. Quasi-Monte [DT18]. Quickest [BMBV22]. Quintana [GPP16, GL16, RF16].

R [DWM^{+}21]. R-INLA [DWM^{+}21]. R. [Ald08]. Radiation [ZJLC10]. radio [AAFS06]. radiocarbon [BB08a, BALO06]. Random [BS14, BLE16, BB24, CLMtH15, CAS^{+}19, DLPS20, FM18, FH17, KDV09, KK07, Pas23, PHC17, QPC24, SLAV13, SC06, ZL24, BVN09, CKS07, GRM^{+}09]. Randomised [DT18]. Randomization [FMM18]. Randomized [MTS^{+}21, OS24]. Randomly [CRAL23]. Rank [BHvD17, KC23, VGS^{+}21, YMX23, GM09, vdL11b]. Rank-Normalization [VGS^{+}21]. Ranking [BSPD23, CAV23, LLPR06]. Ranks [BSPD23]. rapid [FMV11]. Rare [sC16, GM16]. Rate [WM23]. Rates [CGZ16, DRRS17, LL18, NS23, PPG08, RR12, SY19]. Ratio [SCKL22, SF14, TDC^{+}22, VDP15, KS10b]. Rational [KM14]. Rationale [Bic20]. Ratios [BE13]. Razors [Bic20]. Re [BH07, HHC07]. Re-considering [HHC07]. Re-sampling [BH07]. Reagan [AAFS06]. Real [WC18]. Real-Time [WC18]. Reciprocal [NJM18]. Record [GRM22]. Recursive [POWK24, XJC16]. Recycling [ND20]. Reduced [BHvD17, FMO16, vdL11b]. Reduced-Variance [FMO16]. Reduction [TRKS^{+}17]. refer [Chr06]. Reference [JS24, LCS^{+}14]. Regimes [HMZ^{+}22, LTY21, MM14]. Region [MZMK24, Sha14b]. Regions [ZB18]. Registration [CDH16, EH17]. Regression [ADMO24, APRS22, BPSS15, BBG12, BJS23, BGP15, BWD20, CKG20, CS12, CEMR12, DK15, DM07a, GDB20, Gag23, GKSG21, GLM18, GP12, GKMvCT14, GB13, GB17, GL18, GABP19, GSWF19, GS21, HCPH18, HMC20, HBJ14, HSH21a, HSF20, Hut07, KJP24, KK16, KS19, KCK^{+}21, Kob17, Kow21, KGGC10, Kyu11, LML21, LMCD19, LMPS17, LG24, LG12b, MMW15, MW19, MDO18, Nee19, PMS24, PB20, Pra16a, Qia18, Rah16, RVWG24, RV14, SRA23, SK13, SLAV13, SSML20, SRG13, SR17, TK09, TZG10, TK12b, VL20, WPCAV22, WT20, WN21, XLH16, XX20, ZLV24, ZSM07, ZG19, dCJHdC13, AZ10, AVCGG08, CCQ11, GP10, GB10, HH06, LXL10, RB07, vdL11a, vdL11b]. Regressions [PHC17, Woo14]. Regressive [DBHG19]. Regressors [BDW^{+}24]. regret [GD09]. Regular [GC18]. Regularised [MBB^{+}23]. Regularization [CEMR12, HCPH18, HMC20, KMB19, LCL^{+}14]. Regularized [GP12, GKMvCT14, KS19, SOMD23, LXL10]. regularly [AO06]. Regulatory [NJM18]. Rejection [BF21, SOL^{+}12, MBC24]. Rejoinder [ADGJ^{+}12b, Ber06b, BBS15b, BB08b, BD06b, CFRT06b, CCCG16b, CCL^{+}09b, DM15b, DZP^{+}07b, dSFG15, FD14a, GSW^{+}06b, Gel08b, Hof11a, HKLM10b, JN07a, JMW09b, KSLP12b, LG12a, MCW10a, MT09a, MM13b, PQ16a, PS11b, Pra16b, RS14b, Ryd08b, San12a, SFZ08b, SCHT13a, Scu13b, VGB10b, WC14a, WFR11b, Yin09b, vDK06, Gol06b]. Related [SM19]. Relational [GR20, Hof11b]. Relationship [AE17, CI06, Leo11]. Relationships [JP16]. Relative [AE17, BE13]. relevance [YH11]. Reliability [RSSSSL21]. Repairing [GvO17]. Repartitioning [CFH23]. Representation [AJGM22, FLN^{+}16, PJM^{+}21]. Representations [PK24]. Reproducible [HM23]. reproduction [CO08]. Requiring [TAN^{+}18]. Resolution [FWLH06, Ste15]. Resolve [XTMR17]. Resolved [HYDE21]. Resonance [BHJ18, LJCB14]. respect [DZP^{+}07a]. Response [AFRB14, BBB06, Bra22, Chr24, GS21, HH11, MW19, WC18]. Response-Types [Bra22]. Responses [ADMO24, DCKW08, JNBQ13, LMPS17, MRB12, MDO18, PL16, Hoo08, MT09b]. Resting [CSN^{+}15]. Resting-State [CSN^{+}15]. Restricted [LML21, MHSC16, PB20]. Results [AE17, HK18, HdHG21, KM14]. Return [DG11]. Review [KM14, OS09]. rigorous [DT09]. Risk [BGP15, CLMtH15, DG13, GTGC16, GHO^{+}13, Tre08]. RNA [ZWF^{+}18]. RNA-Seq [ZWF^{+}18]. Robert [Bur10, Gel10, Was10]. Robust [BBG12, CAS^{+}19, FD14b, FCP09, GMdPV21, GMS16, Gu19, LV22, MTS^{+}21, PPR17, VG23, WB18]. Robustness [AE17, Des13, GDB20, Gag23, AO06]. ROC [dCJHdC13]. Role [WCO20]. Ronald [AAFS06]. root [KS10b]. Rotation [SHK07]. Route [DWM^{+}21]. Rubio [Ber14, Sco14, WS14, Xu14]. Rules [DM15a, JGP^{+}19, LVW20]. Rydén [FS08, SK08].

Sample [CCY13, LYL24, MJW08, MKIM24, MTM12, PS15, ZS09, BVN09, HCGS15]. Sampled [RCMO22]. Sampler [FT13, NTL19, SSLD23, SCHT13b]. Samplers [SPD19, ZR21]. Samples [CS13, LG17, SM19]. Sampling [BFPT22, BCR20, BF21, CFH23, FSMWG21, GM16, HHHL18, JLM^{+}17, LR16, MCW10b, NS23, PK24, SN18, SPG15, Ski06, TdVPAB17, WS20, dG15, AZ10, BH07, CF10, MBC24, RW08]. Sancetta [Cla12, Lia12]. Sansó [HG08, Rou08]. Scalable [CL24b, CS12, LL23, MNPM20, RCMO22]. Scalar [ZLV24]. Scale [Des13, FWLH06, HOHS24, Hof16, KK16, PS12, RS14a, TAN^{+}18]. Scale-Dependent [KK16]. Scale-Free [Hof16]. Scaled [PPR17]. Scales [PPR17]. Scaling [Wan15]. scattering [HKLM10a]. Scenes [CCL^{+}09a]. Schedule [LTY21]. Scheme [JS24]. Schemes [LR16]. Schmidl [GM13b, Woo13]. Science [O’H06, vDCE^{+}06, BVN09]. Score [SNMS23, US16, WN21, ZSZ18]. Scoring [DM15a, LVW20]. Scott [Han11, MCG11, SYvD11]. Scutari [Dob13, PS13, Wan13]. Sea [ZC20]. Search [SMBS23, Wan15, BR10, Rob10]. Searching [CSN^{+}15]. Season [QMRM08]. Seemingly [CAD^{+}23, PHC17, AZ10]. Segmentation [DHDC12, GDNJ18]. Segments [BF17, WFR11a]. Selection [BF21, BCLM24, CKY20, CL24b, CS12, CVL12, CMG14, CZGV19, Cas21, CS16b, DM15a, DWM^{+}21, FJM14, FND15, GC18, Gu19, HK22, HM23, Joh13, KCK^{+}21, KC23, KMB19, LLW21, LJCB14, LL20, LMPS17, MCW10b, MRB12, MRG19, PKLM10, PHC17, Qia18, RL14, RM21, RC17, SCKGC21, SKN24, VL20, WOJL22, WM23, YN20, ZHG^{+}16, ZB18, ZG19, Bar11, CHIK08, FS11, LZN08, MPK10, OS09, Sco11]. sell [Lad06]. Semi [BGQ21, HAJF23, QMRM08, HS09]. semi-continuous [HS09]. Semi-Hierarchical [BGQ21]. Semi-Markov [HAJF23]. Semi-parametric [QMRM08]. Semiparametric [BWD20, CAD^{+}23, GL17, HJZ12, HD12, JP16, KC23, MHSC16, MQ22, PS15, PCM19, Pol17, TK12b]. sense [Fie06a]. Sensitivity [GLJB23, MPK10, RH11, RMHR15]. Separable [Hof11b, LM16, LM21]. Separated [Sal18]. Septic [MTS^{+}21]. Seq [ZWF^{+}18]. Sequence [vES21, LN08]. Sequencing [XLY^{+}13]. Sequential [APS18, AFRB14, BW15, BCJ21, CBC23, FT13, KDG21, LLW21, SY19, SPD19, YSH18, BM06]. Sequentially [PSC23]. Series [ADP22, AQ17, BF17, DPM16, FWLH06, HHT24, JNBQ13, KEMM19, LJCB14, NBCC14, NGT19, PFS10, FMV11, FS11]. Sets [KD12, LNR19, MCW10b, DM07b]. Setting [RSST17]. Settings [FMM18, Joh13, LC17]. Several [DvdBWL24, GBGTR19]. Shape [GGPM19, HOHS24, MHSC16, PQ15, AVCGG08]. Shape-Restricted [MHSC16]. Shape-Scale [HOHS24]. Sharpened [Bic20]. Shock [MTS^{+}21]. Should [Lad06]. Shoulders [NS23]. Shrinkage [BCLM24, BDW^{+}24, BR13, CBR23, FM18, GMY21, GTGC16, GC17, GP10, GB17, GR23, HIS22, LBLS22, Ma17, NS23, XCPX22, ZL15, ZB18, Sco11]. Shrinking [Pol17]. Shrunken [NS23]. Signals [BDPW17, CHMK22, vES21]. signed [KS10b]. significance [CZ10, dBPSW08]. similarity [FI09]. Simple [DGMQ13, HW13, TDY18, RB07]. Simulating [BR13]. Simulation [Bha07, CCZ17, GP12, MB12, RSV14, Wei12, KS10b, MS07a, WFR11a]. Simulation-based [GP12, MS07a]. simulation-free [WFR11a]. Simulations [WHG^{+}06]. Simultaneous [GW16, HD12, TK12b, AZ10]. Since [Poi06]. Single [DLPS20, RGC20]. Single-Index [RGC20]. Situation [CEMR12]. Situations [HdHG21]. Size [CCY13, MJW08, MKIM24, MTM12, TSL20, ZS09]. Skew [BBB06, GCM24, HD12]. Skew-Normal [HD12]. Skewed [SLAV13, ZKRVA18, AVCGG08, RB07]. Skin [GHO^{+}13]. Sky [SHK07]. Slab [APD19, RM21, XLH16, ZLV24]. Sliced [JYL17]. Small [ADL12, Pol17, HKLM10a]. Smooth [HSH^{+}21b]. Smoothers [KS19]. Smoothing [CS16a, EH17, FM18, LG14, SK17, VDP15, YZCC16, YSH18, YSLR14].

Social [KM14]. software [O’H06]. Soil [TFHP18]. Solution [CCCG16a, RC17, WCO20]. Some [GD09, GB13, Hoe06, KM14, NB18, YPVG22, RM08]. Somewhere [YPVG22]. Sources [BG13]. Space [DHDC12, DEGP22, LC23, QNK23, SMW19, SC17, SN18, WC14b, XS07, DGS09]. Space-Time [DHDC12, DGS09]. Spaces [LAE^{+}09]. Sparse [BCHJ19, BDPW17, BDW^{+}24, GC17, GB13, GABP19, HOHS24, MW15, OK22, SSML20, XCPX22, XJC16, vES21]. Sparsity [GTGC16, OJP23]. Spatial [BJM^{+}22, BHJ18, CRAL23, DBHG19, FSG08, HJZ12, HGXS23, JMKW09, JLM^{+}17, KJP24, KK22, LJCB14, LG17, LM16, LM21, LG12b, MZMK24, OMC19, PQ16b, PBT^{+}21, SLB^{+}21, TFHP18, ZKRVA18, ZLV24, ZKS23, CG10, MPK10]. Spatially [LBBJ16, NGT19, ZSM07]. Spatially-adjusted [ZSM07]. Spatially-Correlated [NGT19]. Spatio [BHW18, RdGvP06, VDF^{+}12, WSDC13, ZC20]. Spatio-Temporal [BHW18, RdGvP06, VDF^{+}12, WSDC13, ZC20]. Spatiotemporal [BHJ18, SC06]. Species [BCR20, JLM^{+}17, TRKS^{+}17, VHV20, ZS09, BB10, GSW^{+}06a, Hoe06]. Specific [NPKC14, PQ15, RRJW20, MBB^{+}23]. Specification [SNMS23, Wil18, AM07]. Spectral [BGQ20, TFHP18]. Spectroscopy [HYDE21]. sphere [Sco11]. Spike [APD19, RM21, XLH16, ZLV24]. Spike-and-Slab [RM21, ZLV24]. Spiked [CZGV19, KDV09, XCPX22]. Spline [CS16a, FJS08, Pac06]. Splines [BS14, Kyu11, SK17, YSLR14, MBB^{+}23]. Splitting [GPL^{+}19, JN07b]. Spread [SHMM23, VDF^{+}12]. Squared [NJ21]. stable [PKL^{+}11]. Stacking [LC17, YVSG18, YPVG22]. Stage [DD07, LLPR06, OS24, SY17]. staged [FS11]. Standard [KGGC10, RC17]. State [CSN^{+}15, CAD^{+}23, DEGP22, QNK23, SN18, WC14b]. State-Space [WC14b]. States [SN18, OGPD19]. Stationary [RCLW17]. Statistical [DMF16, Gin07, GGPM19, WG15, CZ10]. Statistics [CNR15, FMM18, LML21, Poi06, Cla10, Gel08a, Gol06b, Lad06]. Status [PMS24]. Steel [Ber14, Sco14, WS14, Xu14]. Stein [Wen10]. Step [GABP19, HSH^{+}21b, IW19, POWK24]. Steps [ND20]. Stick [BJP12, FLN^{+}16, GLJB23, HZ22, SM19, RD11]. Stick-Breaking [BJP12, FLN^{+}16, GLJB23, HZ22, RD11]. Stiefel [PSMB20, PJM^{+}21]. Stochastic [AKO19, APL24, ADP19, ANRSL16, AKO23, BKD21, DG11, Gu19, HLC20, PL16, SK13, Sha21, TN14, VGE19, Wan15, YSLR14, ZL24, BR10, CO08, DGS09, MT09b]. Stop [Chr06]. Stopping [HdHG21]. Strategies [CHG12]. Strategy [MRB12, LW09]. Stratified [HJZ12, NPKC14]. Structural [DCKW08, PS20, JP08]. Structure [CLMtH15, EDF^{+}19, GLC24, MMN22, MW15, Wan15, YS07, ZM23]. Structured [GKSG21, KK16, KCK^{+}21]. Structures [PNNC17, Vir11]. Student [HSH21a]. Student- [HSH21a]. Studies [CS12, GS21]. Study [DL07, DD07, MNS^{+}20]. Subject [PQ15, CG10]. Subject-Specific [PQ15]. Subjective [Gol06a, WG15, Chr06]. Subjectivity [Gol06b, Dra06, O’H06].

Submodel [LBLS22]. Subnational [BG13]. Subposteriors [NS18]. Subspace [SKN24, TZG10, Hof06]. Substructures [vdBDB23]. Sufficiency [Woo14]. Sufficient [SKG15]. Suggestion [RMP12]. Summaries [RCMO22]. Summary [CNR15]. Summation [Qia18]. Sums [Kad16]. Sun [DL15, MGP15, Rou15, Siv15]. Superiority [EMS13]. Supplemental [Ano11a, Ano12a, Ano13a]. Supplementary [Ano14b, Ano14c]. supplementation [DZP^{+}07a]. Support [BJQ12, PS11a, PS11b]. Sure [AZ13]. Surfaces [BJM^{+}22, Sco11]. Surgery [XTMR17]. surgical [LBBJ16]. Surrogate [JGVM21, FMV11]. Survival [LBLS22, MHSC16, PMS24, DZP^{+}07a]. Switching [SY17, TK09, HS09]. Symmetric [BPH21, NB18, WC14b]. Symmetry [WCO20]. Symptom [LMC20]. Synthetic [HRW18, PSC23]. System [SFZ08a, SHG^{+}10]. Systems [SS08, SCHT13b].

t [VW14]. Table [Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13b, Ano13c, Ano13d, Ano13e, Ano14d, Ano14e, Ano16c, Ano16d, Ano23c, Ano23d, Ano24c, Ano24d]. Tadesse [CGM09, Fra09, Li09, Ste09]. Tail [ALC24, BGP15, RSV14]. tailed [GOO07, Tre08]. Target [Kom15]. Targeted [MCW10b]. Task [SMBS23]. technique [RM08]. Techniques [TAN^{+}18, YS07, AZ10]. Telemetric [HAJF23]. Telescoping [FSMWG21]. Temperature [MNPM20]. Temporal [BHW18, RdGvP06, VDF^{+}12, WSDC13, ZC20]. Temporally [HJZ12]. Temporally-Stratified [HJZ12]. Tensor [GS21]. Term [CLMtH15]. Test [FMM18, LL20, AM07, dBPSW08]. Testing [CB21, DvdBWL24, DD18, FH17, GTGC16, GBGTR19, HCGS15, KDV09, MP18, MF19, Sal18, Spi08, Spi11]. Tests [FMM18, JYL17, LYL24, NJ21, SY17]. Theorem [PS15, SS08, Ald08]. Theorems [HZ22]. Theoretic [OM22, XTMR17]. Theoretical [BG21]. Theory [RSSSSL21, pD20, Cla10, Hoo08]. three [Vir11]. three-way [Vir11]. Threshold [GKMvCT14]. Thresholded [CKY20]. Time [ADP22, AQ17, BF17, DHDC12, DPM16, DD07, FWLH06, HHT24, HK18, HMZ^{+}22, JNBQ13, KR21, KEMM19, Kow21, LJCB14, MBB^{+}23, MHSC16, NBCC14, NGT19, PFS10, SS08, SW22, Sha21, WC18, YHW16, ZWC^{+}16, DGS09, FMV11, FS11, HS09]. Time-Dependent [DD07, MHSC16]. Time-Frequency [YHW16]. Time-Indexed [SW22]. Time-Ordered [Kow21]. Time-Series [LJCB14]. Time-trends [MBB^{+}23]. Time-Varying [KR21, YHW16]. Time-Weighted [HK18]. Times [RRJW20]. Timing [HMZ^{+}22]. Tobit [Kob17]. Tolerance [SCKGC21]. Topic [GTHB19]. Topological [MMN22, Mar24]. Toxicities [LN13]. Tractable [WC14b]. Training [CS13]. Traits [BPSS15, OBS13]. Transformation [Bra22, GLC24]. Transmission [MNS^{+}20]. Treat [MTS^{+}21]. Treatment [HCPH18, HMZ^{+}22, OS24, SM17, VDP15]. Treatments [GBGTR19, XTMR17]. Tree [HMC20, LV22, LBLS22, Ma17, MKIM24, OBS13, Pra16a, ZSM07]. Trees [PK24, FS11]. Trends [TGK^{+}11, MBB^{+}23]. Trial [CBC23, LTY21, MTS^{+}21]. Trials [FCP09, HSC12, LN13, SY17]. Tropical [TGK^{+}11]. true [BP08]. Truncated [CCZ17, HK18]. Trustworthy [MKIM24]. Tucker [Hof11b, Hof16]. Tumor [ZJLC10]. Tuning [BCJ21, RC17]. Tutorial [WSD22]. Two [GHM^{+}23, HCGS15, HSH^{+}21b, LYL24, LLPR06, MRB12, OS24, RS14a, SY17, ZSM07, ZHG^{+}16]. Two-Component [ZHG^{+}16]. Two-Group [GHM^{+}23]. Two-Piece [RS14a]. Two-Pronged [MRB12]. Two-Sample [LYL24, HCGS15]. Two-Stage [LLPR06, OS24, SY17]. Two-Step [HSH^{+}21b]. Type [AJGM22, CCO24, Ma17, PB20, PL16, HR20, SY19]. Types [Bra22].

U.S. [HGXS23]. Ultra [BDPW17]. Ultra-Sparse [BDPW17]. Ultrahigh [Joh13]. Ultrahigh-dimensional [Joh13]. Un-Separated [Sal18]. Unattenuated [WS20]. Uncertain [BSPD23]. Uncertainty [CHG12, CCDT^{+}22, CCCG16a, LV22, VGB10a, YMX23, vdPSvdV17]. Unconstrained [LL18]. Underreported [dOAL^{+}22]. Undirected [ZHA23]. Unified [TSL20]. Uniformly [NJ21]. Unit [JNBQ13]. United [OGPD19]. Univariate [QSF09]. Universality [San12b]. Unknown [MB12]. Unlabeled [KD12]. unobserved [JMKW09]. Unrelated [CAD^{+}23, PHC17, AZ10]. Unseen [ZS09]. Unsettled [CGS22]. Update [TSA20]. upper [MM07]. Use [BR13, GLM18, GSWF19]. Used [Scu13a]. Useful [YPVG22]. Using [APD19, APRS22, BGP15, BHJ18, BWD20, Bra22, BG13, CSN^{+}15, DBHG19, DWM^{+}21, DT18, FD14b, Gop22, GL18, HOHS24, Han06, HSC12, HM23, Joh07, KD12, LV22, LG17, LMC20, LBBJ16, LG12b, MRB12, NSAL^{+}21, OM20, Paj17, PVC20, RS13, RDP16, SFZ08a, SN07, SHK07, SG16, TRKS^{+}17, VDP15, VDF^{+}12, WWACH16, Wil18, WN21, XX20, YVSG18, YSLR14, ZKRVA18, ZL24, ZWF^{+}18, ZG19, AZ10, AO06, BF21, BC11a, Chr06, DEGP22, DGS09, GD09, HKLM10a, KS10a, MM07, RB07, SB11, dTM10]. Utilising [JV23]. Utility [CBC23, LAE^{+}09]. Utility-Based [CBC23, LAE^{+}09].

Validation [BH07, HC17]. valuation [LKF09]. value [SF14]. Valued [BHW18, DPM16]. Values [FMM18]. VAR [PKLM10]. Variable [CL24b, CS12, Cas21, FND15, Gu19, LLW21, LJCB14, LMPS17, MRG19, PHC17, Qia18, RL14, RM21, SKN24, VL20, WOJL22, ZHG^{+}16, ZB18, ZG19, Bar11, CHIK08, HH06, OS09, YH11, vdL11a]. Variables [BBGR21, BG06, FH17, GR20, Kad16, Kom15, vdL11b]. Variance [FMO16, FHK^{+}20, KK16, MRG19, PMG14, SOMD23, Wan17, Gel06, HHC07, KN06]. Variances [DvdBWL24, Pol17]. Variate [CW07, ZWDJ14]. Variates [SOMD23]. Variational [BG06, BJ06, CS12, EH17, Gop22, MW19, QNK23, SK13, TN14, TSA20, WOPF11, ZSZ18, WT06]. vary [DZP^{+}07a]. Varying [KR21, YHW16, AO06]. Vascular [ZJLC10]. Vector [PKL^{+}11, WSD22, PS11a, PS11b]. Velocity [CHG12]. Verbal [LMC20]. Vernon [LH10, Poo10, Ran10, vD10]. Versions [HRW18]. versus [Ryd08a]. Vertical [SHG^{+}10]. Via [BGQ20, PKL^{+}11, CKY20, Chr24, GMdPV21, GGPM19, HC17, Hof11b, JYL17, KS10b, KDG21, LCL^{+}14, MQ22, PQ15, PJM^{+}21, RCMO22, RSV14, Sco11, WRC11, ZSZ18, ZB18]. viewpoint [Spi08]. Vine [GC18, SCHT13b, Wil18]. Vine-copula [SCHT13b]. violence [JP08]. Viot [ALR21]. Viot-driven [ALR21]. Volatility [DG11, VGE19, ZD17]. Votes [WSDC13]. vs [FMM18, SMBS23].

walk [CF10]. Way [LC22, BVN09, CKS07, Vir11]. Weak [CHMK22]. Weight [GHO^{+}13, DZP^{+}07a]. Weight-of-Evidence [GHO^{+}13]. Weighted [CRAL23, HK18, LG24, SPG15]. Weighting [SHMM23]. Weights [BW15, CVCB23]. Where [DWM^{+}21]. which [OS09]. Whittle [KEMM19]. Who [AAFS06]. Whole [Ano06e, Ano06f, Ano06g, Ano06h, Ano07e, Ano07f, Ano07g, Ano07h, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10e, Ano10f, Ano10g, Ano10h, Ano11f, Ano11g, Ano11h, Ano11i, Ano12f, Ano12g, Ano12h, Ano12i, Ano13f, Ano13g, Ano13h, Ano13i, Ano14f, Ano14g, SLB^{+}21]. Whole-Brain [SLB^{+}21]. Windle [Cas14, For14, tHM14]. wise [Kad06]. Within [BCJ21, SHG^{+}10, LG24, WFR11a]. Without [PHG23, DLPS20]. Wombling [MC07]. Women [WPCAV22]. wrote [AAFS06]. Wyse [Fea11, Koo11].

Yin [CK09, Cra09]. Ylvisaker [JB18]. Yor [ADP19, BFPT22, Scr14].

Zag [SSLD23]. Zero [Nee19, PMG14, SOMD23]. Zero-Inflated [Nee19]. Zero-Variance [SOMD23]. Zig [SSLD23]. Zig-Zag [SSLD23].

References

Airoldi:2006:WWR

[AAFS06] Edoardo M. Airoldi, Annelise G. Anderson, Stephen E. Fienberg, and Kiron K. Skinner. Who wrote Ronald Reagan’s radio addresses? Bayesian Analysis, 1(2):289–319, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

            tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
            issue02/airoldi289-320.pdf;   http://projecteuclid.org/
            euclid.ba/1340371064.

Albert:2009:CAJ

[AB09] Jim Albert and Phil Birnbaum. Comment on article by Jensen et al. Bayesian Analysis, 4(4):653–660, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2009/
vol04/issue04/albert.pdf;   http://projecteuclid.org/
euclid.ba/1340369816. See [JMW09a].

Albert:2012:CEO

[ADGJ^{+}12a] Isabelle Albert, Sophie Donnet, Chantal Guihenneuc-Jouyaux, Samantha Low-Choy, Kerrie Mengersen, and Judith Rousseau. Combining expert opinions in prior elicitation. Bayesian Analysis, 7(3):503–532, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue03/albert.pdf;
http://projecteuclid.org/euclid.ba/1346158771. See com-

ments [Gos12, Fre12] and rejoinder [ADGJ^{+}12b].

Albert:2012:R

[ADGJ^{+}12b] Isabelle Albert, Sophie Donnet, Chantal Guihenneuc-Jouyaux, Samantha Low-Choy, Kerrie Mengersen, and Judith Rousseau. Rejoinder. Bayesian Analysis, 7(3):541–546, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue03/albert_rejoinder.pdf;   http://projecteuclid.
org/euclid.ba/1346158774. See [ADGJ^{+}12a].

Arima:2012:OBA

[ADL12] Serena Arima, Gauri S. Datta, and Brunero Liseo. Objective Bayesian analysis of a measurement error small area model. Bayesian Analysis, 7(2):363–384, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2012/vol07/issue02/arima.pdf;
 http://projecteuclid.org/euclid.ba/1339878892.

Ascari:2024:MMR

[ADMO24] Roberto Ascari, Agnese Maria Di Brisco, Sonia Migliorati, and Andrea Ongaro. A multivariate mixture regression model for constrained responses. Bayesian Analysis, 19(2):377–405, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-2/A-Multivariate-Mixture-Regression-Model-for-
Constrained-Responses/10.1214/22-BA1359.full.

Arbel:2019:SAP

[ADP19] Julyan Arbel, Pierpaolo De Blasi, and Igor Prünster. Stochastic approximations to the Pitman–Yor process. Bayesian Analysis, 14(4):1201–1219, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1560240026.

Alexopoulos:2022:BPJ

[ADP22] Angelos Alexopoulos, Petros Dellaportas, and Omiros Papaspiliopoulos. Bayesian prediction of jumps in large panels of time series data. Bayesian Analysis, 17(2):651–683, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-2/Bayesian-Prediction-
of-Jumps-in-Large-Panels-of-Time-Series/10.1214/21-BA1268.
full.

AlLabadi:2017:ORR

[AE17] Luai Al Labadi and Michael Evans. Optimal robustness results for relative belief inferences and the relationship to prior-data conflict. Bayesian Analysis, 12(3):705–728, September 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1473276256.

Azadi:2014:BSE

[AFRB14] Nammam A. Azadi, Paul Fearnhead, Gareth Ridall, and Joleen H. Blok. Bayesian sequential experimental design for binary response data with application to electromyographic experiments. Bayesian Analysis, 9(2):287–306, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue02/azadi.pdf;   http://projecteuclid.org/euclid.
ba/1401148310.

Alexanderian:2016:BOE

[AGG16] Alen Alexanderian, Philip J. Gloor, and Omar Ghattas. On Bayesian A- and D-optimal experimental designs in infinite dimensions. Bayesian Analysis, 11(3):671–695, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1440594948.

Ayala:2022:DPM

[AJGM22] Daniel Ayala, Leonardo Jofré, Luis Gutiérrez, and Ramsés H. Mena. On a Dirichlet process mixture representation of phase-type distributions. Bayesian Analysis, 17(3):765– 790, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-17/issue-3/On-
a-Dirichlet-Process-Mixture-Representation-of-Phase-
Type-Distributions/10.1214/21-BA1272.full.

Alharthi:2019:BFP

[AKO19] Muteb Alharthi, Theodore Kypraios, and Philip D. O’Neill. Bayes factors for partially observed stochastic epidemic models. Bayesian Analysis, 14(3):907–936, September 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1560240033.

Aristotelous:2023:PPC

[AKO23] Georgios Aristotelous, Theodore Kypraios, and Philip D. O’Neill. Posterior predictive checking for partially observed stochastic epidemic models. Bayesian Analysis, 18 (4):1283–1310, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Posterior-Predictive-Checking-for-Partially-
Observed-Stochastic-Epidemic-Models/10.1214/22-BA1336.
full.

Ayed:2024:NGG

[ALC24] Fadhel Ayed, Juho Lee, and François Caron. The normalgeneralised gamma-Pareto process: A novel pure-jump Lévy process with flexible tail and jump-activity properties. Bayesian Analysis, 19(1):123–152, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/The-Normal-Generalised-Gamma-Pareto-Process-
-A-Novel-Pure/10.1214/22-BA1343.full.

Aldrich:2008:RFB

[Ald08] John Aldrich. R. A. Fisher on Bayes and Bayes’ theorem. Bayesian Analysis, 3(1):161–170, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2008/vol03/issue01/aldrich.
pdf;   http://projecteuclid.org/euclid.ba/1340370565.

Allen:2011:CAH

[All11] Genevera I. Allen. Comment on article by Hoff. Bayesian Analysis, 6(2):197–201, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue02/allen.pdf;   http:
//projecteuclid.org/euclid.ba/1339612041. See [Hof11b].

Ascolani:2021:PIF

[ALR21] Filippo Ascolani, Antonio Lijoi, and Matteo Ruggiero. Predictive inference with Fleming–Viot-driven dependent Dirichlet processes. Bayesian Analysis, 16(2):371–395, June 2021. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-2/Predictive-inference-with-FlemingViot-
driven-dependent-Dirichlet-processes/10.1214/20-BA1206.
full.

Almeida:2007:BES

[AM07] Carlos Almeida and Michel Mouchart. Bayesian encompassing specification test under not completely known partial observability. Bayesian Analysis, 2(2):303–318, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2007/vol02/issue02/almeida.
pdf;   http://projecteuclid.org/euclid.ba/1340370802.

Anonymous:2006:TCa

[Ano06a] Anonymous. Table of contents. Bayesian Analysis, 1(1):??, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue01/toc.pdf.

Anonymous:2006:TCb

[Ano06b] Anonymous. Table of contents. Bayesian Analysis, 1(2):??, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue02/toc.pdf.

Anonymous:2006:TCc

[Ano06c] Anonymous. Table of contents. Bayesian Analysis, 1(3):??, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2006/vol01/issue03/toc.pdf.

Anonymous:2006:TCd

[Ano06d] Anonymous. Table of contents. Bayesian Analysis, 1(4):??, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue04/toc.pdf.

Anonymous:2006:WIa

[Ano06e] Anonymous. Whole issue. Bayesian Analysis, 1(1):??, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue01/issue01.pdf.

Anonymous:2006:WIb

[Ano06f] Anonymous. Whole issue. Bayesian Analysis, 1(2):??, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue02/issue02.pdf.

Anonymous:2006:WIc

[Ano06g] Anonymous. Whole issue. Bayesian Analysis, 1(3):??, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue03/issue03.pdf.

Anonymous:2006:WId

[Ano06h] Anonymous. Whole issue. Bayesian Analysis, 1(4):??, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/issue04.pdf.

Anonymous:2007:TCa

[Ano07a] Anonymous. Table of contents. Bayesian Analysis, 2(1):??, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2007/
vol02/issue01/toc.pdf.

Anonymous:2007:TCb

[Ano07b] Anonymous. Table of contents. Bayesian Analysis, 2(2):??, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue02/toc.pdf.

Anonymous:2007:TCc

[Ano07c] Anonymous. Table of contents. Bayesian Analysis, 2(3):??, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2007/vol02/issue03/toc.pdf.

Anonymous:2007:TCd

[Ano07d] Anonymous. Table of contents. Bayesian Analysis, 2(4):??, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2007/
vol02/issue04/toc.pdf.

Anonymous:2007:WIa

[Ano07e] Anonymous. Whole issue. Bayesian Analysis, 2(1):??, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue01/issue01.pdf.

Anonymous:2007:WIb

[Ano07f] Anonymous. Whole issue. Bayesian Analysis, 2(2):??, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue02/issue02.pdf.

Anonymous:2007:WIc

[Ano07g] Anonymous. Whole issue. Bayesian Analysis, 2(3):??, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue03/issue03.pdf.

Anonymous:2007:WId

[Ano07h] Anonymous. Whole issue. Bayesian Analysis, 2(4):??, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue04/issue04.pdf.

Anonymous:2008:TCa

[Ano08a] Anonymous. Table of contents. Bayesian Analysis, 3(1):??, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue01/toc.pdf.

Anonymous:2008:TCb

[Ano08b] Anonymous. Table of contents. Bayesian Analysis, 3(2):??, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue02/toc.pdf.

Anonymous:2008:TCc

[Ano08c] Anonymous. Table of contents. Bayesian Analysis, 3(3):??, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2008/vol03/issue03/toc.pdf.

Anonymous:2008:TCd

[Ano08d] Anonymous. Table of contents. Bayesian Analysis, 3(4):??, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue04/toc.pdf.

Anonymous:2008:WIa

[Ano08e] Anonymous. Whole issue. Bayesian Analysis, 3(1):??, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue01/issue01.pdf.

Anonymous:2008:WIb

[Ano08f] Anonymous. Whole issue. Bayesian Analysis, 3(2):??, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue02/issue02.pdf.

Anonymous:2008:WIc

[Ano08g] Anonymous. Whole issue. Bayesian Analysis, 3(3):??, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue03/issue03.pdf.

Anonymous:2008:WId

[Ano08h] Anonymous. Whole issue. Bayesian Analysis, 3(4):??, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue04/issue04.pdf.

Anonymous:2009:TCa

[Ano09a] Anonymous. Table of contents. Bayesian Analysis, 4(1):??, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2009/
vol04/issue01/toc.pdf.

Anonymous:2009:TCb

[Ano09b] Anonymous. Table of contents. Bayesian Analysis, 4(2):??, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue02/toc.pdf.

Anonymous:2009:TCc

[Ano09c] Anonymous. Table of contents. Bayesian Analysis, 4(3):??, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2009/vol04/issue03/toc.pdf.

Anonymous:2009:TCd

[Ano09d] Anonymous. Table of contents. Bayesian Analysis, 4(4):??, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2009/
vol04/issue04/toc.pdf.

Anonymous:2009:WIa

[Ano09e] Anonymous. Whole issue. Bayesian Analysis, 4(1):??, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue01/issue01.pdf.

Anonymous:2009:WIb

[Ano09f] Anonymous. Whole issue. Bayesian Analysis, 4(2):??, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue02/issue02.pdf.

Anonymous:2009:WIc

[Ano09g] Anonymous. Whole issue. Bayesian Analysis, 4(3):??, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue03/issue03.pdf.

Anonymous:2009:WId

[Ano09h] Anonymous. Whole issue. Bayesian Analysis, 4(4):??, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue04/issue04.pdf.

Anonymous:2010:TCa

[Ano10a] Anonymous. Table of contents. Bayesian Analysis, 5(1):??, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2010/
vol05/issue01/toc.pdf.

Anonymous:2010:TCb

[Ano10b] Anonymous. Table of contents. Bayesian Analysis, 5(2):??, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue02/toc.pdf.

Anonymous:2010:TCc

[Ano10c] Anonymous. Table of contents. Bayesian Analysis, 5(3):??, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2010/vol05/issue03/toc.pdf.

Anonymous:2010:TCd

[Ano10d] Anonymous. Table of contents. Bayesian Analysis, 5(4):??, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2010/
vol05/issue04/toc.pdf.

Anonymous:2010:WIa

[Ano10e] Anonymous. Whole issue. Bayesian Analysis, 5(1):??, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue01/issue01.pdf.

Anonymous:2010:WIb

[Ano10f] Anonymous. Whole issue. Bayesian Analysis, 5(2):??, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue02/issue02.pdf.

Anonymous:2010:WIc

[Ano10g] Anonymous. Whole issue. Bayesian Analysis, 5(3):??, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue03/issue03.pdf.

Anonymous:2010:WId

[Ano10h] Anonymous. Whole issue. Bayesian Analysis, 5(4):??, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue04/issue04.pdf.

Anonymous:2011:SF

[Ano11a] Anonymous. Supplemental file. Bayesian Analysis, 6(4):??, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue04/wand/GEVnorMixParms.txt.

Anonymous:2011:TCa

[Ano11b] Anonymous. Table of contents. Bayesian Analysis, 6(1):??, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue01/toc.pdf.

Anonymous:2011:TCb

[Ano11c] Anonymous. Table of contents. Bayesian Analysis, 6(2):??, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue02/toc.pdf.

Anonymous:2011:TCc

[Ano11d] Anonymous. Table of contents. Bayesian Analysis, 6(3):??, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2011/vol06/issue03/toc.pdf.

Anonymous:2011:TCd

[Ano11e] Anonymous. Table of contents. Bayesian Analysis, 6(4):??, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue04/toc.pdf.

Anonymous:2011:WIa

[Ano11f] Anonymous. Whole issue. Bayesian Analysis, 6(1):??, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue01/issue01.pdf.

Anonymous:2011:WIb

[Ano11g] Anonymous. Whole issue. Bayesian Analysis, 6(2):??, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue02/issue02.pdf.

Anonymous:2011:WIc

[Ano11h] Anonymous. Whole issue. Bayesian Analysis, 6(3):??, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue03/issue03.pdf.

Anonymous:2011:WId

[Ano11i] Anonymous. Whole issue. Bayesian Analysis, 6(4):??, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue04/issue04.pdf.

Anonymous:2012:SF

[Ano12a] Anonymous. Supplemental file. Bayesian Analysis, 7(3):??, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2012/vol07/issue03/morita_supp.pdf.

Anonymous:2012:TCa

[Ano12b] Anonymous. Table of contents. Bayesian Analysis, 7(1):??, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2012/
vol07/issue01/toc.pdf.

Anonymous:2012:TCb

[Ano12c] Anonymous. Table of contents. Bayesian Analysis, 7(2):??, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue02/toc.pdf.

Anonymous:2012:TCc

[Ano12d] Anonymous. Table of contents. Bayesian Analysis, 7(3):??, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2012/vol07/issue03/toc.pdf.

Anonymous:2012:TCd

[Ano12e] Anonymous. Table of contents. Bayesian Analysis, 7(4):??, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2012/
vol07/issue04/toc.pdf.

Anonymous:2012:WIa

[Ano12f] Anonymous. Whole issue. Bayesian Analysis, 7(1):??, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue01/issue01.pdf.

Anonymous:2012:WIb

[Ano12g] Anonymous. Whole issue. Bayesian Analysis, 7(2):??, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue02/issue02.pdf.

Anonymous:2012:WIc

[Ano12h] Anonymous. Whole issue. Bayesian Analysis, 7(3):??, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue03/issue03.pdf.

Anonymous:2012:WId

[Ano12i] Anonymous. Whole issue. Bayesian Analysis, 7(4):??, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue04/issue04.pdf.

Anonymous:2013:SF

[Ano13a] Anonymous. Supplemental file. Bayesian Analysis, 8(1):??, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue01/ogle_supplement.pdf.

Anonymous:2013:TCa

[Ano13b] Anonymous. Table of contents. Bayesian Analysis, 8(1):??, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue01/toc.pdf.

Anonymous:2013:TCb

[Ano13c] Anonymous. Table of contents. Bayesian Analysis, 8(2):??, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue02/toc.pdf.

Anonymous:2013:TCc

[Ano13d] Anonymous. Table of contents. Bayesian Analysis, 8(3):??, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2013/vol08/issue03/toc.pdf.

Anonymous:2013:TCd

[Ano13e] Anonymous. Table of contents. Bayesian Analysis, 8(4):??, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue04/toc.pdf.

Anonymous:2013:WIa

[Ano13f] Anonymous. Whole issue. Bayesian Analysis, 8(1):??, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/issue01.pdf.

Anonymous:2013:WIb

[Ano13g] Anonymous. Whole issue. Bayesian Analysis, 8(2):??, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue02/issue02.pdf.

Anonymous:2013:WIc

[Ano13h] Anonymous. Whole issue. Bayesian Analysis, 8(3):??, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue03/issue03.pdf.

Anonymous:2013:WId

[Ano13i] Anonymous. Whole issue. Bayesian Analysis, 8(4):??, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue04/issue04.pdf.

Anonymous:2014:CDA

[Ano14a] Anonymous. Contributed discussion on article by Finegold and Drton. Bayesian Analysis, 9(3):561–590, September 2014. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1409921105. See

[FD14b].

Anonymous:2014:SMa

[Ano14b] Anonymous. Supplementary material. Bayesian Analysis, 9(1): ??, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2014/vol09/issue01/rubio_supplement.pdf.

Anonymous:2014:SMb

[Ano14c] Anonymous. Supplementary material. Bayesian Analysis, 9(2): ??, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2014/
vol09/issue02/liu_supplement.pdf.

Anonymous:2014:TCa

[Ano14d] Anonymous. Table of contents. Bayesian Analysis, 9(1):??, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2014/
vol09/issue01/toc.pdf.

Anonymous:2014:TCb

[Ano14e] Anonymous. Table of contents. Bayesian Analysis, 9(2):??, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue02/toc.pdf.

Anonymous:2014:WIa

[Ano14f] Anonymous. Whole issue. Bayesian Analysis, 9(1):??, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue01/issue01.pdf.

Anonymous:2014:WIb

[Ano14g] Anonymous. Whole issue. Bayesian Analysis, 9(2):??, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue02/issue02.pdf.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial board. Bayesian Analysis, 11(3):??, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1472829059.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial board. Bayesian Analysis, 11(4):??, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1480474947.

Anonymous:2016:TCa

[Ano16c] Anonymous. Table of contents. Bayesian Analysis, 11(3):??, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1472829058.

Anonymous:2016:TCb

[Ano16d] Anonymous. Table of contents. Bayesian Analysis, 11(4):??, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1480474946.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Bayesian Analysis, 18(3):??, September 2023. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-18/issue-3/Editorial-
Board/ba_18_3p1.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Bayesian Analysis, 18(4):??, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-4/Editorial-Board/
ba_18_4p1.full.

Anonymous:2023:TCa

[Ano23c] Anonymous. Table of contents. Bayesian Analysis, 18(3): ??, September 2023. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-18/issue-3/Table-of-
Contents/ba_18_3co.full.

Anonymous:2023:TCb

[Ano23d] Anonymous. Table of contents. Bayesian Analysis, 18(4):??, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-4/Table-of-Contents/
ba_18_4co.full.

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial board. Bayesian Analysis, 19(1):??, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-1/Editorial-Board/
ba_19_1p1.full.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial board. Bayesian Analysis, 19(2):??, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-2/Editorial-Board/
ba_19_2p1.full.

Anonymous:2024:TCa

[Ano24c] Anonymous. Table of contents. Bayesian Analysis, 19(1):??, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-1/Table-of-Contents/
ba_19_1co.full.

Anonymous:2024:TCb

[Ano24d] Anonymous. Table of contents. Bayesian Analysis, 19(2):??, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-2/Table-of-Contents/
ba_19_2co.full.

Arias-Nicolas:2016:NCP

[ANRSL16] J. Pablo Arias-Nicolás, Fabrizio Ruggeri, and Alfonso Suárez-Llorens. New classes of priors based on stochastic orders and distortion functions. Bayesian Analysis, 11(4):1107–1136, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1448590531.

Andrade:2006:BRM

[AO06] J. A. A. Andrade and A. O’Hagan. Bayesian robustness modeling using regularly varying distributions. Bayesian Analysis, 1(1):169–188, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue01/andrade.pdf;   http://
projecteuclid.org/euclid.ba/1340371079.

Aitken:2013:CDA

[APA^{+}13] Murray Aitken, Julia Polak, Julyan Arbel, Bernardo Nipoti, Bertrand S. Clarke, Gregory E. Holt, Andrew Gelman, Miroslav Kárný, Michalis Kolossiatis, Athanasios Kottas, Maria DeYoreo, Valerie Poynor, Susan M. Paddock, Terrance D. Savitsky, G. Parmigiani, L. Trippa, François Perron, Christian P. Robert, Judith Rousseau, James G. Scott, and Surya T. Tokdar. Contributed discussion on article by Müller and Mitra. Bayesian Analysis, 8(2):323–356, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue02/mueller_comments.
pdf;   http://projecteuclid.org/euclid.ba/1369407554.

See [MM13a].

Antonelli:2019:HDC

[APD19] Joseph Antonelli, Giovanni Parmigiani, and Francesca Dominici. High-dimensional confounding adjustment using continuous spike and slab priors. Bayesian Analysis, 14(3):805–828, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560240029.

Amini:2024:HSB

[APL24] Arash Amini, Marina Paez, and Lizhen Lin. Hierarchical stochastic block model for community detection in multiplex networks. Bayesian Analysis, 19(1):319–345, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/Hierarchical-Stochastic-Block-Model-for-Community-
Detection-in-Multiplex-Networks/10.1214/22-BA1355.full.

Avalos-Pacheco:2022:HLD

[APRS22] Alejandra Avalos-Pacheco, David Rossell, and Richard S. Savage. Heterogeneous large datasets integration using Bayesian factor regression. Bayesian Analysis, 17(1):33–66, ???? 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-1/Heterogeneous-Large-Datasets-Integration-
Using-Bayesian-Factor-Regression/10.1214/20-BA1240.full.

Aktekin:2018:SBA

[APS18] Tevfik Aktekin, Nick Polson, and Refik Soyer. Sequential Bayesian analysis of multivariate count data. Bayesian Analysis, 13(2):385–409, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1490234588.

Anacleto:2017:DCG

[AQ17] Osvaldo Anacleto and Catriona Queen. Dynamic chain graph models for time series network data. Bayesian Analysis, 12(2): 491–509, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1466165926.

Al-Taie:2023:BLB

[ATF23] Wael A. J. Al-Taie and Malcolm Farrow. Bayes linear Bayes networks with an application to prognostic indices. Bayesian Analysis, 18(2):437–463, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Bayes-Linear-Bayes-Networks-with-an-Application-
to-Prognostic-Indices/10.1214/22-BA1314.full.

Arellano-Valle:2008:BIS

[AVCGG08] Reinaldo B. Arellano-Valle, Luis M. Castro, Marc G. Genton, and Héctor W. Gómez. Bayesian inference for shape mixtures of skewed distributions, with application to regression analysis. Bayesian Analysis, 3(3):513–539, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2008/vol03/issue03/arellano.
pdf;   http://projecteuclid.org/euclid.ba/1340370436.

Ando:2010:HBA

[AZ10] Tomohiro Ando and Arnold Zellner. Hierarchical Bayesian analysis of the seemingly unrelated regression and simultaneous equations models using a combination of direct Monte Carlo and importance sampling techniques. Bayesian Analysis, 5(1):65–95, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue01/ando.pdf;   http:/
/projecteuclid.org/euclid.ba/1340369793.

AlLabadi:2013:APA

[AZ13] Luai Al Labadi and Mahmoud Zarepour. On asymptotic properties and almost sure approximation of the normalized inverse-Gaussian process. Bayesian Analysis, 8(3):553– 568, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/zarepour.pdf;   http:/
/projecteuclid.org/euclid.ba/1378729919.

Buck:2006:BNE

[BALO06] Caitlin E. Buck, Delil Gómez Portugal Aguilar, Cliff D. Litton, and Anthony O’Hagan. Bayesian nonparametric estimation of the radiocarbon calibration curve. Bayesian Analysis, 1(2):265–288, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2006/vol01/issue02/buck265-288.pdf;   http://
projecteuclid.org/euclid.ba/1340371063.

Banerjee:2017:HDB

[Ban17] Sudipto Banerjee. High-dimensional Bayesian geostatistics. Bayesian Analysis, 12(2):583–614, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1494921642.

Baragatti:2011:BVS

[Bar11] Meli Baragatti. Bayesian variable selection for probit mixed models applied to gene selection. Bayesian Analysis, 6 (2):209–229, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue02/baragatti.pdf;   http:
//projecteuclid.org/euclid.ba/1339612044.

Buchholz:2023:DCM

[BAR23] Alexander Buchholz, Daniel Ahfock, and Sylvia Richardson. Distributed computation for marginal likelihood based model choice. Bayesian Analysis, 18(2):607–638, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Distributed-Computation-for-Marginal-Likelihood-
based-Model-Choice/10.1214/22-BA1321.full.

Blackwell:2008:ERC

[BB08a] P. G. Blackwell and C. E. Buck. Estimating radiocarbon calibration curves. Bayesian Analysis, 3(2):225–248, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue02/blackwell.pdf;   http://projecteuclid.

org/euclid.ba/1340370544. See comments [HP08, Mil08] and rejoinder [BB08b].

Blackwell:2008:R

[BB08b] P. G. Blackwell and C. E. Buck. Rejoinder. Bayesian Analysis, 3(2):263–268, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2008/vol03/issue02/blackwell_
rejoinder.pdf; http://projecteuclid.org/euclid.ba/1340370547.

See [BB08a].

Barger:2010:OBE

[BB10] Kathryn Barger and John Bunge. Objective Bayesian estimation for the number of species. Bayesian Analysis, 5(4):765–785, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2010/
vol05/issue04/barger.pdf;   http://projecteuclid.org/
euclid.ba/1340110854.

Bodnar:2024:OBM

[BB24] Olha Bodnar and Taras Bodnar. Objective Bayesian metaanalysis based on generalized marginal multivariate random effects model. Bayesian Analysis, 19(2):531–564, June 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/Objective-Bayesian-Meta-Analysis-Based-
on-Generalized-Marginal-Multivariate-Random/10.1214/
23-BA1363.full.

Bazan:2006:SIR

[BBB06] Jorge L. Bazán, Marcia D. Branco, and Heleno Bolfarine. A skew item response model. Bayesian Analysis, 1(4):861–892, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/bazan.pdf;   http://projecteuclid.org/euclid.
ba/1340370945.

Bayes:2012:NRR

[BBG12] Cristian L. Bayes, Jorge L. Bazán, and Catalina García. A new robust regression model for proportions. Bayesian Analysis, 7(4):841–866, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2012/vol07/issue04/bayes.pdf;   http:
//projecteuclid.org/euclid.ba/1354024464.

Barbieri:2021:MPM

[BBGR21] Maria M. Barbieri, James O. Berger, Edward I. George, and Veronika Rocková. The median probability model and correlated variables. Bayesian Analysis, 16(4):1085– 1112, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-16/issue-4/The-
Median-Probability-Model-and-Correlated-Variables/10.
1214/20-BA1249.full.

Berger:2015:OOP

[BBS15a] James O. Berger, Jose M. Bernardo, and Dongchu Sun. Overall objective priors. Bayesian Analysis, 10(1):189–221, March 2015. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-10/issue-1/Overall-

Objective-Priors/10.1214/14-BA915.full. See comments

[Siv15, MGP15, Rou15, DL15] and rejoinder [BBS15b].

Berger:2015:R

[BBS15b] James O. Berger, Jose M. Bernardo, and Dongchu Sun. Rejoinder. Bayesian Analysis, 10(1):243–246, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-10/issue-1/Rejoinder/10.1214/15-BA943.full. See

[BBS15a] and comments [Siv15, MGP15, Rou15, DL15].

Blaauw:2011:FPA

[BC11a] Maarten Blaauw and J. Andrés Christen. Flexible paleoclimate age-depth models using an autoregressive gamma process. Bayesian Analysis, 6(3):457–474, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2011/vol06/issue03/christen.
pdf;   http://projecteuclid.org/euclid.ba/1339616472.

Bornn:2011:BCD

[BC11b] Luke Bornn and François Caron. Bayesian clustering in decomposable graphs. Bayesian Analysis, 6(4):829–846, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue04/bornn.pdf;   http://projecteuclid.org/euclid.
ba/1339616545.

Bashir:2019:PPP

[BCHJ19] Amir Bashir, Carlos M. Carvalho, P. Richard Hahn, and M. Beatrix Jones. Post-processing posteriors over precision matrices to produce sparse graph estimates. Bayesian Analysis, 14(4): 1075–1090, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1545296446.

Buchholz:2021:ATH

[BCJ21] Alexander Buchholz, Nicolas Chopin, and Pierre E. Jacob. Adaptive tuning of Hamiltonian Monte Carlo within sequential Monte Carlo. Bayesian Analysis, 16(3):745–771, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-3/Adaptive-Tuning-

of-Hamiltonian-Monte-Carlo-Within-Sequential-Monte-Carlo/
10.1214/20-BA1222.full.

Bolfarine:2024:DSS

[BCLM24] Henrique Bolfarine, Carlos M. Carvalho, Hedibert F. Lopes, and Jared S. Murray. Decoupling shrinkage and selection in Gaussian linear factor analysis. Bayesian Analysis, 19(1):181–203, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/Decoupling-Shrinkage-and-Selection-in-Gaussian-
Linear-Factor-Analysis/10.1214/22-BA1349.full.

Bassetti:2020:HSS

[BCR20] Federico Bassetti, Roberto Casarin, and Luca Rossini. Hierarchical species sampling models. Bayesian Analysis, 15(3):809–838, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Hierarchical-Species-Sampling-Models/10.1214/
19-BA1168.full.

Briol:2016:CDA

[BCT^{+}16] François-Xavier Briol, Jon Cockayne, Onur Teymur, William Weimin Yoo, Jon Cockayne, Michael Schober, and Philipp Hennig. Contributed discussion on article by chkrebtii, Campbell, calderhead, and girolami. Bayesian Analysis, 11(4):1285–1293, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.

ba/1480474950. See [CCCG16a, CCCG16b].

Browne:2006:CBL

[BD06a] William J. Browne and David Draper. A comparison of Bayesian and likelihood-based methods for fitting multilevel models. Bayesian Analysis, 1(3):473–514, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue03/draper2.
pdf;   http://projecteuclid.org/euclid.ba/1340371047.

See comments [Gel06, KN06, Lam06] and rejoinder [BD06b].

Browne:2006:R

[BD06b] William J. Browne and David Draper. Rejoinder. Bayesian Analysis, 1(3):547–550, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue03/draper_rejoinder.
pdf;   http://projecteuclid.org/euclid.ba/1340371051.

See [BD06a].

Barr:2009:CAC

[BD09] Christopher David Barr and Francesca Dominici. Comment on article by Craigmile et al. Bayesian Analysis, 4(1):37–39, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue01/barr.pdf;   http://projecteuclid.org/euclid.

ba/1340370386. See [CCL^{+}09a].

Bhadra:2017:HEU

[BDPW17] Anindya Bhadra, Jyotishka Datta, Nicholas G. Polson, and Brandon Willard. The horseshoe+ estimator of ultra-sparse signals. Bayesian Analysis, 12(4):1105–1131, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1474572263.

Boss:2024:GIG

[BDW^{+}24] Jonathan Boss, Jyotishka Datta, Xin Wang, Sung Kyun Park, Jian Kang, and Bhramar Mukherjee. Group inversegamma gamma shrinkage for sparse linear models with blockcorrelated regressors. Bayesian Analysis, 19(3):785–814, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-3/Group-Inverse-Gamma-
Gamma-Shrinkage-for-Sparse-Linear-Models-with/10.1214/
23-BA1371.full.

Baskurt:2013:HAI

[BE13] Zeynep Baskurt and Michael Evans. Hypothesis assessment and inequalities for Bayes factors and relative belief ratios. Bayesian Analysis, 8(3):569–590, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2013/vol08/issue03/evans.pdf;
 http://projecteuclid.org/euclid.ba/1378729920.

Berger:2006:COB

[Ber06a] James Berger. The case for objective Bayesian analysis. Bayesian Analysis, 1(3):385–402, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2006/vol01/issue03/berger.pdf;
http://projecteuclid.org/euclid.ba/1340371035. See com-

ments [Chr06, Dra06, Fie06a, Kad06, Kas06, Lad06, O’H06, Was06] and rejoinder [Ber06b].

Berger:2006:R

[Ber06b] James Berger. Rejoinder. Bayesian Analysis, 1(3):457– 464, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue03/berger_rejoinder.pdf;
 http://projecteuclid.org/euclid.ba/1340371045.

Bernardo:2008:CAG

[Ber08] José M. Bernardo. Comment on article by Gelman. Bayesian Analysis, 3(3):451–453, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue03/bernardo.pdf;
 http://projecteuclid.org/euclid.ba/1340370430.  See

[Gel08a].

Bernardo:2014:CAR

[Ber14] José M. Bernardo. Comment on article by Rubio and Steel. Bayesian Analysis, 9(1):23–24, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2014/vol09/issue01/bernardo.
pdf;   http://projecteuclid.org/euclid.ba/1393251765.

See [RS14a].

Bardwell:2017:BDA

[BF17] Lawrence Bardwell and Paul Fearnhead. Bayesian detection of abnormal segments in multiple time series. Bayesian Analysis, 12(1):193–218, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1456235761.

Benson:2021:BIM

[BF21] Alan Benson and Nial Friel. Bayesian inference, model selection and likelihood estimation using fast rejection sampling: The Conway–Maxwell–Poisson distribution. Bayesian Analysis, 16(3):905–931, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-

16/issue-3/Bayesian-Inference-Model-Selection-and-Likelihood-
Estimation-using-Fast-Rejection/10.1214/20-BA1230.full.

Bacallado:2022:PSP

[BFPT22] Sergio Bacallado, Stefano Favaro, Samuel Power, and Lorenzo Trippa. Perfect sampling of the posterior in the hierarchical Pitman–Yor process. Bayesian Analysis, 17(3):685–709, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-3/Perfect-Sampling-
of-the-Posterior-in-the-Hierarchical-PitmanYor-Process/
10.1214/21-BA1269.full.

Beal:2006:VBL

[BG06] Matthew J. Beal and Zoubin Ghahramani. Variational Bayesian learning of directed graphical models with hidden variables. Bayesian Analysis, 1(4):793–831, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2006/vol01/issue04/beal.pdf;
 http://projecteuclid.org/euclid.ba/1340370943.

Bryant:2013:BDA

[BG13] John R. Bryant and Patrick J. Graham. Bayesian demographic accounts: Subnational population estimation using multiple data sources. Bayesian Analysis, 8(3):591–622, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue03/bryant.pdf;   http://projecteuclid.org/
euclid.ba/1378729921.

Beliveau:2021:TIH

[BG21] Audrey Béliveau and Paul Gustafson. A theoretical investigation of how evidence flows in Bayesian network meta-analysis of disconnected networks. Bayesian Analysis, 16(3):803–823, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-3/A-Theoretical-Investigation-
of-How-Evidence-Flows-in-Bayesian-Network/10.1214/20-
BA1224.full.

Bernardi:2015:BTR

[BGP15] Mauro Bernardi, Ghislaine Gayraud, and Lea Petrella. Bayesian tail risk interdependence using quantile regression. Bayesian Analysis, 10(3):553–603, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884983.

Bianchini:2020:DPP

[BGQ20] Ilaria Bianchini, Alessandra Guglielmi, and Fernando A. Quintana. Determinantal point process mixtures via spectral density approach. Bayesian Analysis, 15(1):187–214, March 2020. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1551150619.

Beraha:2021:SHD

[BGQ21] Mario Beraha, Alessandra Guglielmi, and Fernando A. Quintana. The semi-hierarchical Dirichlet process and its application to clustering homogeneous distributions. Bayesian Analysis, 16 (4):1187–1219, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-16/issue-4/The-
Semi-Hierarchical-Dirichlet-Process-and-Its-Application-
to-Clustering/10.1214/21-BA1278.full.

Bhattacharya:2007:IRS

[BH07] S. Bhattacharya and J. Haslett. Importance re-sampling MCMC for cross-validation in inverse problems. Bayesian Analysis, 2(2):385–407, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2007/vol02/issue02/bhattacharya.pdf;   http:/
/projecteuclid.org/euclid.ba/1340393241.

Bove:2011:HPG

[BH11] Daniel Sabanés Bové and Leonhard Held. Hyper-g priors for generalized linear models. Bayesian Analysis, 6(3):387– 410, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue03/sabanes.pdf;   http://
projecteuclid.org/euclid.ba/1339616469.

Bhattacharya:2007:SAB

[Bha07] Sourabh Bhattacharya. A simulation approach to Bayesian emulation of complex dynamic computer models. Bayesian Analysis, 2(4):783–815, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2007/vol02/issue04/bhattacharya.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370715.

Bezener:2018:BSM

[BHJ18] Martin Bezener, John Hughes, and Galin Jones. Bayesian spatiotemporal modeling using hierarchical spatial priors, with applications to functional magnetic resonance imaging (with discussion). Bayesian Analysis, 13(4):1261–1313, December 2018. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-13/issue-4/Bayesian-Spatiotemporal-Modeling-Using-
Hierarchical-Spatial-Priors-with-Applications-to/10.
1214/18-BA1108.full.

Barclay:2014:CEG

[BHS14] Lorna M. Barclay, Jane L. Hutton, and Jim Q. Smith. Chain event graphs for informed missingness. Bayesian Analysis, 9(1):53–76, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2014/vol09/issue01/barclay.pdf;   http://
projecteuclid.org/euclid.ba/1393251770.

Basturk:2017:BAB

[BHvD17] Nalan Bastürk, Lennart Hoogerheide, and Herman K. van Dijk. Bayesian analysis of boundary and near-boundary evidence in econometric models with reduced rank. Bayesian Analysis, 12(3):879–917, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
12/issue-3/Bayesian-Analysis-of-Boundary-and-Near-Boundary-
Evidence-in-Econometric/10.1214/17-BA1061.full.

Bradley:2018:CEM

[BHW18] Jonathan R. Bradley, Scott H. Holan, and Christopher K. Wikle. Computationally efficient multivariate spatio-temporal models for high-dimensional count-valued data (with discussion). Bayesian Analysis, 13(1):253–310, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
13/issue-1/Computationally-Efficient-Multivariate-Spatio-
Temporal-Models-for-High-Dimensional-Count/10.1214/17-
BA1069.full.

Bickel:2020:ERP

[Bic20] David R. Bickel. An explanatory rationale for priors sharpened into Occam’s razors. Bayesian Analysis, 15(4):1299– 1321, December 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-15/issue-4/An-
Explanatory-Rationale-for-Priors-Sharpened-Into-Occams-
Razors/10.1214/19-BA1189.full.

Blei:2006:VID

[BJ06] David M. Blei and Michael I. Jordan. Variational inference for Dirichlet process mixtures. Bayesian Analysis, 1(1):121–143, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2006/vol01/issue01/blei.pdf;   http://projecteuclid.
org/euclid.ba/1340371077.

Berchuck:2022:BNP

[BJM^{+}22] Samuel I. Berchuck, Mark Janko, Felipe A. Medeiros, William Pan, and Sayan Mukherjee. Bayesian non-parametric factor analysis for longitudinal spatial surfaces. Bayesian Analysis, 17(2):435–464, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/Bayesian-Non-Parametric-Factor-Analysis-for-
Longitudinal-Spatial-Surfaces/10.1214/20-BA1253.full.

Broderick:2012:BPS

[BJP12] Tamara Broderick, Michael I. Jordan, and Jim Pitman. Beta processes, stick-breaking, and power laws. Bayesian Analysis, 7(2):439–476, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue02/broderick.pdf;   http:
//projecteuclid.org/euclid.ba/1339878895.

Barrientos:2012:SMD

[BJQ12] Andrés F. Barrientos, Alejandro Jara, and Fernando A. Quintana. On the support of MacEachern’s dependent Dirichlet processes and extensions. Bayesian Analysis, 7(2):277–310, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2012/
vol07/issue02/barrientos.pdf;   http://projecteuclid.
org/euclid.ba/1339878889.

Berman:2023:NAB

[BJS23] Brandon Berman, Wesley O. Johnson, and Weining Shen. Normal approximation for Bayesian mixed effects binomial regression models. Bayesian Analysis, 18(2):415–435, June 2023. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-2/Normal-Approximation-for-Bayesian-
Mixed-Effects-Binomial-Regression-Models/10.1214/22-
BA1312.full.

Botha:2021:PMS

[BKD21] Imke Botha, Robert Kohn, and Christopher Drovandi. Particle methods for stochastic differential equation mixed effects models. Bayesian Analysis, 16(2):575–609, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/Particle-Methods-for-Stochastic-Differential-
Equation-Mixed-Effects-Models/10.1214/20-BA1216.full.

Bodnar:2016:OBI

[BLE16] O. Bodnar, A. Link, and C. Elster. Objective Bayesian inference for a generalized marginal random effects model. Bayesian Analysis, 11(1):25–45, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1423083638.

Balakrishnan:2006:OPS

[BM06] Suhrid Balakrishnan and David Madigan. A one-pass sequential Monte Carlo method for Bayesian analysis of massive datasets. Bayesian Analysis, 1(2):345–361, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue02/madigan345-362.pdf;   http://projecteuclid.org/
euclid.ba/1340371066.

Buonaguidi:2022:BQD

[BMBV22] Bruno Buonaguidi, Antonietta Mira, Herbert Bucheli, and Viton Vitanis. Bayesian quickest detection of credit card fraud. Bayesian Analysis, 17(1):261–290, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-1/Bayesian-Quickest-Detection-of-Credit-Card-
Fraud/10.1214/20-BA1254.full.

Bernardo:2007:CNM

[BP07] José M. Bernardo and Sergio Pérez. Comparing normal means: New methods for an old problem. Bayesian Analysis, 2(1):45–58, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue01/bernardo.pdf;   http:/
/projecteuclid.org/euclid.ba/1340390062.

Barber:2008:CMM

[BP08] Jarrett J. Barber and Steven D. Prager. Combining multiple maps of line features to infer true position. Bayesian Analysis, 3(3):625–658, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue03/barber.pdf;
http://projecteuclid.org/euclid.ba/1340370441.

Barrientos:2020:BBM

[BP20] Andrés F. Barrientos and Víctor Peña. Bayesian bootstraps for massive data. Bayesian Analysis, 15(2):363–388, June 2020. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1557475224.

Burgette:2021:SPM

[BPH21] Lane F. Burgette, David Puelz, and P. Richard Hahn. A symmetric prior for multinomial probit models. Bayesian Analysis, 16(3):991–1008, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-3/A-Symmetric-Prior-for-Multinomial-Probit-
Models/10.1214/20-BA1233.full.

Broderick:2013:FAP

[BPJ13] Tamara Broderick, Jim Pitman, and Michael I. Jordan. Feature allocations, probability functions, and paintboxes. Bayesian Analysis, 8(4):801–836, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2013/vol08/issue04/broderick.
pdf;   http://projecteuclid.org/euclid.ba/1386166314.

Bae:2015:BPR

[BPSS15] Harold Bae, Thomas Perls, Martin Steinberg, and Paola Sebastiani. Bayesian polynomial regression models to fit multiple genetic models for quantitative traits. Bayesian Analysis, 10(1):

53–74, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1422468423.

Bottolo:2010:ESS

[BR10] Leonard Bottolo and Sylvia Richardson. Evolutionary stochastic search for Bayesian model exploration. Bayesian Analysis, 5 (3):583–618, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue03/bottolo.pdf;   http://
projecteuclid.org/euclid.ba/1340380542.

Burgette:2013:MSM

[BR13] Lane F. Burgette and Jerome P. Reiter. Multiple-shrinkage multinomial probit models with applications to simulating geographies in public use data. Bayesian Analysis, 8(2): 453–478, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue02/burgette.pdf;   http:/
/projecteuclid.org/euclid.ba/1369407560.

Bradley:2022:JBA

[Bra22] Jonathan R. Bradley. Joint Bayesian analysis of multiple response-types using the hierarchical generalized transformation model. Bayesian Analysis, 17(1):127–164, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-1/Joint-Bayesian-Analysis-of-Multiple-
Response-Types-Using-the-Hierarchical/10.1214/20-BA1246.
full.

Belitser:2014:APB

[BS14] Eduard Belitser and Paulo Serra. Adaptive priors based on splines with random knots. Bayesian Analysis, 9(4):859–882, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1416579182.

Bunnin:2021:BHM

[BS21] F. O. Bunnin and J. Q. Smith. A Bayesian hierarchical model for criminal investigations. Bayesian Analysis, 16(1):1–30, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-1/A-Bayesian-Hierarchical-
Model-for-Criminal-Investigations/10.1214/19-BA1192.
full.

Barrientos:2023:BIU

[BSPD23] Andrés F. Barrientos, Deborshee Sen, Garritt L. Page, and David B. Dunson. Bayesian inferences on uncertain ranks and orderings: Application to ranking players and lineups. Bayesian Analysis, 18(3):777–806, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Bayesian-Inferences-on-Uncertain-Ranks-and-
Orderings--Application-to/10.1214/22-BA1324.full.

Burdzy:2010:CAR

[Bur10] Krzysztof Burdzy. Comment on article by Robert. Bayesian Analysis, 5(2):233–236, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue02/burdzy.pdf;
http://projecteuclid.org/euclid.ba/1340218337. See [Rob10].

Bingham:2009:BOS

[BVN09] Melissa A. Bingham, Stephen B. Vardeman, and Daniel J. Nordman. Bayes one-sample and one-way random effects analyses for 3-D orientations with application to materials science. Bayesian Analysis, 4(3):607–629, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue03/bingham.
pdf;   http://projecteuclid.org/euclid.ba/1340369857.

Bonassi:2015:SMC

[BW15] Fernando V. Bonassi and Mike West. Sequential Monte Carlo with adaptive weights for approximate Bayesian computation. Bayesian Analysis, 10(1):171–187, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422468427.

Bishoyi:2020:LSR

[BWD20] Abhishek Bishoyi, Xiaojing Wang, and Dipak K. Dey. Learning semiparametric regression with missing covariates using Gaussian process models. Bayesian Analysis, 15(1):215–239, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1554775288.

Cremaschi:2023:SUM

[CAD^{+}23] Andrea Cremaschi, Raffaele Argiento, Maria De Iorio, Cai Shirong, Yap Seng Chong, Michael Meaney, and Michelle Kee. Seemingly unrelated multi-state processes: a Bayesian semiparametric approach. Bayesian Analysis, 18(3):753–775, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-3/Seemingly-Unrelated-
Multi-State-Processes-A-Bayesian-Semiparametric-Approach/
10.1214/22-BA1326.full.

Carlin:2006:CAC

[Car06] Bradley P. Carlin. Comment on article by Celeux et al. Bayesian Analysis, 1(4):675–676, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue04/carlin.
pdf;   http://projecteuclid.org/euclid.ba/1340370934.

See [CFRT06a].

Carlin:2008:ECN

[Car08] Bradley P. Carlin. Editor in Chief’s note. Bayesian Analysis, 3(3):443–444, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue03/eic.pdf;   http://
projecteuclid.org/euclid.ba/1340370428.

Carlin:2009:ECN

[Car09] Bradley P. Carlin. Editor-in-chief’s note. Bayesian Analysis, 4(4):847–850, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue04/eic.pdf;   http://
projecteuclid.org/euclid.ba/1340369827.

Casarin:2014:CAW

[Cas14] Roberto Casarin. Comment on article by Windle and Carvalho. Bayesian Analysis, 9(4):793–804, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1416579177. See [WC14b].

Cremaschi:2019:HNC

[CAS^{+}19] Andrea Cremaschi, Raffaele Argiento, Katherine Shoemaker, Christine Peterson, and Marina Vannucci. Hierarchical normalized completely random measures for robust graphical modeling. Bayesian Analysis, 14(4):1271–1301, December 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1553738429.

Castellanos:2021:MSA

[Cas21] María Eugenia Castellanos. A model selection approach for variable selection with censored data. Bayesian Analysis, 16(1): 271–300, March 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-1/A-Model-
Selection-Approach-for-Variable-Selection-with-Censored-
Data/10.1214/20-BA1207.full.

Crispino:2023:IPC

[CAV23] Marta Crispino and Isadora Antoniano-Villalobos. Informative priors for the consensus ranking in the Bayesian Mallows model. Bayesian Analysis, 18(2):391–414, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Informative-Priors-for-the-Consensus-Ranking-
in-the-Bayesian-Mallows/10.1214/22-BA1307.full.

Caron:2014:CAF

[CB14] François Caron and Luke Bornn. Comment on article by Finegold and Drton. Bayesian Analysis, 9(3):551–556, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1409921103. See

[FD14b].

Chang:2021:CBF

[CB21] Sean Chang and James O. Berger. Comparison of Bayesian and frequentist multiplicity correction for testing mutually exclusive hypotheses under data dependence. Bayesian Analysis, 16(1):111–128, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/Comparison-of-Bayesian-and-Frequentist-Multiplicity-
Correction-for-Testing-Mutually/10.1214/20-BA1196.full.

Chapple:2023:MAB

[CBC23] Andrew Chapple, Yussef Bennani, and Meredith Clement. A multi-armed Bayesian ordinal outcome utility-based sequential trial with a pairwise null clustering prior. Bayesian Analysis, 18 (2):519–546, June 2023. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-18/issue-2/A-Multi-
Armed-Bayesian-Ordinal-Outcome-Utility-Based-Sequential-
Trial/10.1214/22-BA1316.full.

Cabral:2023:CFN

[CBR23] Rafael Cabral, David Bolin, and Håvard Rue. Controlling the flexibility of non-Gaussian processes through shrinkage priors. Bayesian Analysis, 18(4):1223–1246, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-4/Controlling-the-Flexibility-of-Non-
Gaussian-Processes-Through-Shrinkage-Priors/10.1214/
22-BA1342.full.

Cressie:2015:CAF

[CC15] Noel Cressie and Raymond L. Chambers. Comment on article by Ferreira and Gamerman. Bayesian Analysis, 10(3):741–748, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1429880217. See [dG15].

Castelletti:2021:BCI

[CC21] Federico Castelletti and Guido Consonni. Bayesian causal inference in probit graphical models. Bayesian Analysis, 16(4):1113–1137, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-4/Bayesian-Causal-Inference-in-Probit-Graphical-
Models/10.1214/21-BA1260.full.

Chkrebtii:2016:BSU

[CCCG16a] Oksana A. Chkrebtii, David A. Campbell, Ben Calderhead, and Mark A. Girolami. Bayesian solution uncertainty quantification for differential equations. Bayesian Analysis, 11(4): 1239–1267, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid. org/euclid.ba/1473276259. See comments [Lys16, Das16, MYGE16, BCT^{+}16] and rejoinder [CCCG16b].

Chkrebtii:2016:R

[CCCG16b] Oksana A. Chkrebtii, David A. Campbell, Ben Calderhead, and Mark A. Girolami. Rejoinder. Bayesian Analysis, 11(4): 1295–1299, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid. org/euclid.ba/1480129462. See [CCCG16a, Lys16, Das16, MYGE16, BCT^{+}16].

Capistran:2022:ECN

[CCDT^{+}22] Marcos A. Capistrán, J. Andrés Christen, María L. Daza-Torres, Hugo Flores-Arguedas, and J. Cricelio Montesinos-López. Error control of the numerical posterior with Bayes factors in Bayesian uncertainty quantification. Bayesian Analysis, 17(2):381–403, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/Error-Control-of-the-Numerical-Posterior-
with-Bayes-Factors-in/10.1214/20-BA1255.full.

Craigmile:2009:HMB

[CCL^{+}09a] Peter F. Craigmile, Catherine A. Calder, Hongfei Li, Rajib Paul, and Noel Cressie. Hierarchical model building, fitting, and checking: A behind-the-scenes look at a Bayesian analysis of arsenic exposure pathways. Bayesian Analysis, 4(1):1–35, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2009/vol04/issue01/craigmile.pdf;   http:
//projecteuclid.org/euclid.ba/1340370385. See comments

[BD09, Dun09, Sch09] and rejoinder [CCL^{+}09b].

Craigmile:2009:R

[CCL^{+}09b] Peter F. Craigmile, Catherine A. Calder, Hongfei Li, Rajib Paul, and Noel Cressie. Rejoinder. Bayesian Analysis, 4(1):55–62, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2009/vol04/issue01/craigmile_
rejoinder.pdf; http://projecteuclid.org/euclid.ba/1340370389.

See [CCL^{+}09a].

Camerlenghi:2024:CGT

[CCO24] Federico Camerlenghi, Riccardo Corradin, and Andrea Ongaro. Contaminated Gibbs-type priors. Bayesian Analysis, 19(2):347–376, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-2/Contaminated-Gibbs-Type-Priors/10.1214/22-
BA1358.full.

Cabras:2011:GFC

[CCQ11] Stefano Cabras, María Eugenia Castellanos, and Alicia Quirós. Goodness-of-fit of conditional regression models for multiple imputation. Bayesian Analysis, 6(3):429–455, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue03/quiros.pdf;   http://projecteuclid.org/
euclid.ba/1339616471.

Castelletti:2018:LME

[CCVP18] Federico Castelletti, Guido Consonni, Marco L. Della Vedova, and Stefano Peluso. Learning Markov equivalence classes of directed acyclic graphs: An objective Bayes approach. Bayesian Analysis, 13(4):1235–1260, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1521079250.

Clarke:2013:PCP

[CCY13] Jennifer Lynn Clarke, Bertrand Clarke, and Chi-Wai Yu. Prediction in M-complete problems with limited sample size. Bayesian Analysis, 8(3):647–690, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2013/vol08/issue03/clarke.pdf;
http://projecteuclid.org/euclid.ba/1378729923.

Cong:2017:FSH

[CCZ17] Yulai Cong, Bo Chen, and Mingyuan Zhou. Fast simulation of hyperplane-truncated multivariate normal distributions. Bayesian Analysis, 12(4):1017–1037, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1488337478.

Chipeta:2015:CAF

[CD15] Michael Chipeta and Peter J. Diggle. Comment on article by Ferreira and Gamerman. Bayesian Analysis, 10(3):737–739, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1432213203. See [dG15].

Cheng:2016:BRF

[CDH16] Wen Cheng, Ian L. Dryden, and Xianzheng Huang. Bayesian registration of functions and curves. Bayesian Analysis, 11(2): 447–475, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1433162661.

Camerlenghi:2019:LNN

[CDL^{+}19] Federico Camerlenghi, David B. Dunson, Antonio Lijoi, Igor Prünster, and Abel Rodríguez. Latent nested nonparametric priors (with discussion). Bayesian Analysis, 14(4):1303–1356, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1561601089.

Celeux:2012:RRC

[CEMR12] Gilles Celeux, Mohammed El Anbari, Jean-Michel Marin, and Christian P. Robert. Regularization in regression: Comparing Bayesian and frequentist methods in a poorly informative situation. Bayesian Analysis, 7(2):477–502, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue02/celeux.
pdf;   http://projecteuclid.org/euclid.ba/1339878896.

Christen:2010:GPS

[CF10] J. Andrés Christen and Colin Fox. A general purpose sampling algorithm for continuous distributions (the t-walk). Bayesian Analysis, 5(2):263–281, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue02/christen.pdf;   http:/
/projecteuclid.org/euclid.ba/1340218339.

Chen:2023:BPR

[CFH23] Xi Chen, Farhan Feroz, and Michael Hobson. Bayesian posterior repartitioning for nested sampling. Bayesian Analysis, 18(3):695–721, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Bayesian-Posterior-Repartitioning-for-Nested-
Sampling/10.1214/22-BA1323.full.

Consonni:2018:PDO

[CFLN18] Guido Consonni, Dimitris Fouskakis, Brunero Liseo, and Ioannis Ntzoufras. Prior distributions for objective Bayesian analysis. Bayesian Analysis, 13(2):627–679, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1523671250.

Celeux:2006:DIC

[CFRT06a] G. Celeux, F. Forbes, C. P. Robert, and D. M. Titterington. Deviance information criteria for missing data models. Bayesian Analysis, 1(4):651–673, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2006/vol01/issue04/celeux.pdf;
http://projecteuclid.org/euclid.ba/1340370933. See com-

ments [Car06, Che06, Plu06, MV06, vdL06] and rejoinder [CFRT06b].

Celeux:2006:R

[CFRT06b] G. Celeux, F. Forbes, C. P. Robert, and D. M. Titterington. Rejoinder. Bayesian Analysis, 1(4):701–705, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue04/rejoinder.pdf;   http://projecteuclid.

org/euclid.ba/1340370939. See [CFRT06a] [MR2282197].

Chakraborty:2010:ASP

[CG10] Avishek Chakraborty and Alan E. Gelfand. Analyzing spatial point patterns subject to measurement error. Bayesian Analysis, 5(1):97–122, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue01/avishek.pdf;   http://
projecteuclid.org/euclid.ba/1340369794.

Chipman:2009:CAM

[CGM09] Hugh Chipman, Edward George, and Robert McCulloch. Comment on article by Monni and Tadesse. Bayesian Analysis, 4 (3):437–438, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2009/vol04/issue03/chipman.pdf;   http://
projecteuclid.org/euclid.ba/1340369846. See [MT09b].

Chipman:2022:MMM

[CGMS22] Hugh A. Chipman, Edward I. George, Robert E. McCulloch, and Thomas S. Shively. mBART: Multidimensional monotone BART. Bayesian Analysis, 17(2):515–544, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/mBART-Multidimensional-Monotone-BART/10.1214/
21-BA1259.full.

Cote:2022:BAM

[CGS22] Marie-Pier Côté, Christian Genest, and David A. Stephens. A Bayesian approach to modeling multivariate multilevel insurance claims in the presence of unsettled claims. Bayesian Analysis, 17(1):67–93, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-1/A-Bayesian-Approach-to-Modeling-Multivariate-
Multilevel-Insurance-Claims-in/10.1214/20-BA1243.full.

Chen:2016:PCR

[CGZ16] Mengjie Chen, Chao Gao, and Hongyu Zhao. Posterior contraction rates of the phylogenetic Indian buffet processes. Bayesian Analysis, 11(2):477–497, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1433511994.

Clark:2009:III

[CH09] James S. Clark and Michelle H. Hersh. Inference in incidence, infection, and impact: Co-infection of multiple hosts by multiple pathogens. Bayesian Analysis, 4(2):337–365, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue02/clark.pdf;   http://projecteuclid.org/euclid.
ba/1340370281.

Chen:2006:CAC

[Che06] Ming-Hui Chen. Comments on article by Celeux et al. Bayesian Analysis, 1(4):677–680, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2006/vol01/issue04/chen.pdf;

 http://projecteuclid.org/euclid.ba/1340370935.  See

[CFRT06a].

Caiado:2012:BSA

[CHG12] Camila C. S. Caiado, Richard W. Hobbs, and Michael Goldstein. Bayesian strategies to assess uncertainty in velocity models. Bayesian Analysis, 7(1):211–234, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue01/caiado.
pdf;   http://projecteuclid.org/euclid.ba/1339616730.

Chen:2008:BVS

[CHIK08] Ming-Hui Chen, Lan Huang, Joseph G. Ibrahim, and Sungduk Kim. Bayesian variable selection and computation for generalized linear models with conjugate priors. Bayesian Analysis, 3(3):585–613, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue03/chen.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370439.

Cui:2022:IBN

[CHMK22] Tianyu Cui, Aki Havulinna, Pekka Marttinen, and Samuel Kaski. Informative Bayesian neural network priors for weak signals. Bayesian Analysis, 17(4):1121–1151, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Informative-Bayesian-Neural-Network-
Priors-for-Weak-Signals/10.1214/21-BA1291.full.

Christen:2006:SUS

[Chr06] J. Andrés Christen. Stop using ‘subjective’ to refer to Bayesian analyses (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1(3):421–422, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2006/vol01/issue03/christen.
pdf;   http://projecteuclid.org/euclid.ba/1340371037.

See [Ber06a, Gol06a].

Christensen:2009:IBE

[Chr09] Ronald Christensen. Inconsistent Bayesian estimation. Bayesian Analysis, 4(4):759–762, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2009/vol04/issue04/christensen.
pdf;   http://projecteuclid.org/euclid.ba/1340369823.

Christensen:2024:IBN

[Chr24] Dennis Christensen. Inference for Bayesian nonparametric models with binary response data via permutation counting. Bayesian Analysis, 19(1):293–318, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/Inference-for-Bayesian-Nonparametric-Models-
with-Binary-Response-Data-via/10.1214/22-BA1353.full.

Chen:2006:RBP

[CI06] Ming-Hui Chen and Joseph G. Ibrahim. The relationship between the power prior and hierarchical models. Bayesian Analysis, 1(3):551–574, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2006/vol01/issue03/chenibra.pdf;
 http://projecteuclid.org/euclid.ba/1340371052.

Chen:2009:CAY

[CK09] Ming-Hui Chen and Sungduk Kim. Comments on article by Yin. Bayesian Analysis, 4(2):209–212, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2009/vol04/issue02/chen.pdf;
 http://projecteuclid.org/euclid.ba/1340370273.  See

[Yin09a].

Cao:2020:HDP

[CKG20] Xuan Cao, Kshitij Khare, and Malay Ghosh. High-dimensional posterior consistency for hierarchical non-local priors in regression. Bayesian Analysis, 15(1):241–262, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1555660820.

Cano:2007:IPO

[CKS07] Juan Antonio Cano, Mathieu Kessler, and Diego Salmerón. Integral priors for the one way random effects model. Bayesian Analysis, 2(1):59–67, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue01/cano.pdf;   http:/
/projecteuclid.org/euclid.ba/1340390063.

Cai:2020:BNM

[CKY20] Qingpo Cai, Jian Kang, and Tianwei Yu. Bayesian network marker selection via the thresholded graph Laplacian Gaussian prior. Bayesian Analysis, 15(1):79–102, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1546657330.

Cao:2024:BIH

[CL24a] Xuan Cao and Kyoungjae Lee. Bayesian inference on hierarchical nonlocal priors in generalized linear models. Bayesian Analysis, 19(1):99–122, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-1/Bayesian-Inference-on-Hierarchical-
Nonlocal-Priors-in-Generalized-Linear-Models/10.1214/
22-BA1350.full.

Cao:2024:CSB

[CL24b] Xuan Cao and Kyoungjae Lee. Consistent and scalable Bayesian joint variable and graph selection for disease diagnosis leveraging functional brain network. Bayesian Analysis, 19(3):895–923, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-3/Consistent-and-Scalable-Bayesian-Joint-Variable-
and-Graph-Selection-for/10.1214/23-BA1376.full.

Clarke:2010:DPT

[Cla10] Bertrand Clarke. Desiderata for a predictive theory of statistics. Bayesian Analysis, 5(2):283–318, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue02/clarke.pdf;
http://projecteuclid.org/euclid.ba/1340218340.

Clarke:2012:CAS

[Cla12] Bertrand Clarke. Comment on article by Sancetta. Bayesian Analysis, 7(1):37–44, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue01/clarke.pdf;   http://
projecteuclid.org/euclid.ba/1339616722. See [San12b].

Chkrebtii:2016:CDA

[CLH^{+}16] Oksana A. Chkrebtii, Scotland Leman, Andrew Hoegh, Reihaneh Entezari, Radu V. Craiu, Jeffrey S. Rosenthal, Abdolreza Mohammadi, Maurits Kaptein, Luca Martino, Rafael B. Stern, and Francisco Louzada. Contributed discussion on article by Pratola. Bayesian Analysis, 11(3):929–943, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1472829062. See [Pra16a].

Cowell:2007:GMD

[CLM07] R. G. Cowell, S. L. Lauritzen, and J. Mortera. A gamma model for DNA mixture analyses. Bayesian Analysis, 2(2):333–348, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2007/
vol02/issue02/cowell.pdf;   http://projecteuclid.org/
euclid.ba/1340393238.

Casarin:2015:BBM

[CLMtH15] Roberto Casarin, Fabrizio Leisen, German Molina, and Enrique ter Horst. A Bayesian beta Markov random field calibration of the term structure of implied risk neutral densities. Bayesian Analysis, 10(4):791–819, December 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1434980946.

Carvalho:2010:PLG

[CLPT10] Carlos M. Carvalho, Hedibert F. Lopes, Nicholas G. Polson, and Matt A. Taddy. Particle learning for general mixtures. Bayesian Analysis, 5(4):709–740, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue04/carvalho.pdf;
 http://projecteuclid.org/euclid.ba/1340110852.

Carlin:2013:CAM

[CM13] Bradley P. Carlin and Thomas A. Murray. Comment on article by Müller and Mitra. Bayesian Analysis, 8(2):303–310, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue02/carlin.pdf;   http://projecteuclid.org/
euclid.ba/1369407551. See [MM13a].

Casella:2014:CAM

[CMG14] George Casella, Elías Moreno, and F. Javier Girón. Cluster analysis, model selection, and prior distributions on models. Bayesian Analysis, 9(3):613–658, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1409921108.

Cabras:2015:ABC

[CNR15] Stefano Cabras, Maria Eugenia Castellanos Nueda, and Erlis Ruli. Approximate Bayesian computation by modelling summary statistics in a quasi-likelihood framework. Bayesian Analysis, 10(2):411–439, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1422884980.

Clancy:2008:BEB

[CO08] Damian Clancy and Philip D. O’Neill. Bayesian estimation of the basic reproduction number in stochastic epidemic models. Bayesian Analysis, 3(4):737–757, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2008/vol03/issue04/clancy.
pdf;   http://projecteuclid.org/euclid.ba/1340370407.

Cockayne:2019:BCG

[COIG19] Jon Cockayne, Chris J. Oates, Ilse C. F. Ipsen, and Mark Girolami. A Bayesian conjugate gradient method (with discussion). Bayesian Analysis, 14(3):937–1012, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-14/issue-3/A-Bayesian-Conjugate-Gradient-Method-
with-Discussion/10.1214/19-BA1145.full. Includes 6 discus-

sions and a rejoinder from the authors.

Crainiceanu:2009:CAY

[Cra09] Ciprian M. Crainiceanu. Comments on article by Yin. Bayesian Analysis, 4(2):213–215, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2009/vol04/issue02/crainiceanu.
pdf;   http://projecteuclid.org/euclid.ba/1340370274.

See [Yin09a].

Cruz-Reyes:2023:IHS

[CRAL23] Danna L. Cruz-Reyes, Renato M. Assunção, and Rosangela H. Loschi. Inducing high spatial correlation with randomly edge-weighted neighborhood graphs. Bayesian Analysis, 18(4):1247–1281, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Inducing-High-Spatial-Correlation-with-Randomly-
Edge-Weighted-Neighborhood-Graphs/10.1214/23-BA1390.
full.

Cook:2007:CAD

[CS07] Samantha R. Cook and Elizabeth A. Stuart. Comment on article by Dominici et al. Bayesian Analysis, 2(1):31–35, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue01/cook.pdf;   http://projecteuclid.org/euclid.

ba/1340390059. See [DZP^{+}07a].

Carbonetto:2012:SVI

[CS12] Peter Carbonetto and Matthew Stephens. Scalable variational inference for Bayesian variable selection in regression, and its accuracy in genetic association studies. Bayesian Analysis, 7(1):73–108, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2012/vol07/issue01/carbonetto.pdf;   http://
projecteuclid.org/euclid.ba/1339616726.

Cano:2013:IPC

[CS13] Juan Antonio Cano and Diego Salmerón. Integral priors and constrained imaginary training samples for nested and non-nested Bayesian model comparison. Bayesian Analysis, 8(2):361–380, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue02/cano.pdf;   http:/
/projecteuclid.org/euclid.ba/1369407556.

Cheng:2016:BFS

[CS16a] Chin-I. Cheng and Paul L. Speckman. Bayes factors for smoothing spline ANOVA. Bayesian Analysis, 11(4):957–975, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1444653650.

Collazo:2016:NFN

[CS16b] Rodrigo A. Collazo and Jim Q. Smith. A new family of non-local priors for chain event graph model selection. Bayesian Analysis, 11(4):1165–1201, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1448852254.

Costa:2015:SMD

[CSN^{+}15] Lilia Costa, Jim Smith, Thomas Nichols, James Cussens, Eugene P. Duff, and Tamar R. Makin. Searching multiregression dynamic models of resting-state fMRI networks using integer programming. Bayesian Analysis, 10(2):441–478, June 2015. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1422884981.

Crepet:2011:BNM

[CT11] Amélie Crépet and Jessica Tressou. Bayesian nonparametric model for clustering individual co-exposure to pesticides found in the French diet. Bayesian Analysis, 6(1):127–144, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue01/crepet.pdf;   http://projecteuclid.org/
euclid.ba/1339611943.

Carvalho:2023:BIW

[CVCB23] Luiz M. Carvalho, Daniel A. M. Villela, Flavio C. Coelho, and Leonardo S. Bastos. Bayesian inference for the weights in logarithmic pooling. Bayesian Analysis, 18(1):223–251, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-1/Bayesian-Inference-
for-the-Weights-in-Logarithmic-Pooling/10.1214/22-BA1311.
full.

Casarin:2012:BMS

[CVL12] Roberto Casarin, Luciana Dalla Valle, and Fabrizio Leisen. Bayesian model selection for beta autoregressive processes. Bayesian Analysis, 7(2):385–410, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue02/casarin.
pdf;   http://projecteuclid.org/euclid.ba/1339878893.

Carvalho:2007:DMV

[CW07] Carlos M. Carvalho and Mike West. Dynamic matrix-variate graphical models. Bayesian Analysis, 2(1):69–97, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue01/carvalho.pdf;   http://projecteuclid.org/euclid.
ba/1340390064.

Cao:2010:MSS

[CZ10] Jing Cao and Song Zhang. Measuring statistical significance for full Bayesian methods in microarray analyses. Bayesian Analysis, 5(2):413–427, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue02/cao.pdf;   http://
projecteuclid.org/euclid.ba/1340218344.

Cassese:2019:BNS

[CZGV19] Alberto Cassese, Weixuan Zhu, Michele Guindani, and Marina Vannucci. A Bayesian nonparametric spiked process prior for dynamic model selection. Bayesian Analysis, 14(2):553–572, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1533866669.

Dahl:2007:CAJ

[Dah07] David B. Dahl. Comment on article by Jain and Neal. Bayesian Analysis, 2(3):473–477, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue03/dahl.pdf;
 http://projecteuclid.org/euclid.ba/1340370721.  See

[JN07b].

Dahl:2009:MCC

[Dah09] David B. Dahl. Modal clustering in a class of product partition models. Bayesian Analysis, 4(2):243–264, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue02/dahl.pdf;   http://projecteuclid.org/euclid.
ba/1340370277.

Dass:2016:CAC

[Das16] Sarat C. Dass. Comment on article by Chkrebtii, Campbell, Calderhead, and Girolami. Bayesian Analysis, 11(4):1275–1277, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.

ba/1480474949. See [CCCG16a, CCCG16b].

Datta:2019:SDM

[DBHG19] Abhirup Datta, Sudipto Banerjee, James S. Hodges, and Leiwen Gao. Spatial disease mapping using directed acyclic graph auto-regressive (DAGAR) models. Bayesian Analysis, 14(4): 1221–1244, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1570068455.

Pereira:2008:CST

[dBPSW08] Carlos A. de B. Pereira, Julio Michael Stern, and Sergio Wechsler. Can a significance test be genuinely Bayesian? Bayesian Analysis, 3(1):79–100, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2008/vol03/issue01/pereira.
pdf;   http://projecteuclid.org/euclid.ba/1340370562.

deCarvalho:2013:BNR

[dCJHdC13] Vanda Inácio de Carvalho, Alejandro Jara, Timothy E. Hanson, and Miguel de Carvalho. Bayesian nonparametric ROC regression modeling. Bayesian Analysis, 8(3):623– 646, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/carvalho.pdf;   http:/
/projecteuclid.org/euclid.ba/1378729922.

Das:2008:BSE

[DCKW08] Sonali Das, Ming-Hui Chen, Sungduk Kim, and Nicholas Warren. A Bayesian structural equations model for multilevel data with missing responses and missing covariates. Bayesian Analysis, 3(1):197–224, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue01/chen.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370567.

deCarvalho:2019:GBI

[dCPB19] Miguel de Carvalho, Garritt L. Page, and Bradley J. Barney. On the geometry of Bayesian inference. Bayesian Analysis, 14 (4):1013–1036, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1533866667.

Dukic:2007:BHM

[DD07] Vanja Dukić and James Dignam. Bayesian hierarchical multiresolution hazard model for the study of time-dependent failure patterns in early stage breast cancer. Bayesian Analysis, 2(3):591–609, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue03/dukic.pdf;   http:
//projecteuclid.org/euclid.ba/1340370728.

Durante:2018:BIT

[DD18] Daniele Durante and David B. Dunson. Bayesian inference and testing of group differences in brain networks. Bayesian Analysis, 13(1):29–58, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1479179031.

Drovandi:2022:EMA

[DEGP22] Christopher Drovandi, Richard G. Everitt, Andrew Golightly, and Dennis Prangle. Ensemble MCMC: Accelerating pseudomarginal MCMC for state space models using the ensemble Kalman filter. Bayesian Analysis, 17(1):223–260, ???? 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-1/Ensemble-MCMC--Accelerating-Pseudo-
Marginal-MCMC-for-State-Space/10.1214/20-BA1251.full.

Didelot:2011:LFE

[DEJL11] Xavier Didelot, Richard G. Everitt, Adam M. Johansen, and Daniel J. Lawson. Likelihood-free estimation of model evidence. Bayesian Analysis, 6(1):49–76, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue01/didelot.
pdf;   http://projecteuclid.org/euclid.ba/1339611941.

Desgagne:2013:FRB

[Des13] Alain Desgagné. Full robustness in Bayesian modelling of a scale parameter. Bayesian Analysis, 8(1):187–220, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/

issue01/desgagne.pdf;   http://projecteuclid.org/euclid.
ba/1362406657.

Delatola:2011:BNM

[DG11] Eleni-Ioanna Delatola and Jim E. Griffin. Bayesian nonparametric modelling of the return distribution with stochastic volatility. Bayesian Analysis, 6(4):901–926, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue04/griffin.
pdf;   http://projecteuclid.org/euclid.ba/1339616547.

Datta:2013:APB

[DG13] Jyotishka Datta and Jayanta K. Ghosh. Asymptotic properties of Bayes risk for the horseshoe prior. Bayesian Analysis, 8(1):111– 132, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2013/vol08/issue01/datta.pdf;   http://projecteuclid.
org/euclid.ba/1362406654.

daSilvaFerreira:2015:ODG

[dG15] Gustavo da Silva Ferreira and Dani Gamerman. Optimal design in geostatistics under preferential sampling. Bayesian Analysis, 10(3):711–735, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1424441439. See comments [CD15, CC15,

Zid15] and rejoinder [dSFG15].

DiLucca:2013:SCB

[DGMQ13] Maria Anna Di Lucca, Alessandra Guglielmi, Peter Müller, and Fernando A. Quintana. A simple class of Bayesian nonparametric autoregression models. Bayesian Analysis, 8(1):63–88, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/lucca.pdf;   http://projecteuclid.org/euclid.
ba/1362406652.

Duan:2009:MST

[DGS09] Jason A. Duan, Alan E. Gelfand, and C. F. Sirmans. Modeling space-time data using stochastic differential equations. Bayesian Analysis, 4(4):733–758, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2009/vol04/issue04/duan.pdf;
 http://projecteuclid.org/euclid.ba/1340369822.

Ding:2012:NBS

[DHDC12] Mingtao Ding, Lihan He, David Dunson, and Lawrence Carin. Nonparametric Bayesian segmentation of a multivariate inhomogeneous space-time Poisson process. Bayesian Analysis, 7 (4):813–840, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2012/vol07/issue04/carin.pdf;   http:
//projecteuclid.org/euclid.ba/1354024463.

DeYoreo:2015:FNM

[DK15] Maria DeYoreo and Athanasios Kottas. A fully nonparametric modeling approach to binary regression. Bayesian Analysis, 10 (4):821–847, December 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1437137636.

Dey:2007:QSQ

[DL07] Dipak K. Dey and Junfeng Liu. A quantitative study of quantile based direct prior elicitation from expert opinion. Bayesian Analysis, 2(1):137–166, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue01/dey.pdf;   http://
projecteuclid.org/euclid.ba/1340390066.

Datta:2015:CAB

[DL15] Gauri Sankar Datta and Brunero Liseo. Comment on article by Berger, Bernardo, and Sun. Bayesian Analysis, 10(1):237–241, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). See [BBS15a].

Dhara:2020:NBS

[DLPS20] Kumaresh Dhara, Stuart Lipsitz, Debdeep Pati, and Debajyoti Sinha. A new Bayesian single index model with or without covariates missing at random. Bayesian Analysis, 15(3):759–780, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/A-New-Bayesian-Single-Index-Model-with-or-
without-Covariates/10.1214/19-BA1170.full.

Denham:2007:GAE

[DM07a] Robert Denham and Kerrie Mengersen. Geographically assisted elicitation of expert opinion for regression models. Bayesian Analysis, 2(1):99–135, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2007/vol02/issue01/denham.pdf;
http://projecteuclid.org/euclid.ba/1340390065.

Druilhet:2007:IHC

[DM07b] Pierre Druilhet and Jean-Michel Marin. Invariant HPD credible sets and MAP estimators. Bayesian Analysis, 2(4): 681–691, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue04/druilhet.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370710.

Dawid:2015:BMS

[DM15a] A. Philip Dawid and Monica Musio. Bayesian model selection based on proper scoring rules. Bayesian Analysis, 10(2):479– 499, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.

ba/1423083641. See comments [KB15, HP15, GMR15] and rejoinder [DM15b].

Dawid:2015:R

[DM15b] A. Philip Dawid and Monica Musio. Rejoinder. Bayesian Analysis, 10(2):517–521, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1430830145. See [DM15a].

Dawid:2016:SEE

[DMF16] A. Philip Dawid, Monica Musio, and Stephen E. Fienberg. From statistical evidence to evidence of causality. Bayesian Analysis, 11(3):725–752, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1440594950.

deOliveira:2022:BCC

[dOAL^{+}22] Guilherme Lopes de Oliveira, Raffaele Argiento, Rosangela Helena Loschi, Renato Martins Assunção, Fabrizio Ruggeri, and Márcia D’Elia Branco. Bias correction in clustered underreported data. Bayesian Analysis, 17(1):95–126, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-1/Bias-Correction-
in-Clustered-Underreported-Data/10.1214/20-BA1244.full.

Dobra:2013:CAS

[Dob13] Adrian Dobra. Comment on article by Scutari. Bayesian Analysis, 8(3):533–538, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue03/dobra.pdf;   http:
//projecteuclid.org/euclid.ba/1378729915. See [Scu13a].

Druilhet:2012:ICA

[DP12] Pierre Druilhet and Denys Pommeret. Invariant conjugate analysis for exponential families. Bayesian Analysis, 7(4): 903–916, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue04/druilhet.pdf;   http:/
/projecteuclid.org/euclid.ba/1354024467.

Drovandi:2016:EAB

[DPM16] Christopher C. Drovandi, Anthony N. Pettitt, and Roy A. Mc-Cutchan. Exact and approximate Bayesian inference for low integer-valued time series models with intractable likelihoods. Bayesian Analysis, 11(2):325–352, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1429543852.

Donnet:2016:BIP

[DR16] Sophie Donnet and Judith Rousseau. Bayesian inference for partially observed multiplicative intensity processes. Bayesian Analysis, 11(1):151–190, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1425492492.

Draper:2006:CCC

[Dra06] David Draper. Coherence and calibration: comments on subjectivity and “objectivity” in Bayesian analysis (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1 (3):423–428, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue03/draper.pdf;   http://
projecteuclid.org/euclid.ba/1340371038. See [Ber06a,

Gol06a].

DeYoreo:2017:BMM

[DRH17] Maria DeYoreo, Jerome P. Reiter, and D. Sunshine Hillygus. Bayesian mixture models with focused clustering for mixed ordinal and nominal data. Bayesian Analysis, 12(3):679–703, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1471454533.

Donnet:2017:PCR

[DRRS17] Sophie Donnet, Vincent Rivoirard, Judith Rousseau, and Catia Scricciolo. Posterior concentration rates for counting processes with Aalen multiplicative intensities. Bayesian Analysis, 12(1): 53–87, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1451333725.

Ferreira:2015:R

[dSFG15] Gustavo da Silva Ferreira and Dani Gamerman. Rejoinder. Bayesian Analysis, 10(3):753–758, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1431607820. See [dG15].

Dupre:2009:NAR

[DT09] Maurice J. Dupré and Frank J. Tipler. New axioms for rigorous Bayesian probability. Bayesian Analysis, 4(3):599–606, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue03/dupre.pdf;   http://projecteuclid.org/euclid.
ba/1340369856.

Drovandi:2018:IEF

[DT18] Christopher C. Drovandi and Minh-Ngoc Tran. Improving the efficiency of fully Bayesian optimal design of experiments using randomised quasi-Monte Carlo. Bayesian Analysis, 13(1):139– 162, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1483066880.

deTibeiro:2010:CAI

[dTM10] Jules J. S. de Tibeiro and Duncan J. Murdoch. Correspondence analysis with incomplete paired data using Bayesian imputation. Bayesian Analysis, 5(3):519–532, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue03/tibeiro.
pdf;   http://projecteuclid.org/euclid.ba/1340380539.

Dunson:2009:CAC

[Dun09] David B. Dunson. Comment on article by Craigmile et al. Bayesian Analysis, 4(1):41–43, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue01/dunson.
pdf;   http://projecteuclid.org/euclid.ba/1340370387.

See [CCL^{+}09a].

Dablander:2024:DBF

[DvdBWL24] Fabian Dablander, Don van den Bergh, Eric-Jan Wagenmakers, and Alexander Ly. Default Bayes factors for testing the (in)equality of several population variances. Bayesian Analysis, 19(3):699–723, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-3/Default-Bayes-Factors-for-Testing-the-Inequality-
of-Several-Population/10.1214/23-BA1369.full.

DeBlasi:2013:BED

[DW13] Pierpaolo De Blasi and Stephen G. Walker. Bayesian estimation of the discrepancy with misspecified parametric models. Bayesian Analysis, 8(4):781–800, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2013/vol08/issue04/blasi.pdf;
 http://projecteuclid.org/euclid.ba/1386166313.

Dawkins:2021:WCA

[DWM^{+}21] Laura C. Dawkins, Daniel B. Williamson, Kerrie L. Mengersen, Lidia Morawska, Rohan Jayaratne, and Gavin Shaddick. Where is the clean air? A Bayesian decision framework for personalised cyclist route selection using R-INLA. Bayesian Analysis, 16(1):61–91, March 2021. ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-1/Where-Is-the-Clean-
Air-A-Bayesian-Decision-Framework-for/10.1214/19-BA1193.
full.

Dominici:2007:DEM

[DZP^{+}07a] Francesca Dominici, Scott L. Zeger, Giovanni Parmigiani, Joanne Katz, and Parul Christian. Does the effect of micronutrient supplementation on neonatal survival vary with respect to the percentiles of the birth weight distribution? Bayesian Analysis, 2(1):1–30, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue01/dominici.pdf;   http:/
/projecteuclid.org/euclid.ba/1340390058. See comments

[CS07, RC07].

Dominici:2007:R

[DZP^{+}07b] Francesca Dominici, Scott L. Zeger, Giovanni Parmigiani, Joanne Katz, and Parul Christian. Rejoinder. Bayesian Analysis, 2(1):43–44, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2007/vol02/issue01/dominici_
rejoinder.pdf; http://projecteuclid.org/euclid.ba/1340390061.

Elliott:2019:MPS

[EDF^{+}19] Lloyd T. Elliott, Maria De Iorio, Stefano Favaro, Kaustubh Adhikari, and Yee Whye Teh. Modeling population structure under hierarchical Dirichlet processes. Bayesian Analysis, 14(2):313– 339, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1526695351.

Earls:2017:VBF

[EH17] Cecilia Earls and Giles Hooker. Variational Bayes for functional data registration, smoothing, and prediction. Bayesian Analysis, 12(2):557–582, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1469553352.

Evans:2006:CPD

[EM06] Michael Evans and Hadas Moshonov. Checking for priordata conflict. Bayesian Analysis, 1(4):893–914, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/evans.pdf;   http://projecteuclid.org/euclid.
ba/1340370946.

Eaton:2013:LBP

[EMS13] Morris L. Eaton, Robb J. Muirhead, and Adina I. Soaita. On the limiting behavior of the “Probability of Claiming Superiority” in a Bayesian context. Bayesian Analysis, 8(1):221–232, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/

issue01/eaton.pdf;   http://projecteuclid.org/euclid.
ba/1362406658.

Fuquene:2009:CRB

[FCP09] Jairo A. Fúquene, John D. Cook, and Luis R. Pericchi. A case for robust Bayesian priors with applications to clinical trials. Bayesian Analysis, 4(4):817–846, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue04/fuquene.
pdf;   http://projecteuclid.org/euclid.ba/1340369826.

Finegold:2014:R

[FD14a] Michael Finegold and Mathias Drton. Rejoinder. Bayesian Analysis, 9(3):591–596, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1409921106. See [FD14b].

Finegold:2014:RBG

[FD14b] Michael Finegold and Mathias Drton. Robust Bayesian graphical modeling using Dirichlet t-distributions. Bayesian Analysis, 9(3):521–550, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid. org/euclid.ba/1409921102. See comments [CB14, Sha14a, Ano14a] and rejoinder [FD14a].

Fearnhead:2011:CAW

[Fea11] Paul Fearnhead. Comment on article by Wyse et al. Bayesian Analysis, 6(4):529–532, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2011/vol06/issue04/fearnhead.
pdf;   http://projecteuclid.org/euclid.ba/1339616533.

See [WFR11a].

Ferreira:2012:CAL

[Fer12] Marco A. R. Ferreira. Comment on article by Lum and Gelfand. Bayesian Analysis, 7(2):271–272, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2012/vol07/issue02/ferreira.
pdf;   http://projecteuclid.org/euclid.ba/1339878887.

See [LG12b].

Filippi:2017:BNA

[FH17] Sarah Filippi and Chris C. Holmes. A Bayesian nonparametric approach to testing for dependence between random variables. Bayesian Analysis, 12(4):919–938, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1474463236.

Fuglstad:2020:IJP

[FHK^{+}20] Geir-Arne Fuglstad, Ingeborg Gullikstad Hem, Alexander Knight, Håvard Rue, and Andrea Riebler. Intuitive joint priors for variance parameters. Bayesian Analysis, 15(4):1109– 1137, December 2020. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-15/issue-4/Intuitive-
Joint-Priors-for-Variance-Parameters/10.1214/19-BA1185.
full.

Fritsch:2009:ICC

[FI09] Arno Fritsch and Katja Ickstadt. Improved criteria for clustering based on the posterior similarity matrix. Bayesian Analysis, 4(2):367–391, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2009/vol04/issue02/fritsch.pdf;   http://
projecteuclid.org/euclid.ba/1340370282.

Fienberg:2006:DIM

[Fie06a] Stephen E. Fienberg. Does it make sense to be an “objective Bayesian”? (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1(3):429–432, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2006/vol01/issue03/fienberg.
pdf;   http://projecteuclid.org/euclid.ba/1340371039.

See [Ber06a, Gol06a].

Fienberg:2006:WDB

[Fie06b] Stephen E. Fienberg. When did Bayesian inference become “Bayesian”? Bayesian Analysis, 1(1):1–40, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue01/fienberg.pdf;   http://projecteuclid.org/euclid.
ba/1340371071.

Fitch:2014:PCS

[FJM14] A. Marie Fitch, M. Beatrix Jones, and Hélène Massam. The performance of covariance selection methods that consider decomposable models only. Bayesian Analysis, 9(3):659–684, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1409921109.

Ferreira:2008:DLS

[FJS08] José T. A. S. Ferreira, Miguel A. Juárez, and Mark F. J. Steel. Directional log-spline distributions. Bayesian Analysis, 3(2):297–316, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2008/vol03/issue02/ferreira.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370549.

Favaro:2016:SBR

[FLN^{+}16] S. Favaro, A. Lijoi, C. Nava, B. Nipoti, I. Prünster, and Y. W. Teh. On the stick-breaking representation for homogeneous NR-MIs. Bayesian Analysis, 11(3):697–724, September 2016. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1440594949.

Faulkner:2018:LAS

[FM18] James R. Faulkner and Vladimir N. Minin. Locally adaptive smoothing with Markov random fields and shrinkage priors. Bayesian Analysis, 13(1):225–252, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1487905413.

Forastiere:2018:PPP

[FMM18] Laura Forastiere, Fabrizia Mealli, and Luke Miratrix. Posterior predictive p-values with Fisher randomization tests in noncompliance settings: Test statistics vs discrepancy measures. Bayesian Analysis, 13(3):681–701, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1502935324.

Friel:2016:EMC

[FMO16] Nial Friel, Antonietta Mira, and Chris. J. Oates. Exploiting multi-core architectures for reduced-variance estimation with intractable likelihoods. Bayesian Analysis, 11(1):215–245, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1428516724.

Ford:2011:BSM

[FMV11] Eric B. Ford, Althea V. Moorhead, and Dimitri Veras. A Bayesian surrogate model for rapid time series analysis and application to exoplanet observations. Bayesian Analysis, 6 (3):475–499, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue03/ford.pdf;   http:/
/projecteuclid.org/euclid.ba/1339616473.

Fouskakis:2022:PEP

[FN22] Dimitris Fouskakis and Ioannis Ntzoufras. Power-expected-posterior priors as mixtures of g-priors in normal linear models. Bayesian Analysis, 17(4):1073–1099, December 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Power-Expected-Posterior-Priors-as-
Mixtures-of-g-Priors-in/10.1214/21-BA1288.full.

Fouskakis:2015:PEP

[FND15] Dimitris Fouskakis, Ioannis Ntzoufras, and David Draper. Powerexpected-posterior priors for variable selection in Gaussian linear models. Bayesian Analysis, 10(1):75–107, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1422468424.

Fouskakis:2018:PEP

[FNP18] Dimitris Fouskakis, Ioannis Ntzoufras, and Konstantinos Perrakis. Power-expected-posterior priors for generalized linear models. Bayesian Analysis, 13(3):721–748, September 2018. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1507341641.

Forbes:2014:CAW

[For14] Catherine Scipione Forbes. Comment on article by Windle and Carvalho. Bayesian Analysis, 9(4):805–808, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1416579178. See

[WC14b].

Fraley:2009:CAM

[Fra09] Chris Fraley. Comment on article by Monni and Tadesse. Bayesian Analysis, 4(3):439–447, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue03/fraley.
pdf;   http://projecteuclid.org/euclid.ba/1340369847.

See [MT09b].

FreitasLopes:2011:CAH

[Fre11] Hedibert Freitas Lopes. Comment on article by Hoff. Bayesian Analysis, 6(2):203–204, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue02/lopes.pdf;   http:
//projecteuclid.org/euclid.ba/1339612042. See [Hof11b].

French:2012:CAA

[Fre12] Simon French. Comment on article by Albert et al. Bayesian Analysis, 7(3):533–536, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue03/french.
pdf;   http://projecteuclid.org/euclid.ba/1346158772.

See [ADGJ^{+}12a].

Fruhwirth-Schnatter:2008:CAR

[FS08] Sylvia Frühwirth-Schnatter. Comment on article by Rydén. Bayesian Analysis, 3(4):689–697, December 2008. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue04/fruehwirth.pdf;   http://projecteuclid.org/euclid.
ba/1340370403. See [Ryd08a].

Freeman:2011:DST

[FS11] Guy Freeman and Jim Q. Smith. Dynamic staged trees for discrete multivariate time series: forecasting, model selection and causal analysis. Bayesian Analysis, 6(2): 279–305, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue02/freeman.pdf;   http://
projecteuclid.org/euclid.ba/1339612047.

FreitasLopes:2008:SDF

[FSG08] Hedibert Freitas Lopes, Esther Salazar, and Dani Gamerman. Spatial dynamic factor analysis. Bayesian Analysis, 3(4):759–792, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue04/lopes.pdf;   http:
//projecteuclid.org/euclid.ba/1340370408.

Fruhwirth-Schnatter:2021:GMF

[FSMWG21] Sylvia Frühwirth-Schnatter, Gertraud Malsiner-Walli, and Bettina Grün. Generalized mixtures of finite mixtures and telescoping sampling. Bayesian Analysis, 16(4):1279–1307, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-4/Generalized-Mixtures-
of-Finite-Mixtures-and-Telescoping-Sampling/10.1214/
21-BA1294.full.

Fabrizi:2012:BEL

[FT12] Enrico Fabrizi and Carlo Trivisano. Bayesian estimation of lognormal means with finite quadratic expected loss. Bayesian Analysis, 7(4):975–996, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue04/fabrizi.pdf;
 http://projecteuclid.org/euclid.ba/1354024469.

Fearnhead:2013:ASM

[FT13] Paul Fearnhead and Benjamin M. Taylor. An adaptive sequential Monte Carlo sampler. Bayesian Analysis, 8(2):411–438, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue02/fearnhead.pdf;   http://projecteuclid.
org/euclid.ba/1369407558.

Ferreira:2006:MSH

[FWLH06] Marco A. R. Ferreira, Mike West, Herbert K. H. Lee, and David M. Higdon. Multi-scale and hidden resolution time series models. Bayesian Analysis, 1(4):947–967, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/ferreira.pdf;   http://projecteuclid.org/euclid.
ba/1340370948.

Grollemund:2019:BFL

[GABP19] Paul-Marie Grollemund, Christophe Abraham, Meïli Baragatti, and Pierre Pudlo. Bayesian functional linear regression with sparse step functions. Bayesian Analysis, 14(1):111–135, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1524103229.

Gagnon:2023:RAC

[Gag23] Philippe Gagnon. Robustness against conflicting prior information in regression. Bayesian Analysis, 18(3):841–864, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-3/Robustness-Against-
Conflicting-Prior-Information-in-Regression/10.1214/
22-BA1330.full.

Griffin:2010:ING

[GB10] Jim E. Griffin and Philip J. Brown. Inference with normalgamma prior distributions in regression problems. Bayesian Analysis, 5(1):171–188, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue01/griffin2.pdf;
 http://projecteuclid.org/euclid.ba/1340369797.

Guhaniyogi:2012:CAL

[GB12] Rajarshi Guhaniyogi and Sudipto Banerjee. Comment on article by Lum and Gelfand. Bayesian Analysis, 7(2):259–262, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2012/
vol07/issue02/guhaniyogi.pdf;   http://projecteuclid.
org/euclid.ba/1339878885. See [LG12b].

Griffin:2013:SPS

[GB13] Jim E. Griffin and Philip J. Brown. Some priors for sparse regression modelling. Bayesian Analysis, 8(3):691– 702, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/griffin.pdf;   http://
projecteuclid.org/euclid.ba/1378729924.

Griffin:2017:HSP

[GB17] Jim Griffin and Phil Brown. Hierarchical shrinkage priors for regression models. Bayesian Analysis, 12(1):135–159, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1453211963.

Gutierrez:2019:BNM

[GBGTR19] Luis Gutiérrez, Andrés F. Barrientos, Jorge González, and Daniel Taylor-Rodríguez. A Bayesian nonparametric multiple testing procedure for comparing several treatments against a control. Bayesian Analysis, 14(2):649–675, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1537258138.

Ghosh:2017:AOO

[GC17] Prasenjit Ghosh and Arijit Chakrabarti. Asymptotic optimality of one-group shrinkage priors in sparse high-dimensional problems. Bayesian Analysis, 12(4):1133–1161, December 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1475266758.

Gruber:2018:BMS

[GC18] Lutz F. Gruber and Claudia Czado. Bayesian model selection of regular vine copulas. Bayesian Analysis, 13(4):1111–1135, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1514516431.

Gorsky:2024:CMH

[GCM24] Shai Gorsky, Cliburn Chan, and Li Ma. Coarsened mixtures of hierarchical skew normal kernels for flow and mass cytometry analyses. Bayesian Analysis, 19(2):439–463, June 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/Coarsened-Mixtures-of-Hierarchical-
Skew-Normal-Kernels-for-Flow-and/10.1214/22-BA1356.full.

Gomez-Deniz:2009:SBC

[GD09] E. Gómez-Déniz. Some Bayesian credibility premiums obtained by using posterior regret Γ-minimax methodology. Bayesian Analysis, 4(2):223–242, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue02/deniz.pdf;   http:
//projecteuclid.org/euclid.ba/1340370276.

Gagnon:2020:NBA

[GDB20] Philippe Gagnon, Alain Desgagné, and Mylène Bédard. A new Bayesian approach to robustness against outliers in linear regression. Bayesian Analysis, 15(2):389–414, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1558598428.

Ghanta:2018:LMP

[GDNJ18] Sindhu Ghanta, Jennifer G. Dy, Donglin Niu, and Michael I. Jordan. Latent marked Poisson process with applications to object segmentation. Bayesian Analysis, 13(1):85–113, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1480129463.

Gelman:2006:PDV

[Gel06] Andrew Gelman. Prior distributions for variance parameters in hierarchical models (comment on article by Browne and Draper). Bayesian Analysis, 1(3):515–534, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue03/gelman.pdf;   http://projecteuclid.org/
euclid.ba/1340371048. See [BD06a].

Gelman:2008:OBS

[Gel08a] Andrew Gelman. Objections to Bayesian statistics. Bayesian Analysis, 3(3):445–449, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue03/gelman.pdf;
http://projecteuclid.org/euclid.ba/1340370429. See com-

ments [Ber08, Kad08, Sen08, Was08] and rejoinder [Gel08b].

Gelman:2008:R

[Gel08b] Andrew Gelman. Rejoinder. Bayesian Analysis, 3(3):467– 477, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2008/vol03/issue03/gelman_rejoinder.pdf;
 http://projecteuclid.org/euclid.ba/1340370434.  See

[Gel08a].

Gelman:2010:CAR

[Gel10] Andrew Gelman. Comment on article by Robert. Bayesian Analysis, 5(2):229–232, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue02/gelman.pdf;   http://
projecteuclid.org/euclid.ba/1340218336. See [Rob10].

Gutierrez:2019:BAS

[GGPM19] Luis Gutiérrez, Eduardo Gutiérrez-Peña, and Ramsés H. Mena. A Bayesian approach to statistical shape analysis via the projected normal distribution. Bayesian Analysis, 14(2):427–447, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1529719227.

Greengard:2023:FMP

[GHM^{+}23] Philip Greengard, Jeremy Hoskins, Charles C. Margossian, Jonah Gabry, Andrew Gelman, and Aki Vehtari. Fast methods for posterior inference of two-group normal-normal models. Bayesian Analysis, 18(3):889–907, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-3/Fast-Methods-for-Posterior-Inference-
of-Two-Group-Normal-Normal/10.1214/22-BA1329.full.

Gosling:2013:BLA

[GHO^{+}13] John Paul Gosling, Andy Hart, Helen Owen, Michael Davies, Jin Li, and Cameron MacKay. A Bayes linear approach to weightof-evidence risk assessment for skin allergy. Bayesian Analysis, 8(1):169–186, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue01/gosling.pdf;   http://
projecteuclid.org/euclid.ba/1362406656.

Ginebra:2007:MIS

[Gin07] Josep Ginebra. On the measure of the information in a statistical experiment. Bayesian Analysis, 2(1):167–211, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue01/ginebra.pdf;   http://projecteuclid.org/euclid.
ba/1340390067.

Greb:2014:RBE

[GKMvCT14] Friederike Greb, Tatyana Krivobokova, Axel Munk, and Stephan von Cramon-Taubadel. Regularized Bayesian estimation of generalized threshold regression models. Bayesian Analysis, 9(1):171–196, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue01/greb.pdf;   http:/
/projecteuclid.org/euclid.ba/1393251775.

Gao:2021:IMR

[GKSG21] Yuxiang Gao, Lauren Kennedy, Daniel Simpson, and Andrew Gelman. Improving multilevel regression and poststratification with structured priors. Bayesian Analysis, 16(3):719–744, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-3/Improving-Multilevel-Regression-and-Poststratification-
with-Structured-Priors/10.1214/20-BA1223.full.

Gramacy:2016:CAPa

[GL16] Robert B. Gramacy and Herbert K. H. Lee. Comment on article by Page and Quintana. Bayesian Analysis, 11(1):299– 302, March 2016. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-11/issue-1/Comment-
on-Article-by-Page-and-Quintana/10.1214/15-BA971A.full.

See [PQ16b].

Grazian:2017:ABI

[GL17] Clara Grazian and Brunero Liseo. Approximate Bayesian inference in semiparametric copula models. Bayesian Analysis, 12 (4):991–1016, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1510110045.

Griffin:2018:MCU

[GL18] Jim Griffin and Fabrizio Leisen. Modelling and computation using NCoRM mixtures for density regression. Bayesian Analysis, 13(3):897–916, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1508983454.

Gu:2022:GOL

[GL22] Mengyang Gu and Hanmo Li. Gaussian orthogonal latent factor processes for large incomplete matrices of correlated data. Bayesian Analysis, 17(4):1219–1244, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Gaussian-Orthogonal-Latent-Factor-
Processes-for-Large-Incomplete-Matrices-of/10.1214/21-
BA1295.full.

Gramatica:2024:SIM

[GLC24] Marco Gramatica, Silvia Liverani, and Peter Congdon. Structure induced by a multiple membership transformation on the conditional autoregressive model. Bayesian Analysis, 19(3):725–749, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-3/Structure-Induced-by-a-Multiple-Membership-
Transformation-on-the-Conditional/10.1214/23-BA1370.
full.

Glickman:2009:CAJ

[Gli09] Mark E. Glickman. Comment on article by Jensen et al. Bayesian Analysis, 4(4):661–664, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2009/vol04/issue04/glickman.
pdf;   http://projecteuclid.org/euclid.ba/1340369817.

See [JMW09a].

Giordano:2023:ESS

[GLJB23] Ryan Giordano, Runjing Liu, Michael I. Jordan, and Tamara Broderick. Evaluating sensitivity to the stick-breaking prior in Bayesian nonparametrics (with discussion). Bayesian Analysis, 18(1):287–366, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-1/Evaluating-Sensitivity-to-the-Stick-Breaking-
Prior-in-Bayesian-Nonparametrics/10.1214/22-BA1309.full.

Ghosh:2018:UCP

[GLM18] Joyee Ghosh, Yingbo Li, and Robin Mitra. On the use of Cauchy prior distributions for Bayesian logistic regression. Bayesian Analysis, 13(2):359–383, June 2018. CODEN ????

ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1488855634.

Gormley:2009:GMM

[GM09] Isobel Claire Gormley and Thomas Brendan Murphy. A grade of membership model for rank data. Bayesian Analysis, 4 (2):265–295, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2009/vol04/issue02/gormley.pdf;   http://
projecteuclid.org/euclid.ba/1340370278.

Geyer:2013:ACD

[GM13a] Charles Geyer and Glen Meeden. Asymptotics for constrained Dirichlet distributions. Bayesian Analysis, 8(1):89–110, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/geyer.pdf;   http://projecteuclid.org/euclid.
ba/1362406653.

Girolami:2013:CAS

[GM13b] Mark Girolami and Antonietta Mira. Comment on article by Schmidl et al. Bayesian Analysis, 8(1):27–32, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/girolami.pdf;   http://projecteuclid.org/euclid.

ba/1362406649. See [SCHT13b].

Goncalves:2016:MMR

[GM16] Kelly C. M. Gonçalves and Fernando A. S. Moura. A mixture model for rare and clustered populations under adaptive cluster sampling. Bayesian Analysis, 11(2):519–544, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1435585113.

Goncalves:2020:DQL

[GMB20] Kelly C. M. Gonçalves, Hélio S. Migon, and Leonardo S. Bastos. Dynamic quantile linear models: A Bayesian approach. Bayesian Analysis, 15(2):335–362, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1556244057.

Gonzalez:2021:REC

[GMdPV21] Miguel González, Carmen Minuesa, Inés del Puerto, and Anand N. Vidyashankar. Robust estimation in controlled branching processes: Bayesian estimators via disparities. Bayesian Analysis, 16(3):1009–1037, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-3/Robust-Estimation-in-Controlled-Branching-
Processes--Bayesian-Estimators-via/10.1214/20-BA1239.
full.

Ghosh:2021:BEC

[GMP21] Riddhi Pratim Ghosh, Bani Mallick, and Mohsen Pourahmadi. Bayesian estimation of correlation matrices of longitudinal data. Bayesian Analysis, 16(3):1039–1058, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-3/Bayesian-Estimation-of-Correlation-
Matrices-of-Longitudinal-Data/10.1214/20-BA1237.full.

Grazian:2015:CAD

[GMR15] C. Grazian, I. Masiani, and C. P. Robert. Comment on article by Dawid and Musio. Bayesian Analysis, 10(2):511–515, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1427806911. See

[DM15a].

Graham:2016:ABI

[GMS16] Daniel J. Graham, Emma J. McCoy, and David A. Stephens. Approximate Bayesian inference for doubly robust estimation. Bayesian Analysis, 11(1):47–69, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1423083639.

George:2021:OSE

[GMY21] Edward George, Gourab Mukherjee, and Keisuke Yano. Optimal shrinkage estimation of predictive densities under α-divergences. Bayesian Analysis, 16(4):1139–1155, December 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-4/Optimal-
Shrinkage-Estimation-of-Predictive-Densities-Under-%
ce%b1-Divergences/10.1214/21-BA1264.full.

Goldstein:2006:SBA

[Gol06a] Michael Goldstein. Subjective Bayesian analysis: Principles and practice. Bayesian Analysis, 1(3):403–420, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue03/goldstein.pdf;   http://projecteuclid.

org/euclid.ba/1340371036. See comments [Chr06, Dra06, Fie06a, Kad06, Kas06, Lad06, O’H06, Was06] and rejoinder [Ber06b].

Goldstein:2006:SOB

[Gol06b] Michael Goldstein. Subjectivity and objectivity in Bayesian statistics: rejoinder to the discussion. Bayesian Analysis, 1(3):465–472, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2006/vol01/issue03/goldstein_
rejoinder.pdf; http://projecteuclid.org/euclid.ba/1340371046.

Gosling:2007:NEH

[GOO07] John Paul Gosling, Jeremy E. Oakley, and Anthony O’Hagan. Nonparametric elicitation for heavy-tailed prior distributions. Bayesian Analysis, 2(4):693–718, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2007/vol02/issue04/gosling.
pdf;   http://projecteuclid.org/euclid.ba/1340370711.

Goplerud:2022:FAE

[Gop22] Max Goplerud. Fast and accurate estimation of non-nested binomial hierarchical models using variational inference. Bayesian Analysis, 17(2):623–650, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/Fast-and-Accurate-Estimation-of-Non-Nested-
Binomial-Hierarchical-Models/10.1214/21-BA1266.full.

Gosling:2012:CAA

[Gos12] John Paul Gosling. Comment on article by Albert et al. Bayesian Analysis, 7(3):537–540, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue03/gosling.
pdf;   http://projecteuclid.org/euclid.ba/1346158773.

See [ADGJ^{+}12a].

Gramacy:2010:SRM

[GP10] Robert B. Gramacy and Ester Pantaleo. Shrinkage regression for multivariate inference with missing data, and an application to portfolio balancing. Bayesian Analysis, 5 (2):237–262, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue02/gramacy.pdf;   http://
projecteuclid.org/euclid.ba/1340218338.

Gramacy:2012:SBR

[GP12] Robert B. Gramacy and Nicholas G. Polson. Simulationbased regularized logistic regression. Bayesian Analysis, 7(3): 567–590, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue03/gramacy.pdf;   http://
projecteuclid.org/euclid.ba/1346158776.

Goudie:2019:JSM

[GPL^{+}19] Robert J. B. Goudie, Anne M. Presanis, David Lunn, Daniela De Angelis, and Lorenz Wernisch. Joining and splitting models with Markov melding. Bayesian Analysis, 14(1):81–109, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1523671251.

Gaetan:2016:CAP

[GPP16] Carlo Gaetan, Simone A. Padoan, and Igor Prünster. Comment on article by Page and Quintana. Bayesian Analysis, 11 (1):307–314, March 2016. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-11/issue-1/Comment-
on-Article-by-Page-and-Quintana/10.1214/16-BA971C.full.

See [PQ16b] and rejoinder [PQ16a].

Guhaniyogi:2020:JML

[GR20] Rajarshi Guhaniyogi and Abel Rodriguez. Joint modeling of longitudinal relational data and exogenous variables. Bayesian Analysis, 15(2):477–503, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1559959375.

Guha:2023:HDB

[GR23] Sharmistha Guha and Abel Rodriguez. High-dimensional Bayesian network classification with network global-local shrinkage priors. Bayesian Analysis, 18(4):1131–1160, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-4/High-Dimensional-
Bayesian-Network-Classification-with-Network-Global-
Local-Shrinkage/10.1214/23-BA1378.full.

Gramacy:2016:CAPb

[Gra16] Robert B. Gramacy. Comment on article by Pratola. Bayesian Analysis, 11(3):913–919, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1472829060. See [Pra16a].

Griffin:2010:DPD

[Gri10] J. E. Griffin. Default priors for density estimation with mixture models. Bayesian Analysis, 5(1):45–64, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue01/griffin.
pdf;   http://projecteuclid.org/euclid.ba/1340369792.

Grelaud:2009:ALF

[GRM^{+}09] Aude Grelaud, Christian P. Robert, Jean-Michel Marin, François Rodolphe, and Jean-François Taly. ABC likelihood-free methods for model choice in Gibbs random fields. Bayesian Analysis, 4(2):317–335, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2009/vol04/issue02/robert.pdf;
http://projecteuclid.org/euclid.ba/1340370280.

Guha:2022:BCI

[GRM22] Sharmistha Guha, Jerome P. Reiter, and Andrea Mercatanti. Bayesian causal inference with bipartite record linkage. Bayesian Analysis, 17(4):1275–1299, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-4/Bayesian-Causal-Inference-with-Bipartite-
Record-Linkage/10.1214/21-BA1297.full.

Guhaniyogi:2021:BTR

[GS21] Rajarshi Guhaniyogi and Daniel Spencer. Bayesian tensor response regression with an application to brain activation studies. Bayesian Analysis, 16(4):1221–1249, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-4/Bayesian-Tensor-
Response-Regression-with-an-Application-to-Brain-Activation/
10.1214/21-BA1280.full.

Gelfand:2006:ESD

[GSW^{+}06a] Alan E. Gelfand, John A. Silander, Jr., Shanshan Wu, Andrew Latimer, Paul O. Lewis, Anthony G. Rebelo, and Mark Holder. Explaining species distribution patterns through hierarchical modeling. Bayesian Analysis, 1(1):41–92, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue01/gelfand.
pdf;   http://projecteuclid.org/euclid.ba/1340371072.

See comments [Hoe06, Ver06].

Gelfand:2006:R

[GSW^{+}06b] Alan E. Gelfand, John A. Silander, Jr., Shanshan Wu, Andrew Latimer, Paul O. Lewis, Anthony G. Rebelo, and Mark Holder. Rejoinder. Bayesian Analysis, 1(1):103–104, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue01/rejoinder.pdf;   http://projecteuclid.
org/euclid.ba/1340371075.

Gruber:2019:EUP

[GSWF19] Lutz F. Gruber, Erica F. Stuber, Lyndsie S. Wszola, and Joseph J. Fontaine. Estimating the use of public lands: Integrated modeling of open populations with convolution likelihood ecological abundance regression. Bayesian Analysis, 14 (4):1173–1199, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1551949261.

Ghosh:2016:APB

[GTGC16] Prasenjit Ghosh, Xueying Tang, Malay Ghosh, and Arijit Chakrabarti. Asymptotic properties of Bayes risk of a general class of shrinkage priors in multiple hypothesis testing under sparsity. Bayesian Analysis, 11(3):753–796, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1442364340.

Glynn:2019:BAD

[GTHB19] Chris Glynn, Surya T. Tokdar, Brian Howard, and David L. Banks. Bayesian analysis of dynamic linear topic models. Bayesian Analysis, 14(1):53–80, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1523671249.

Gu:2019:JRP

[Gu19] Mengyang Gu. Jointly robust prior for Gaussian stochastic process in emulation, calibration and variable selection. Bayesian Analysis, 14(3):857–885, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1560240031.

Grunwald:2017:IBI

[GvO17] Peter Grünwald and Thijs van Ommen. Inconsistency of Bayesian inference for misspecified linear models, and a proposal for repairing it. Bayesian Analysis, 12(4):1069–1103, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1510974325.

Gruber:2016:GAB

[GW16] Lutz Gruber and Mike West. GPU-accelerated Bayesian learning and forecasting in simultaneous graphical dynamic linear models. Bayesian Analysis, 11(1):125–149, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1425304898.

Hadj-Amar:2023:BAH

[HAJF23] Beniamino Hadj-Amar, Jack Jewson, and Mark Fiecas. Bayesian approximations to hidden semi-Markov models for telemetric monitoring of physical activity. Bayesian Analysis, 18(2):547– 577, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-2/Bayesian-Approximations-
to-Hidden-Semi-Markov-Models-for-Telemetric-Monitoring/
10.1214/22-BA1318.full.

Hanson:2006:MCL

[Han06] Timothy E. Hanson. Modeling censored lifetime data using a mixture of gammas baseline. Bayesian Analysis, 1(3):

575–594, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue03/hanson.pdf;   http://
projecteuclid.org/euclid.ba/1340371053.

Hans:2011:CAP

[Han11] Chris Hans. Comment on article by Polson and Scott. Bayesian Analysis, 6(1):37–41, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue01/hans.pdf;   http:/
/projecteuclid.org/euclid.ba/1339611939. See [PS11a].

Hans:2016:CAP

[Han16] Christopher M. Hans. Comment on article by Pratola. Bayesian Analysis, 11(3):921–927, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1472829061. See [Pra16a].

Hanson:2014:IPL

[HBJ14] Timothy E. Hanson, Adam J. Branscum, and Wesley O. Johnson. Informative g-priors for logistic regression. Bayesian Analysis, 9 (3):597–612, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1409921107.

Hart:2017:NGF

[HC17] Jeffrey D. Hart and Taeryon Choi. Nonparametric goodness of fit via cross-validation Bayes factors. Bayesian Analysis, 12 (3):653–677, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1471454532.

Holmes:2015:TSB

[HCGS15] Chris C. Holmes, François Caron, Jim E. Griffin, and David A. Stephens. Two-sample Bayesian nonparametric hypothesis testing. Bayesian Analysis, 10(2):297–320, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884976.

House:2006:BID

[HCH06] Leanna L. House, Merlise A. Clyde, and Yuh-Chin T. Huang. Bayesian identification of differential gene expression induced by metals in human bronchial epithelial cells. Bayesian Analysis, 1(1):105–120, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue01/house.pdf;   http:
//projecteuclid.org/euclid.ba/1340371076.

Hahn:2018:RCL

[HCPH18] P. Richard Hahn, Carlos M. Carvalho, David Puelz, and Jingyu He. Regularization and confounding in linear regression for treatment effect estimation. Bayesian Analysis, 13(1):163–182, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1484103680.

Huang:2012:SBI

[HD12] Yangxin Huang and Getachew A. Dagne. Simultaneous Bayesian inference for skew-normal semiparametric nonlinear mixedeffects models with covariate measurement errors. Bayesian Analysis, 7(1):189–210, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue01/yhuang.pdf;
http://projecteuclid.org/euclid.ba/1339616729.

Hendriksen:2021:OSB

[HdHG21] Allard Hendriksen, Rianne de Heide, and Peter Grünwald. Optional stopping with Bayes factors: a categorization and extension of folklore results, with an application to invariant situations. Bayesian Analysis, 16(3):961–989, September 2021. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-3/Optional-Stopping-with-Bayes-Factors-
-A-Categorization-and-Extension/10.1214/20-BA1234.full.

Hengartner:2010:CAH

[Hen10] Nick Hengartner. Comment on article by Hogg et al. Bayesian Analysis, 5(1):35–37, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2010/vol05/issue01/hengartner.pdf;   http://
projecteuclid.org/euclid.ba/1340369789. See [HKLM10a].

Higdon:2008:CAS

[HG08] Dave Higdon and James Gattiker. Comment on article by Sansó et al. [MR2383247]. Bayesian Analysis, 3(1):39–44, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue01/higdon.pdf;   http://projecteuclid.org/
euclid.ba/1340370558. See [SFZ08a].

Hu:2023:BSH

[HGXS23] Guanyu Hu, Junxian Geng, Yishu Xue, and Huiyan Sang. Bayesian spatial homogeneity pursuit of functional data: An application to the U.S. income distribution. Bayesian Analysis, 18(2):579–605, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Bayesian-Spatial-Homogeneity-Pursuit-of-Functional-
Data--An-Application/10.1214/22-BA1320.full.

Holmes:2006:BAV

[HH06] Chris C. Holmes and Leonhard Held. Bayesian auxiliary variable models for binary and multinomial regression. Bayesian Analysis, 1(1):145–168, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue01/held.pdf;   http:/
/projecteuclid.org/euclid.ba/1340371078. See comments

[vdL11a] and response [HH11].

Holmes:2011:RVL

[HH11] Chris Holmes and Leonhard Held. Response to van der Lans. Bayesian Analysis, 6(2):357–358, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue02/holmes.
pdf;   http://projecteuclid.org/euclid.ba/1339612051.

See [HH06].

He:2007:RCV

[HHC07] Yi He, James S. Hodges, and Bradley P. Carlin. Re-considering the variance parameterization in multiple precision models. Bayesian Analysis, 2(3):529–556, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue03/he.pdf;
http://projecteuclid.org/euclid.ba/1340370726.

Hui:2008:MBA

[HHG08] Sam K. Hui, Yanliu Huang, and Edward I. George. Modelbased analysis of concept maps. Bayesian Analysis, 3(3):

479–512, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue03/hui.pdf;   http://
projecteuclid.org/euclid.ba/1340370435.

Higson:2018:SEN

[HHHL18] Edward Higson, Will Handley, Mike Hobson, and Anthony Lasenby. Sampling errors in nested sampling parameter estimation. Bayesian Analysis, 13(3):873–896, September 2018. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1508897094.

Hallgren:2024:CDG

[HHT24] Karl L. Hallgren, Nicholas A. Heard, and Melissa J. M. Turcotte. Changepoint detection on a graph of time series. Bayesian Analysis, 19(2):649–676, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-2/Changepoint-Detection-on-a-Graph-of-Time-
Series/10.1214/23-BA1365.full.

Hamura:2022:GLS

[HIS22] Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. On global-local shrinkage priors for count data. Bayesian Analysis, 17(2):545–564, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-17/issue-2/On-
Global-Local-Shrinkage-Priors-for-Count-Data/10.1214/
21-BA1263.full.

Hanson:2012:BST

[HJZ12] Timothy E. Hanson, Alejandro Jara, and Luping Zhao. A Bayesian semiparametric temporally-stratified proportional hazards model with spatial frailties. Bayesian Analysis, 7(1): 147–188, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue01/hanson.pdf;   http://
projecteuclid.org/euclid.ba/1339616728.

Henderson:2018:CTT

[HK18] Daniel A. Henderson and Liam J. Kirrane. A comparison of truncated and time-weighted Plackett–Luce models for probabilistic forecasting of formula one results. Bayesian Analysis, 13(2):335– 358, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1488250819.

Heiner:2022:BND

[HK22] Matthew Heiner and Athanasios Kottas. Bayesian nonparametric density autoregression with lag selection. Bayesian Analysis, 17(4):1245–1273, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-4/Bayesian-Nonparametric-Density-Autoregression-
with-Lag-Selection/10.1214/21-BA1296.full.

Hogg:2010:EAS

[HKLM10a] Charles R. Hogg, Joseph B. Kadane, Jong Soo Lee, and Sara A. Majetich. Error analysis for small angle neutron scattering datasets using Bayesian inference. Bayesian Analysis, 5(1):1–33, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue01/hogg.pdf;   http:/
/projecteuclid.org/euclid.ba/1340369788. See comments

[Hen10, SS10] and rejoinder [HKLM10b].

Hogg:2010:R

[HKLM10b] Charles R. Hogg, Joseph B. Kadane, Jong Soo Lee, and Sara A. Majetich. Rejoinder. Bayesian Analysis, 5(1): 41–43, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue01/hogg_rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1340369791. See [HKLM10a].

Huang:2020:MMS

[HLC20] Weihong Huang, Yan Liu, and Yuguo Chen. Mixed membership stochastic blockmodels for heterogeneous networks. Bayesian Analysis, 15(3):711–736, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Mixed-Membership-Stochastic-Blockmodels-for-
Heterogeneous-Networks/10.1214/19-BA1163.full.

Huggins:2023:RMS

[HM23] Jonathan H. Huggins and Jeffrey W. Miller. Reproducible model selection using bagged posteriors. Bayesian Analysis, 18(1):79–104, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-1/Reproducible-Model-Selection-Using-Bagged-
Posteriors/10.1214/21-BA1301.full.

Holan:2009:BAE

[HMC09] Scott Holan, Tucker McElroy, and Sounak Chakraborty. A Bayesian approach to estimating the long memory parameter. Bayesian Analysis, 4(1):159–190, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2009/vol04/issue01/holan.pdf;
 http://projecteuclid.org/euclid.ba/1340370394.

Hahn:2020:BRT

[HMC20] P. Richard Hahn, Jared S. Murray, and Carlos M. Carvalho. Bayesian regression tree models for causal inference: Regularization, confounding, and heterogeneous effects (with discussion). Bayesian Analysis, 15(3):965–1056, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-15/issue-3/Bayesian-Regression-Tree-Models-for-
Causal-Inference--Regularization-Confounding/10.1214/
19-BA1195.full.

Hua:2022:PDT

[HMZ^{+}22] William Hua, Hongyuan Mei, Sarah Zohar, Magali Giral, and Yanxun Xu. Personalized dynamic treatment regimes in continuous time: a Bayesian approach for optimizing clinical decisions with timing. Bayesian Analysis, 17(3):849–878, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-3/Personalized-Dynamic-Treatment-Regimes-in-
Continuous-Time--A-Bayesian/10.1214/21-BA1276.full.

Hoeting:2006:SPM

[Hoe06] Jennifer A. Hoeting. Some perspectives on modeling species distributions (comment on article by Gelfand et al.). Bayesian Analysis, 1(1):93–97, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue01/hoeting.pdf;   http://
projecteuclid.org/euclid.ba/1340371073. See [GSW^{+}06a].

Hoff:2006:MBS

[Hof06] Peter D. Hoff. Model-based subspace clustering. Bayesian Analysis, 1(2):321–344, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2006/vol01/issue02/hoff321-344.pdf;   http://
projecteuclid.org/euclid.ba/1340371065.

Hoff:2011:RCA

[Hof11a] Peter D. Hoff. Rejoinder: “Comment on article by Hoff”. Bayesian Analysis, 6(2):205–207, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2011/vol06/issue02/hoff_rejoinder.
pdf;   http://projecteuclid.org/euclid.ba/1339612043.

See [Hof11b].

Hoff:2011:SCA

[Hof11b] Peter D. Hoff. Separable covariance arrays via the Tucker product, with applications to multivariate relational data. Bayesian Analysis, 6(2):179–196, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue02/hoff.pdf;   http:/
/projecteuclid.org/euclid.ba/1339612040. See comments

[All11, Fre11] and rejoinder [Hof11a].

Hoff:2013:CAM

[Hof13] Peter D. Hoff. Comment on article by Müller and Mitra. Bayesian Analysis, 8(2):311–318, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2013/vol08/issue02/hoff.pdf;
 http://projecteuclid.org/euclid.ba/1369407552.  See

[MM13a].

Hoff:2016:ESF

[Hof16] Peter D. Hoff. Equivariant and scale-free Tucker decomposition models. Bayesian Analysis, 11(3):627–648, September 2016. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1423576317.

Hamura:2024:SBI

[HOHS24] Yasuyuki Hamura, Takahiro Onizuka, Shintaro Hashimoto, and Shonosuke Sugasawa. Sparse Bayesian inference on gammadistributed observations using shape-scale inverse-gamma mixtures. Bayesian Analysis, 19(1):77–97, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-1/Sparse-Bayesian-Inference-on-Gamma-
Distributed-Observations-Using-Shape-Scale/10.1214/22-
BA1348.full.

Hooper:2008:EDT

[Hoo08] Peter M. Hooper. Exact distribution theory for belief net responses. Bayesian Analysis, 3(3):615–624, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue03/hooper.pdf;   http://projecteuclid.org/
euclid.ba/1340370440.

Haslett:2008:CAB

[HP08] John Haslett and Andrew Parnell. Comment on article by Blackwell and Buck. Bayesian Analysis, 3(2):249–254, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue02/haslett.pdf;   http://projecteuclid.org/euclid.
ba/1340370545. See [BB08a].

Hans:2015:CAD

[HP15] Christopher M. Hans and Mario Peruggia. Comment on article by Dawid and Musio. Bayesian Analysis, 10(2):505–509, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1427806910. See [DM15a].

Heaukulani:2020:GTI

[HR20] Creighton Heaukulani and Daniel M. Roy. Gibbs-type Indian buffet processes. Bayesian Analysis, 15(3):683–710, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-15/issue-3/Gibbs-type-Indian-
Buffet-Processes/10.1214/19-BA1166.full.

Hammer:2024:ABI

[HRT24] Hugo L. Hammer, Michael A. Riegler, and Håkon Tjelmeland. Approximate Bayesian inference based on expected evaluation. Bayesian Analysis, 19(3):677–698, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-3/Approximate-Bayesian-Inference-Based-
on-Expected-Evaluation/10.1214/23-BA1368.full.

Hu:2018:DPM

[HRW18] Jingchen Hu, Jerome P. Reiter, and Quanli Wang. Dirichlet process mixture models for modeling and generating synthetic versions of nested categorical data. Bayesian Analysis, 13(1):183– 200, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1485227030.

Hahn:2009:PEC

[HS09] Markus Hahn and Jörn Sassy. Parameter estimation in continuous time Markov switching models: a semi-continuous Markov chain Monte Carlo approach. Bayesian Analysis, 4(1):63–84, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2009/vol04/issue01/hahn.pdf;   http://projecteuclid.
org/euclid.ba/1340370390.

Hernandez-Stumpfhauser:2017:GPN

[HSBvdW17] Daniel Hernandez-Stumpfhauser, F. Jay Breidt, and Mark J. van der Woerd. The general projected normal distribution of arbitrary dimension: Modeling and Bayesian inference. Bayesian Analysis, 12(1):113–133, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1453211962.

Hobbs:2012:CPI

[HSC12] Brian P. Hobbs, Daniel J. Sargent, and Bradley P. Carlin. Commensurate priors for incorporating historical information in clinical trials using general and generalized linear models. Bayesian Analysis, 7(3):639–674, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2012/vol07/issue03/hobbs.pdf;
 http://projecteuclid.org/euclid.ba/1346158779.

Hubin:2020:NAA

[HSF20] Aliaksandr Hubin, Geir Storvik, and Florian Frommlet. A novel algorithmic approach to Bayesian logic regression (with discussion). Bayesian Analysis, 15(1):263–333, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1545296448.

He:2021:OBA

[HSH21a] Daojiang He, Dongchu Sun, and Lei He. Objective Bayesian analysis for the Student-t linear regression. Bayesian Analysis, 16(1):129–145, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/Objective-Bayesian-Analysis-for-the-Student-
-t-Linear-Regression/10.1214/20-BA1198.full.

Hrafnkelsson:2021:MST

[HSH^{+}21b] Birgir Hrafnkelsson, Stefan Siegert, Raphaël Huser, Haakon Bakka, and Árni V. Jóhannesson. Max-and-smooth: a two-step approach for approximate Bayesian inference in latent Gaussian models. Bayesian Analysis, 16(2):611–638, June 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-2/Max-and-
Smooth--A-Two-Step-Approach-for-Approximate/10.1214/
20-BA1219.full.

Hutter:2007:EBR

[Hut07] Marcus Hutter. Exact Bayesian regression of piecewise constant functions. Bayesian Analysis, 2(4):635–664, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2007/
vol02/issue04/hutter.pdf;   http://projecteuclid.org/
euclid.ba/1340370708.

Horrocks:2009:PPJ

[HvDH09] Julie Horrocks and Marianne J. van Den Heuvel. Prediction of pregnancy: a joint model for longitudinal and binary data. Bayesian Analysis, 4(3):523–538, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2009/vol04/issue03/horrocks.
pdf;   http://projecteuclid.org/euclid.ba/1340369853.

Huang:2013:SMN

[HW13] Alan Huang and M. P. Wand. Simple marginally noninformative prior distributions for covariance matrices. Bayesian Analysis, 8(2):439–452, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue02/huang.pdf;   http:
//projecteuclid.org/euclid.ba/1369407559.

Heinecke:2021:BDQ

[HYDE21] Andreas Heinecke, Lifeng Ye, Maria De Iorio, and Timothy Ebbels. Bayesian deconvolution and quantification of metabolites from J-resolved NMR spectroscopy. Bayesian Analysis, 16(2):425–458, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/Bayesian-Deconvolution-and-Quantification-
of-Metabolites-from-J-Resolved-NMR/10.1214/20-BA1208.
full.

Huo:2012:BDF

[HYY12] Lin Huo, Ying Yuan, and Guosheng Yin. Bayesian dose finding for combined drugs with discrete and continuous doses. Bayesian Analysis, 7(4):1035–1052, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue04/huo.pdf;
 http://projecteuclid.org/euclid.ba/1354024471.

Hu:2022:FCL

[HZ22] Yaozhong Hu and Junxi Zhang. Functional central limit theorems for stick-breaking priors. Bayesian Analysis, 17 (4):1101–1120, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-4/Functional-Central-Limit-Theorems-for-Stick-
Breaking-Priors/10.1214/21-BA1290.full.

Irie:2019:BEM

[IW19] Kaoru Irie and Mike West. Bayesian emulation for multi-step optimization in decision problems. Bayesian Analysis, 14(1):137– 160, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1524103230.

Johndrow:2018:OGA

[JB18] James Johndrow and Anirban Bhattacharya. Optimal Gaussian approximations to the posterior for log-linear models with Diaconis–Ylvisaker priors. Bayesian Analysis, 13(1):201–223, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1487646097.

Jiang:2008:BMC

[JD08] Thomas J. Jiang and James M. Dickey. Bayesian methods for categorical data under informative censoring. Bayesian Analysis, 3(3):541–553, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue03/jiang.pdf;   http:
//projecteuclid.org/euclid.ba/1340370437.

Jarvenpaa:2019:EAR

[JGP^{+}19] Marko Järvenpää, Michael U. Gutmann, Arijus Pleska, Aki Vehtari, and Pekka Marttinen. Efficient acquisition rules for modelbased approximate Bayesian computation. Bayesian Analysis, 14(2):595–622, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1537258134.

Jarvenpaa:2021:PGP

[JGVM21] Marko Järvenpää, Michael U. Gutmann, Aki Vehtari, and Pekka Marttinen. Parallel Gaussian process surrogate Bayesian inference with noisy likelihood evaluations. Bayesian Analysis, 16(1):147–178, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/Parallel-Gaussian-Process-Surrogate-Bayesian-
Inference-with-Noisy-Likelihood-Evaluations/10.1214/
20-BA1200.full.

Johnson:2022:BIE

[JHB22] Stephen R. Johnson, Daniel A. Henderson, and Richard J. Boys. On Bayesian inference for the extended Plackett–Luce model. Bayesian Analysis, 17(2):465–490, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/On-Bayesian-inference-for-the-Extended-Plackett-
Luce-model/10.1214/21-BA1258.full.

Jewell:2009:BAE

[JKNR09] Chris P. Jewell, Theodore Kypraios, Peter Neal, and Gareth O. Roberts. Bayesian analysis for emerging infectious diseases. Bayesian Analysis, 4(3):465–496, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue03/jewell.
pdf;   http://projecteuclid.org/euclid.ba/1340369851.

Jiang:2019:BOP

[JL19] Wenxin Jiang and Cheng Li. On Bayesian oracle properties. Bayesian Analysis, 14(1):235–260, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1526457619.

Jo:2017:DSS

[JLM^{+}17] Seongil Jo, Jaeyong Lee, Peter Müller, Fernando A. Quintana, and Lorenzo Trippa. Dependent species sampling models for spatial density estimation. Bayesian Analysis, 12(2):379–406, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1462297334.

Ji:2009:SMM

[JMKW09] Chunlin Ji, Daniel Merl, Thomas B. Kepler, and Mike West. Spatial mixture modelling for unobserved point processes: examples in immunofluorescence histology. Bayesian Analysis, 4(2):297–315, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue02/ji.pdf;   http://
projecteuclid.org/euclid.ba/1340370279.

Jensen:2009:HBM

[JMW09a] Shane T. Jensen, Blakeley B. McShane, and Abraham J. Wyner. Hierarchical Bayesian modeling of hitting performance in baseball. Bayesian Analysis, 4(4):631–652, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2009/
vol04/issue04/jensen.pdf;   http://projecteuclid.org/

euclid.ba/1340369815. See comments [AB09, Gli09, QM09] and rejoinder [JMW09b].

Jensen:2009:R

[JMW09b] Shane T. Jensen, Blakeley B. McShane, and Abraham J. Wyner. Rejoinder. Bayesian Analysis, 4(4):669–674, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue04/jensen_rejoinder.pdf;   http://projecteuclid.
org/euclid.ba/1340369819. See [JMW09a].

Jain:2007:R

[JN07a] Sonia Jain and Radford M. Neal. Rejoinder. Bayesian Analysis, 2(3):495–500, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue03/rejoinder.pdf;   http:
//projecteuclid.org/euclid.ba/1340370724. See [JN07b]

[MR2342168].

Jain:2007:SMC

[JN07b] Sonia Jain and Radford M. Neal. Splitting and merging components of a nonconjugate Dirichlet process mixture model. Bayesian Analysis, 2(3):445–472, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue03/jain.pdf;
 http://projecteuclid.org/euclid.ba/1340370720.  See

comments [Dah07, Rob07, Mac07] and rejoinder [JN07a]..

Jara:2013:TSM

[JNBQ13] A. Jara, L. E. Nieto-Barajas, and F. Quintana. A time series model for responses on the unit interval. Bayesian Analysis, 8(3):723–740, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue03/jara.pdf;   http:/
/projecteuclid.org/euclid.ba/1378729926.

Johnson:2007:BMA

[Joh07] Valen E. Johnson. Bayesian model assessment using pivotal quantities. Bayesian Analysis, 2(4):719–733, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue04/johnson.pdf;   http://projecteuclid.org/euclid.
ba/1340370712.

Johnson:2013:NAB

[Joh13] Valen E. Johnson. On numerical aspects of Bayesian model selection in high and ultrahigh-dimensional settings. Bayesian Analysis, 8(4):741–758, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue04/johnson.pdf;   http://
projecteuclid.org/euclid.ba/1386166311.

Jeliazkov:2008:DSF

[JP08] Ivan Jeliazkov and Dale J. Poirier. Dynamic and structural features of intifada violence: a Markov process approach. Bayesian Analysis, 3(1):63–77, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2008/vol03/issue01/jeliazkov.pdf;   http:
//projecteuclid.org/euclid.ba/1340370561.

Jeong:2016:BSI

[JP16] Seonghyun Jeong and Taeyoung Park. Bayesian semiparametric inference on functional relationships in linear mixed models. Bayesian Analysis, 11(4):1137–1163, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1448852253.

Jha:2024:GSD

[JS24] Chetkar Jha and Dongchu Sun. A general scheme for deriving conditional reference priors. Bayesian Analysis, 19(1):153–179, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-1/A-General-Scheme-
for-Deriving-Conditional-Reference-Priors/10.1214/22-
BA1345.full.

Jones:2022:QOP

[JTC22] David E. Jones, Robert N. Trangucci, and Yang Chen. Quantifying observed prior impact. Bayesian Analysis, 17(3):737–764, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-3/Quantifying-Observed-Prior-Impact/10.1214/
21-BA1271.full.

Jackson:2023:EEC

[JV23] Samuel E. Jackson and Ian Vernon. Efficient emulation of computer models utilising multiple known boundaries of differing dimension. Bayesian Analysis, 18(1):165–191, March 2023. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-1/Efficient-Emulation-of-Computer-Models-
Utilising-Multiple-Known-Boundaries-of/10.1214/22-BA1304.
full.

Jiang:2017:BNT

[JYL17] Bo Jiang, Chao Ye, and Jun S. Liu. Bayesian nonparametric tests via sliced inverse modeling. Bayesian Analysis, 12(1):89– 112, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1453211961.

Kadane:2006:OBA

[Kad06] Joseph B. Kadane. Is “objective Bayesian analysis” objective, Bayesian, or wise? (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1(3):433–436, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue03/kadane.pdf;   http://projecteuclid.org/
euclid.ba/1340371040. See [Ber06a, Gol06a].

Kadane:2008:CAG

[Kad08] Joseph B. Kadane. Comment on article by Gelman. Bayesian Analysis, 3(3):455–457, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2008/vol03/issue03/kadane.
pdf;   http://projecteuclid.org/euclid.ba/1340370431.

See [Gel08a].

Kadane:2016:SPA

[Kad16] Joseph B. Kadane. Sums of possibly associated Bernoulli variables: The Conway–Maxwell-binomial distribution. Bayesian Analysis, 11(2):403–420, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1431607821.

Kim:2012:BME

[KAL12] Joungyoun Kim, Nicola M. Anthony, and Bret R. Larget. A Bayesian method for estimating evolutionary history. Bayesian Analysis, 7(4):917–974, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2012/vol07/issue04/kim.pdf;
http://projecteuclid.org/euclid.ba/1354024468.

Kass:2006:KBC

[Kas06] Robert E. Kass. Kinds of Bayesians (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1(3):437–440, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2006/vol01/issue03/kass.pdf;   http://projecteuclid.
org/euclid.ba/1340371041. See [Ber06a, Gol06a].

Katzfuss:2015:CAD

[KB15] Matthias Katzfuss and Anirban Bhattacharya. Comment on article by Dawid and Musio. Bayesian Analysis, 10(2):501– 504, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1428419112. See [DM15a].

Kowal:2023:SFF

[KC23] Daniel R. Kowal and Antonio Canale. Semiparametric functional factor models with Bayesian rank selection. Bayesian Analysis, 18(4):1161–1189, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Semiparametric-Functional-Factor-Models-with-
Bayesian-Rank-Selection/10.1214/23-BA1410.full.

Ko:2015:DPH

[KCG15] Stanley I. M. Ko, Terence T. L. Chong, and Pulak Ghosh. Dirichlet process hidden Markov multiple change-point model. Bayesian Analysis, 10(2):275–296, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884975.

Klein:2021:BES

[KCK^{+}21] Nadja Klein, Manuel Carlan, Thomas Kneib, Stefan Lang, and Helga Wagner. Bayesian effect selection in structured additive distributional regression models. Bayesian Analysis, 16(2):545– 573, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-2/Bayesian-Effect-
Selection-in-Structured-Additive-Distributional-Regression-
Models/10.1214/20-BA1214.full.

King:2019:BFF

[KCR19] Guillaume Kon Kam King, Antonio Canale, and Matteo Ruggiero. Bayesian functional forecasting with locally-autoregressive dependent processes. Bayesian Analysis, 14(4):1121–1141, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1545296447.

Kenobi:2012:BMU

[KD12] Kim Kenobi and Ian L. Dryden. Bayesian matching of unlabeled point sets using Procrustes and configuration models. Bayesian Analysis, 7(3):547–566, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue03/kenobi.pdf;
http://projecteuclid.org/euclid.ba/1346158775.

Kleinegesse:2021:SBE

[KDG21] Steven Kleinegesse, Christopher Drovandi, and Michael U. Gutmann. Sequential Bayesian experimental design for implicit models via mutual information. Bayesian Analysis, 16(3):773–802, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-3/Sequential-Bayesian-Experimental-Design-for-
Implicit-Models-via-Mutual-Information/10.1214/20-BA1225.
full.

Kim:2009:SDP

[KDV09] Sinae Kim, David B. Dahl, and Marina Vannucci. Spiked Dirichlet process prior for Bayesian multiple hypothesis testing in random effects models. Bayesian Analysis, 4(4): 707–732, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue04/kim.pdf;   http://
projecteuclid.org/euclid.ba/1340369821.

Kirch:2019:BWN

[KEMM19] Claudia Kirch, Matthew C. Edwards, Alexander Meier, and Renate Meyer. Beyond Whittle: Nonparametric correction of a parametric likelihood with a focus on Bayesian time series analysis. Bayesian Analysis, 14(4):1037–1073, December 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1540865702.

Keefe:2019:OBA

[KFF19] Matthew J. Keefe, Marco A. R. Ferreira, and Christopher T. Franck. Objective Bayesian analysis for Gaussian hierarchical models with intrinsic conditional autoregressive priors. Bayesian Analysis, 14(1):181–209, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1524124869.

Kyung:2009:BID

[KG09] Minjung Kyung and Sujit K. Ghosh. Bayesian inference for directional conditionally autoregressive models. Bayesian Analysis, 4(4):675–706, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue04/kyung.pdf;   http:
//projecteuclid.org/euclid.ba/1340369820.

Kyung:2010:PRS

[KGGC10] Minjung Kyung, Jeff Gill, Malay Ghosh, and George Casella. Penalized regression, standard errors, and Bayesian lassos. Bayesian Analysis, 5(2):369–411, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue02/casella.
pdf;   http://projecteuclid.org/euclid.ba/1340218343.

Kang:2024:FBF

[KJP24] Hyun Bin Kang, Yeo Jin Jung, and Jaewoo Park. Fast Bayesian functional regression for non-Gaussian spatial data. Bayesian Analysis, 19(2):407–438, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-2/Fast-Bayesian-Functional-Regression-for-Non-
Gaussian-Spatial-Data/10.1214/22-BA1354.full.

Kokolakis:2007:MRP

[KK07] George Kokolakis and George Kouvaras. On the multimodality of random probability measures. Bayesian Analysis, 2 (1):213–219, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue01/kokolakis.pdf;   http:
//projecteuclid.org/euclid.ba/1340390068.

Klein:2016:SDP

[KK16] Nadja Klein and Thomas Kneib. Scale-dependent priors for variance parameters in structured additive distributional regression. Bayesian Analysis, 11(4):1071–1106, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1448323525.

Kidd:2022:BNN

[KK22] Brian Kidd and Matthias Katzfuss. Bayesian nonstationary and nonparametric covariance estimation for large spatial data (with discussion). Bayesian Analysis, 17(1):291–351, March 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-1/Bayesian-Nonstationary-and-Nonparametric-
Covariance-Estimation-for-Large-Spatial-Data/10.1214/
21-BA1273.full.

Kadane:2014:TRS

[KM14] Joseph B. Kadane and Steven N. MacEachern. Toward rational social decisions: A review and some results. Bayesian Analysis, 9(3):685–698, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1409921110.

Kundu:2019:EBR

[KMB19] Suprateek Kundu, Bani K. Mallick, and Veera Baladandayuthapani. Efficient Bayesian regularization for graphical model selection. Bayesian Analysis, 14(2):449–476, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1531274648.

Kass:2006:DCP

[KN06] Robert E. Kass and Ranjini Natarajan. A default conjugate prior for variance components in generalized linear mixed models (comment on article by Browne and Draper). Bayesian Analysis, 1(3):535–542, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue03/kass2.pdf;   http:
//projecteuclid.org/euclid.ba/1340371049. See [BD06a].

Kobayashi:2017:BET

[Kob17] Genya Kobayashi. Bayesian endogenous tobit quantile regression. Bayesian Analysis, 12(1):161–191, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1455559718.

Komaki:2015:APB

[Kom15] Fumiyasu Komaki. Asymptotic properties of Bayesian predictive densities when the distributions of data and target variables are different. Bayesian Analysis, 10(1):31–51, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1422468422.

Koop:2011:CAW

[Koo11] Gary Koop. Comment on article by Wyse et al. Bayesian Analysis, 6(4):533–540, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue04/koop.pdf;   http:/
/projecteuclid.org/euclid.ba/1339616534. See [WFR11a].

Kowal:2021:DRM

[Kow21] Daniel R. Kowal. Dynamic regression models for timeordered functional data. Bayesian Analysis, 16(2):459–487, June 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-2/Dynamic-
Regression-Models-for-Time-Ordered-Functional-Data/10.
1214/20-BA1213.full.

Karmakar:2021:BMT

[KR21] Sayar Karmakar and Arkaprava Roy. Bayesian modelling of time-varying conditional heteroscedasticity. Bayesian Analysis, 16(4):1157–1185, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-4/Bayesian-Modelling-of-Time-Varying-Conditional-
Heteroscedasticity/10.1214/21-BA1267.full.

Kaufman:2010:BFA

[KS10a] Cari G. Kaufman and Stephan R. Sain. Bayesian functional ANOVA modeling using Gaussian process prior distributions. Bayesian Analysis, 5(1):123–149, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue01/kaufman.
pdf;   http://projecteuclid.org/euclid.ba/1340369795.

Kharroubi:2010:PSS

[KS10b] S. A. Kharroubi and T. J. Sweeting. Posterior simulation via the signed root log-likelihood ratio. Bayesian Analysis, 5(4): 787–815, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue04/kharroubi.pdf;   http:
//projecteuclid.org/euclid.ba/1340110855.

Klein:2019:ICB

[KS19] Nadja Klein and Michael Stanley Smith. Implicit copulas from Bayesian regularized regression smoothers. Bayesian Analysis, 14 (4):1143–1171, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1545296445.

Kim:2012:FBH

[KSLP12a] Sungduk Kim, Rajeshwari Sundaram, Germaine M. Buck Louis, and Cecilia Pyper. Flexible Bayesian human fecundity models. Bayesian Analysis, 7(4):771–800, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2012/vol07/issue04/kim2.pdf;
 http://projecteuclid.org/euclid.ba/1354024459.  See

comments [Sta12, Sca12].

Kim:2012:R

[KSLP12b] Sungduk Kim, Rajeshwari Sundaram, Germaine M. Buck Louis, and Cecilia Pyper. Rejoinder. Bayesian Analysis, 7(4): 809–812, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2012/vol07/issue04/kim_rejoinder.pdf;   http:
//projecteuclid.org/euclid.ba/1354024462.

Kadane:2006:CAC

[KSM^{+}06] Joseph B. Kadane, Galit Shmueli, Thomas P. Minka, Sharad Borle, and Peter Boatwright. Conjugate analysis of the Conway–Maxwell–Poisson distribution. Bayesian Analysis, 1 (2):363–374, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2006/vol01/issue02/kadane363-374.pdf;   http:
//projecteuclid.org/euclid.ba/1340371067. See correction

[KSM^{+}18].

Kadane:2018:NCC

[KSM^{+}18] Joseph B. Kadane, Galit Shmueli, Thomas P. Minka, Sharad Borle, and Peter Boatwright. Note of correction: “Conjugate Analysis of the Conway–Maxwell–Poisson Distribution”. Bayesian Analysis, 13(3):1009, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1532937627. See [KSM^{+}06].

Kyung:2011:CBM

[Kyu11] Minjung Kyung. A computational Bayesian method for estimating the number of knots in regression splines. Bayesian Analysis, 6(4):793–828, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue04/kyung.pdf;   http:
//projecteuclid.org/euclid.ba/1339616544.

Lad:2006:OBS

[Lad06] Frank Lad. Objective Bayesian statistics . . .. Do you buy it? Should we sell it? (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1(3):441–444, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue03/lad.pdf;   http://projecteuclid.org/euclid.ba/

1340371042. See [Ber06a, Gol06a].

Liverani:2009:EUB

[LAE^{+}09] Silvia Liverani, Paul E. Anderson, Kieron D. Edwards, Andrew J. Millar, and Jim Q. Smith. Efficient utility-based clustering over high dimensional partition spaces. Bayesian Analysis, 4 (3):539–571, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2009/vol04/issue03/liverani.pdf;   http:/
/projecteuclid.org/euclid.ba/1340369854.

Lambert:2006:CAB

[Lam06] Paul C. Lambert. (comment on article by Browne and Draper). Bayesian Analysis, 1(3):543–546, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue03/lambert.
pdf;   http://projecteuclid.org/euclid.ba/1340371050.

See [BD06a].

Liu:2009:MBA

[LBB09] F. Liu, M. J. Bayarri, and J. O. Berger. Modularization in Bayesian analysis, with emphasis on analysis of computer models. Bayesian Analysis, 4(1):119–150, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2009/vol04/issue01/liu.pdf;
http://projecteuclid.org/euclid.ba/1340370392.

Liu:2016:PSF

[LBBJ16] Zhuqing Liu, Veronica J. Berrocal, Andreas J. Bartsch, and Timothy D. Johnson. Pre-surgical fMRI data analysis using a spatially adaptive conditionally autoregressive model. Bayesian Analysis, 11(2):599–625, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1440594946.

Linero:2022:BST

[LBLS22] Antonio R. Linero, Piyali Basak, Yinpu Li, and Debajyoti Sinha. Bayesian survival tree ensembles with submodel shrinkage. Bayesian Analysis, 17(3):997–1020, September 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-3/Bayesian-Survival-Tree-Ensembles-with-
Submodel-Shrinkage/10.1214/21-BA1285.full.

Lin:2012:CAL

[LC12] Nan Lin and Chao Chang. Comment on article by Lum and Gelfand. Bayesian Analysis, 7(2):263–270, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue02/lin.pdf;
 http://projecteuclid.org/euclid.ba/1339878886.  See

[LG12b].

Le:2017:BIS

[LC17] Tri Le and Bertrand Clarke. A Bayes interpretation of stacking for M-complete and M-open settings. Bayesian Analysis, 12 (3):807–829, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1473276261.

Levi:2022:FOW

[LC22] Evgeny Levi and Radu V. Craiu. Finding our way in the dark: Approximate MCMC for approximate Bayesian methods. Bayesian Analysis, 17(1):193–221, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-1/Finding-our-Way-in-the-Dark--Approximate-
MCMC-for/10.1214/20-BA1250.full.

Loyal:2023:BNL

[LC23] Joshua Daniel Loyal and Yuguo Chen. A Bayesian nonparametric latent space approach to modeling evolving communities in dynamic networks. Bayesian Analysis, 18(1):49–77, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-1/A-Bayesian-Nonparametric-
Latent-Space-Approach-to-Modeling-Evolving-Communities/
10.1214/21-BA1300.full.

Liu:2014:BRG

[LCL^{+}14] Fei Liu, Sounak Chakraborty, Fan Li, Yan Liu, and Aurelie C. Lozano. Bayesian regularization via graph Laplacian. Bayesian Analysis, 9(2):449–474, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2014/vol09/issue02/sounak.pdf;
http://projecteuclid.org/euclid.ba/1401148316.

Liu:2014:DML

[LCS^{+}14] Ruitao Liu, Arijit Chakrabarti, Tapas Samanta, Jayanta K. Ghosh, and Malay Ghosh. On divergence measures leading to Jeffreys and other reference priors. Bayesian Analysis, 9(2):331–370, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue02/liu.pdf;   http://
projecteuclid.org/euclid.ba/1401148312.

Leonard:2011:EBL

[Leo11] David Leonard. Estimating a bivariate linear relationship. Bayesian Analysis, 6(4):727–754, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue04/leonard.
pdf;   http://projecteuclid.org/euclid.ba/1339616542.

Loh:2006:CAV

[LG06] Ji Meng Loh and Andrew Gelman. Comment on article by van Dyk et al. Bayesian Analysis, 1(2):237–240, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue02/loh237-240.pdf;   http://projecteuclid.
org/euclid.ba/1340371060. See [vDCE^{+}06].

Lum:2012:R

[LG12a] Kristian Lum and Alan E. Gelfand. Rejoinder. Bayesian Analysis, 7(2):273–276, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.edu/

journal/2012/vol07/issue02/lum_rejoinder.pdf;   http:
//projecteuclid.org/euclid.ba/1339878888. See [LG12b].

Lum:2012:SQM

[LG12b] Kristian Lum and Alan E. Gelfand. Spatial quantile multiple regression using the asymmetric Laplace process. Bayesian Analysis, 7(2):235–258, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2012/vol07/issue02/lum.pdf;   http://
projecteuclid.org/euclid.ba/1339878884. See comments

[GB12, LC12, Fer12] and rejoinder [LG12a].

Li:2014:BMS

[LG14] Meng Li and Subhashis Ghosal. Bayesian multiscale smoothing of Gaussian noised images. Bayesian Analysis, 9(3):733–758, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1409921112.

Leininger:2017:BIM

[LG17] Thomas J. Leininger and Alan E. Gelfand. Bayesian inference and model assessment for spatial point patterns using posterior predictive samples. Bayesian Analysis, 12(1):1–30, March 2017.

CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1448899901.

Liu:2024:GGW

[LG24] Yang Liu and Robert J. B. Goudie. Generalized geographically weighted regression model within a modularized Bayesian framework. Bayesian Analysis, 19(2):465–500, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/Generalized-Geographically-Weighted-
Regression-Model-within-a-Modularized-Bayesian-Framework/
10.1214/22-BA1357.full.

Lawrence:2010:CAV

[LH10] Earl Lawrence and David M. Higdon. Comment on article by Vernon et al. Bayesian Analysis, 5(4):683–689, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue04/lawrence.pdf;   http://projecteuclid.org/euclid.
ba/1340110849. See [VGB10a].

Lan:2020:FBD

[LHE^{+}20] Shiwei Lan, Andrew Holbrook, Gabriel A. Elias, Norbert J. Fortin, Hernando Ombao, and Babak Shahbaba. Flexible Bayesian dynamic modeling of correlation and covariance matrices. Bayesian Analysis, 15(4):1199–1228, December 2020. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-15/issue-4/Flexible-Bayesian-Dynamic-Modeling-
of-Correlation-and-Covariance-Matrices/10.1214/19-BA1173.
full.

Li:2009:CAM

[Li09] Hongzhe Li. Comment on article by Monni and Tadesse. Bayesian Analysis, 4(3):449–452, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://ba.stat.cmu.edu/journal/2009/vol04/issue03/li.
pdf;   http://projecteuclid.org/euclid.ba/1340369848.

See [MT09b].

Liang:2012:CAS

[Lia12] Feng Liang. Comment on article by Sancetta. Bayesian Analysis, 7(1):45–46, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2012/vol07/issue01/liang.pdf;   http:
//projecteuclid.org/euclid.ba/1339616723. See [San12b].

Lee:2014:SBV

[LJCB14] Kuo-Jung Lee, Galin L. Jones, Brian S. Caffo, and Susan S. Bassett. Spatial Bayesian variable selection models on functional magnetic resonance imaging time-series data. Bayesian Analysis, 9(3):699–732, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1409921111.

Leslie:2009:NED

[LKF09] David S. Leslie, Robert Kohn, and Denzil G. Fiebig. Nonparametric estimation of the distribution function in contingent valuation models. Bayesian Analysis, 4(3):573–597, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2009/
vol04/issue03/leslie.pdf;   http://projecteuclid.org/
euclid.ba/1340369855.

Letham:2019:CBO

[LKOB19] Benjamin Letham, Brian Karrer, Guilherme Ottoni, and Eytan Bakshy. Constrained Bayesian optimization with noisy experiments. Bayesian Analysis, 14(2):495–519, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1533866666.

Li:2010:BEN

[LL10] Qing Li and Nan Lin. The Bayesian elastic net. Bayesian Analysis, 5(1):151–170, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue01/lin.pdf;   http://
projecteuclid.org/euclid.ba/1340369796.

Lee:2018:OBM

[LL18] Kyoungjae Lee and Jaeyong Lee. Optimal Bayesian minimax rates for unconstrained large covariance matrices. Bayesian Analysis, 13(4):1215–1233, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1519355083.

Lee:2020:BBT

[LL20] Kyoungjae Lee and Lizhen Lin. Bayesian bandwidth test and selection for high-dimensional banded precision matrices. Bayesian Analysis, 15(3):737–758, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Bayesian-Bandwidth-Test-and-Selection-for-
High-dimensional-Banded-Precision/10.1214/19-BA1167.
full.

Lee:2023:SBH

[LL23] Kyoungjae Lee and Lizhen Lin. Scalable Bayesian highdimensional local dependence learning. Bayesian Analysis, 18(1):25–47, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-1/Scalable-Bayesian-High-dimensional-Local-
Dependence-Learning/10.1214/21-BA1299.full.

Lee:2023:PPP

[LLL23] Kwangmin Lee, Kyoungjae Lee, and Jaeyong Lee. Postprocessed posteriors for banded covariances. Bayesian Analysis, 18(3):1017–1040, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Post-Processed-Posteriors-for-Banded-Covariances/
10.1214/22-BA1333.full.

Lin:2006:LFB

[LLPR06] Rongheng Lin, Thomas A. Louis, Susan M. Paddock, and Greg Ridgeway. Loss function based ranking in two-stage, hierarchical models. Bayesian Analysis, 1(4):915–946, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/lin.pdf;   http://projecteuclid.org/euclid.ba/
1340370947.

Lavine:2021:AVS

[LLW21] Isaac Lavine, Michael Lindon, and Mike West. Adaptive variable selection for sequential prediction in multivariate dynamic models. Bayesian Analysis, 16(4):1059–1083, December 2021. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-4/Adaptive-Variable-Selection-for-Sequential-
Prediction-in-Multivariate-Dynamic-Models/10.1214/20-
BA1245.full.

Leorato:2016:SPD

[LM16] Samantha Leorato and Maura Mezzetti. Spatial panel data model with error dependence: A Bayesian separable covariance approach. Bayesian Analysis, 11(4):1035–1069, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1446124569.

Leorato:2021:BFM

[LM21] Samantha Leorato and Maura Mezzetti. A Bayesian factor model for spatial panel data with a separable covariance approach. Bayesian Analysis, 16(2):489–519, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/A-Bayesian-Factor-Model-for-Spatial-Panel-
Data-with-a/10.1214/20-BA1215.full.

Li:2020:UBL

[LMC20] Zehang Richard Li, Tyler H. McComick, and Samuel J. Clark. Using Bayesian latent Gaussian graphical models to infer symptom associations in verbal autopsies. Bayesian Analysis, 15(3):781–807, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Using-Bayesian-Latent-Gaussian-Graphical-
Models-to-Infer-Symptom-Associations/10.1214/19-BA1172.
full.

Lin:2019:EGP

[LMCD19] Lizhen Lin, Niu Mu, Pokman Cheung, and David Dunson. Extrinsic Gaussian processes for regression and classification on manifolds. Bayesian Analysis, 14(3):887–906, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1560240032.

Lewis:2021:BRL

[LML21] John R. Lewis, Steven N. MacEachern, and Yoonkyung Lee. Bayesian restricted likelihood methods: Conditioning on insufficient statistics in Bayesian regression (with discussion). Bayesian Analysis, 16(4):1393–1462, December 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-4/Bayesian-Restricted-Likelihood-Methods--Conditioning-
on-Insufficient-Statistics-in/10.1214/21-BA1257.full.

Lee:2014:LMP

[LMLM14] Juhee Lee, Steven N. MacEachern, Yiling Lu, and Gordon B. Mills. Local-mass preserving prior distributions for nonparametric Bayesian models. Bayesian Analysis, 9(2):307–330, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue02/juhee.pdf;   http://projecteuclid.org/euclid.
ba/1401148311.

Liquet:2017:BVS

[LMPS17] B. Liquet, K. Mengersen, A. N. Pettitt, and M. Sutton. Bayesian variable selection regression of multivariate responses for group data. Bayesian Analysis, 12(4):1039–1067, December 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1508983455.

Li:2008:CPB

[LN08] Longhai Li and Radford M. Neal. Compressing parameters in Bayesian high-order models with application to logistic sequence models. Bayesian Analysis, 3(4):793–821, December 2008. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://ba.stat.cmu.edu/journal/2008/vol03/issue04/li.
pdf;   http://projecteuclid.org/euclid.ba/1340370409.

Liu:2013:BDF

[LN13] Suyu Liu and Jing Ning. A Bayesian dose-finding design for drug combination trials with delayed toxicities. Bayesian Analysis, 8(3):703–722, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue03/liu.pdf;   http://
projecteuclid.org/euclid.ba/1378729925.

Lee:2019:CPA

[LNR19] Jeong Eun Lee, Geoff K. Nicholls, and Robin J. Ryder. Calibration procedures for approximate Bayesian credible sets. Bayesian Analysis, 14(4):1245–1269, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1570089912.

Lee:2016:ISS

[LR16] Jeong Eun Lee and Christian P. Robert. Importance sampling schemes for evidence approximation in mixture models. Bayesian Analysis, 11(2):573–597, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1440507475.

Lee:2006:IPD

[LSZH06] Herbert K. H. Lee, Bruno Sansó, Weining Zhou, and David M. Higdon. Inferring particle distribution in a proton accelerator experiment. Bayesian Analysis, 1(2):249–264, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2006/
vol01/issue02/lee249-264.pdf;   http://projecteuclid.
org/euclid.ba/1340371062.

Lin:2021:PIB

[LTY21] Ruitao Lin, Peter F. Thall, and Ying Yuan. A phase i-II basket trial design to optimize dose-schedule regimes based on delayed outcomes. Bayesian Analysis, 16(1):179–202, March 2021. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-1/A-Phase-III-Basket-Trial-Design-to-
Optimize-Dose-Schedule/10.1214/20-BA1205.full.

Lambert:2022:RRM

[LV22] Ben Lambert and Aki Vehtari. R^{∗}: a robust MCMC convergence diagnostic with uncertainty using decision tree classifiers. Bayesian Analysis, 17(2):353–379, June 2022. ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://doi.org/

10.1214/20-ba1252.

Leisen:2020:COP

[LVW20] Fabrizio Leisen, Cristiano Villa, and Stephen G. Walker. On a class of objective priors from scoring rules (with discussion). Bayesian Analysis, 15(4):1345–1423, December 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-15/issue-4/On-a-Class-of-Objective-Priors-from-
Scoring-Rules-with/10.1214/19-BA1187.full.

Liu:2009:DMS

[LW09] Fei Liu and Mike West. A dynamic modelling strategy for Bayesian computer model emulation. Bayesian Analysis, 4(2):393–411, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue02/liu.pdf;   http://
projecteuclid.org/euclid.ba/1340370283.

Li:2010:BRQ

[LXL10] Qing Li, Ruibin Xi, and Nan Lin. Bayesian regularized quantile regression. Bayesian Analysis, 5(3):533–556, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue03/lixin.pdf;   http://projecteuclid.org/euclid.
ba/1340380540.

Lee:2024:BOT

[LYL24] Kyoungjae Lee, Kisung You, and Lizhen Lin. Bayesian optimal two-sample tests for high-dimensional Gaussian populations. Bayesian Analysis, 19(3):869–893, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-3/Bayesian-Optimal-Two-Sample-Tests-
for-High-Dimensional-Gaussian-Populations/10.1214/23-
BA1373.full.

Lysy:2016:CAC

[Lys16] Martin Lysy. Comment on article by Chkrebtii, Campbell, Calderhead, and Girolami. Bayesian Analysis, 11(4):1269–1273, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.

ba/1480474948. See [CCCG16a, CCCG16b].

Li:2008:MAB

[LZN08] Longhai Li, Jianguo Zhang, and Radford M. Neal. A method for avoiding bias from feature selection with application to naive Bayes classification models. Bayesian Analysis, 3(1):171–196, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue01/li.pdf;   http://
projecteuclid.org/euclid.ba/1340370566.

Ma:2017:ASP

[Ma17] Li Ma. Adaptive shrinkage in Pólya tree type models. Bayesian Analysis, 12(3):779–805, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1473276260.

MacEachern:2007:CAJ

[Mac07] Steven N. MacEachern. Comment on article by Jain and Neal. Bayesian Analysis, 2(3):483–494, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2007/vol02/issue03/snm.pdf;
 http://projecteuclid.org/euclid.ba/1340370723.  See

[JN07b].

Madrigal:2007:CAD

[Mad07] Ana Maria Madrigal. Cluster allocation design networks. Bayesian Analysis, 2(3):557–589, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue03/madrigal.
pdf;   http://projecteuclid.org/euclid.ba/1340370727.

Monteiro:2011:PPM

[MAL11] João V. D. Monteiro, Renato M. Assunção, and Rosangela H. Loschi. Product partition models with correlated parameters. Bayesian Analysis, 6(4):691–726, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2011/vol06/issue04/monteiro.
pdf;   http://projecteuclid.org/euclid.ba/1339616541.

Martinez:2024:BET

[Mar24] Asael Fabian Martínez. Bayesian estimation of topological features of persistence diagrams. Bayesian Analysis, 19(1):1–20, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-1/Bayesian-Estimation-
of-Topological-Features-of-Persistence-Diagrams/10.1214/
22-BA1341.full.

Mukhopadhyay:2012:PSM

[MB12] Sabyasachi Mukhopadhyay and Sourabh Bhattacharya. Perfect simulation for mixtures with known and unknown number of components. Bayesian Analysis, 7(3):675–714, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue03/sabya.pdf;   http://projecteuclid.org/euclid.
ba/1346158780.

Monod:2023:RBS

[MBB^{+}23] Mélodie Monod, Alexandra Blenkinsop, Andrea Brizzi, Yu Chen, Carlos Cardoso Correia Perello, Vidoushee Jogarah, Yuanrong Wang, Seth Flaxman, Samir Bhatt, and Oliver Ratmann. Regularised B-splines projected Gaussian process priors to estimate time-trends in age-specific COVID-19 deaths. Bayesian Analysis, 18(3):957–987, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Regularised-B-splines-Projected-Gaussian-
Process-Priors-to-Estimate-Time/10.1214/22-BA1334.full.

Martinez-Beneito:2017:TMA

[MBBRB17] Miguel A. Martinez-Beneito, Paloma Botella-Rocamora, and Sudipto Banerjee. Towards a multidimensional approach to Bayesian disease mapping. Bayesian Analysis, 12(1):239–259, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1458324098.

Modi:2024:DRH

[MBC24] Chirag Modi, Alex Barnett, and Bob Carpenter. Delayed rejection Hamiltonian Monte Carlo for sampling multiscale distributions. Bayesian Analysis, 19(3):815–842, September 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-3/Delayed-rejection-Hamiltonian-Monte-
Carlo-for-sampling-multiscale-distributions/10.1214/
23-BA1360.full.

Ma:2007:BMA

[MC07] Haijun Ma and Bradley P. Carlin. Bayesian multivariate areal wombling for multiple disease boundary analysis. Bayesian Analysis, 2(2):281–302, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue02/ma.pdf;   http://
projecteuclid.org/euclid.ba/1340370801.

Mlan:2015:OBI

[MC15] Cyr Emile M’lan and Ming-Hui Chen. Objective Bayesian inference for bilateral data. Bayesian Analysis, 10(1):139–170, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1422468426.

Moran:2023:PPN

[MCB23] Gemma E. Moran, John P. Cunningham, and David M. Blei. The posterior predictive null. Bayesian Analysis, 18(4):1071– 1097, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-18/issue-4/The-
Posterior-Predictive-Null/10.1214/22-BA1313.full.

Mallick:2011:CAP

[MCG11] Bani K. Mallick, Sounak Chakraborty, and Malay Ghosh. Comment on article by Polson and Scott. Bayesian Analysis, 6 (1):25–29, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue01/mallick.pdf;   http://
projecteuclid.org/euclid.ba/1339611937. See [PS11a].

Moore:2020:DPM

[MCMK20] Camille M. Moore, Nichole E. Carlson, Samantha MaWhinney, and Sarah Kreidler. A Dirichlet process mixture model for non-ignorable dropout. Bayesian Analysis, 15 (4):1139–1167, December 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-4/A-Dirichlet-Process-Mixture-Model-for-Non-
Ignorable-Dropout/10.1214/19-BA1181.full.

Manolopoulou:2010:R

[MCW10a] Ioanna Manolopoulou, Cliburn Chan, and Mike West. Rejoinder. Bayesian Analysis, 5(3):461–463, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue03/manolopoulou_rejoinder.pdf; http://projecteuclid.
org/euclid.ba/1340380536. See [MCW10b].

Manolopoulou:2010:SSL

[MCW10b] Ioanna Manolopoulou, Cliburn Chan, and Mike West. Selection sampling from large data sets for targeted inference in mixture modeling. Bayesian Analysis, 5(3):429–449, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue03/manolopoulou.pdf;   http://projecteuclid.org/

euclid.ba/1340369756. See comments [Rig10, Whi10] and rejoinder [MCW10a].

Migliorati:2018:NRM

[MDO18] Sonia Migliorati, Agnese Maria Di Brisco, and Andrea Ongaro. A new regression model for bounded responses. Bayesian Analysis, 13(3):845–872, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1508897093.

Mulder:2019:BFT

[MF19] Joris Mulder and Jean-Paul Fox. Bayes factor testing of multiple intraclass correlations. Bayesian Analysis, 14(2):521–552, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1533866668.

Marmin:2022:DGP

[MF22] Sébastien Marmin and Maurizio Filippone. Deep Gaussian processes for calibration of computer models (with discussion). Bayesian Analysis, 17(4):1301–1350, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Deep-Gaussian-Processes-for-Calibration-
of-Computer-Models-with-Discussion/10.1214/21-BA1293.
full.

Mulgrave:2020:BIN

[MG20] Jami J. Mulgrave and Subhashis Ghosal. Bayesian inference in nonparanormal graphical models. Bayesian Analysis, 15(2):449– 475, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1559721629.

Manderson:2023:CCB

[MG23] Andrew A. Manderson and Robert J. B. Goudie. Combining chains of Bayesian models with Markov melding. Bayesian Analysis, 18(3):807–840, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Combining-Chains-of-Bayesian-Models-with-
Markov-Melding/10.1214/22-BA1327.full.

Mendoza:2015:CAB

[MGP15] Manuel Mendoza and Eduardo Gutiérrez-Peña. Comment on article by Berger, Bernardo, and Sun. Bayesian Analysis, 10(1): 227–231, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). See [BBS15a].

Murray:2016:FBS

[MHSC16] Thomas A. Murray, Brian P. Hobbs, Daniel J. Sargent, and Bradley P. Carlin. Flexible Bayesian survival modeling with semiparametric time-dependent and shape-restricted covariate effects. Bayesian Analysis, 11(2):381–402, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1431607819.

Millard:2008:CAB

[Mil08] Andrew R. Millard. Comment on article by Blackwell and Buck. Bayesian Analysis, 3(2):255–261, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2008/vol03/issue02/millard.
pdf;   http://projecteuclid.org/euclid.ba/1340370546.

See [BB08a].

MLan:2008:BSS

[MJW08] Cyr E. M’Lan, Lawrence Joseph, and David B. Wolfson. Bayesian sample size determination for binomial proportions. Bayesian Analysis, 3(2):269–296, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue02/mlan.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370548.

Magee:2024:HTY

[MKIM24] Andrew Magee, Michael Karcher, Frederick A. Matsen IV, and Volodymyr M. Minin. How trustworthy is your tree?

Bayesian phylogenetic effective sample size through the lens of Monte Carlo error. Bayesian Analysis, 19(2):565–593, June 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-2/How-Trustworthy-
Is-Your-Tree-Bayesian-Phylogenetic-Effective-Sample-
Size/10.1214/22-BA1339.full.

Ma:2022:PCB

[ML22] Yucong Ma and Jun S. Liu. On posterior consistency of Bayesian factor models in high dimensions. Bayesian Analysis, 17(3): 901–929, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-17/issue-3/On-
Posterior-Consistency-of-Bayesian-Factor-Models-in-High-
Dimensions/10.1214/21-BA1281.full.

Moller:2007:EAM

[MM07] Jesper Møller and Kerrie Mengersen. Ergodic averages for monotone functions using upper and lower dominating processes. Bayesian Analysis, 2(4):761–781, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2007/vol02/issue04/moller.
pdf;   http://projecteuclid.org/euclid.ba/1340370714.

Muller:2013:BNI

[MM13a] Peter Müller and Riten Mitra. Bayesian nonparametric inference – why and how. Bayesian Analysis, 8(2):269–302, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue02/mueller.pdf;   http://projecteuclid.org/euclid.

ba/1369407550. See comments [CM13, Hof13, O’H13, O’H13, APA^{+}13] and rejoinder [MM13b].

Muller:2013:R

[MM13b] Peter Müller and Riten Mitra. Rejoinder. Bayesian Analysis, 8(2):357–360, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue02/mueller_rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1369407555. See [MM13a].

Martinez:2014:NCP

[MM14] Asael Fabian Martínez and Ramsés H. Mena. On a nonparametric change point detection model in Markovian regimes. Bayesian Analysis, 9(4):823–858, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1416579181.

Michalak:2016:PPH

[MM16] Sarah E. Michalak and Carl N. Morris. Posterior propriety for hierarchical models with log-likelihoods that have norm bounds. Bayesian Analysis, 11(2):545–571, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1437137635.

Mitra:2016:BGM

[MMJ16] Riten Mitra, Peter Müller, and Yuan Ji. Bayesian graphical models for differential pathways. Bayesian Analysis, 11(1):99–124, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1423839169.

Maroulas:2022:BTL

[MMN22] Vasileios Maroulas, Cassie Putman Micucci, and Farzana Nasrin. Bayesian topological learning for classifying the structure of biological networks. Bayesian Analysis, 17(3):711–736, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-3/Bayesian-Topological-
Learning-for-Classifying-the-Structure-of-Biological-
Networks/10.1214/21-BA1270.full.

McKinley:2015:BMC

[MMW15] Trevelyan J. McKinley, Michelle Morters, and James L. N. Wood. Bayesian model choice in cumulative link ordinal regression models. Bayesian Analysis, 10(1):1–30, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422468421.

Moores:2020:SBI

[MNPM20] Matthew Moores, Geoff Nicholls, Anthony Pettitt, and Kerrie Mengersen. Scalable Bayesian inference for the inverse temperature of a hidden Potts model. Bayesian Analysis, 15(1):

1–27, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1544583756.

McKinley:2020:EBM

[MNS^{+}20] Trevelyan J. McKinley, Peter Neal, Simon E. F. Spencer, Andrew J. K. Conlan, and Laurence Tiley. Efficient Bayesian model choice for partially observed processes: With application to an experimental transmission study of an infectious disease. Bayesian Analysis, 15(3):839–870, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-15/issue-3/Efficient-Bayesian-Model-Choice-for-
Partially-Observed-Processes--With/10.1214/19-BA1174.
full.

Mulder:2018:MPE

[MP18] Joris Mulder and Luis Raúl Pericchi. The matrix-F prior for estimating and testing covariance matrices. Bayesian Analysis, 13(4):1193–1214, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1515747744.

Majumdar:2010:SAM

[MPK10] Anandamayee Majumdar, Debashis Paul, and Jason Kaye. Sensitivity analysis and model selection for a generalized convolution model for spatial processes. Bayesian Analysis, 5 (3):493–518, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue03/paul.pdf;   http:/
/projecteuclid.org/euclid.ba/1340380538.

Murua:2022:BSB

[MQ22] Alejandro Murua and Fernando Andrés Quintana. Biclustering via semiparametric Bayesian inference. Bayesian Analysis, 17(3):969–995, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-3/Biclustering-via-Semiparametric-Bayesian-
Inference/10.1214/21-BA1284.full.

Mason:2012:TPS

[MRB12] Alexina Mason, Sylvia Richardson, and Nicky Best. Twopronged strategy for using DIC to compare selection models with non-ignorable missing responses. Bayesian Analysis, 7(1):109– 146, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2012/vol07/issue01/mason.pdf;   http://projecteuclid.
org/euclid.ba/1339616727.

Moran:2019:VPF

[MRG19] Gemma E. Moran, Veronika Rocková, and Edward I. George. Variance prior forms for high-dimensional Bayesian variable selection. Bayesian Analysis, 14(4):1091–1119, December 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1551927862.

Marshall:2007:IOB

[MS07a] E. C. Marshall and D. J. Spiegelhalter. Identifying outliers in Bayesian hierarchical models: a simulation-based approach. Bayesian Analysis, 2(2):409–444, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2007/vol02/issue02/marshall.pdf;
 http://projecteuclid.org/euclid.ba/1340393242.

Millar:2007:ALI

[MS07b] Russell B. Millar and Wayne S. Stewart. Assessment of locally influential observations in Bayesian models. Bayesian Analysis, 2(2):365–383, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2007/vol02/issue02/millar.pdf;
http://projecteuclid.org/euclid.ba/1340393240.

Monni:2009:R

[MT09a] Stefano Monni and Mahlet G. Tadesse. Rejoinder. Bayesian Analysis, 4(3):457–464, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2009/vol04/issue03/monni_rejoinder.
pdf;   http://projecteuclid.org/euclid.ba/1340369850.

See [MT09b].

Monni:2009:SPM

[MT09b] Stefano Monni and Mahlet G. Tadesse. A stochastic partitioning method to associate high-dimensional responses and covariates. Bayesian Analysis, 4(3):413–436, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/

issue03/monni.pdf;   http://projecteuclid.org/euclid.

ba/1340369845. See comments [CGM09, Fra09, Li09, Ste09] and rejoinder [MT09a].

Morita:2012:PES

[MTM12] Satoshi Morita, Peter F. Thall, and Peter Müller. Prior effective sample size in conditionally independent hierarchical models. Bayesian Analysis, 7(3):591–614, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue03/morita.
pdf;   http://projecteuclid.org/euclid.ba/1346158777.

Murray:2021:RAI

[MTS^{+}21] Thomas A. Murray, Peter F. Thall, Frederique Schortgen, Pierre Asfar, Sarah Zohar, and Sandrine Katsahian. Robust adaptive incorporation of historical control data in a randomized trial of external cooling to treat septic shock. Bayesian Analysis, 16(3): 825–844, September 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-3/Robust-
Adaptive-Incorporation-of-Historical-Control-Data-in-
a-Randomized/10.1214/20-BA1229.full.

Meng:2006:CAC

[MV06] Xiao-Li Meng and Florin Vaida. Comment on article by Celeux et al. Bayesian Analysis, 1(4):687–698, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/meng.pdf;   http://projecteuclid.org/euclid.

ba/1340370937. See [CFRT06a].

Murphy:2020:IMI

[MVG20] Keefe Murphy, Cinzia Viroli, and Isobel Claire Gormley. Infinite mixtures of infinite factor analysers. Bayesian Analysis, 15(3):937–963, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Infinite-Mixtures-of-Infinite-Factor-Analysers/
10.1214/19-BA1179.full.

Mohammadi:2015:BSL

[MW15] A. Mohammadi and E. C. Wit. Bayesian structure learning in sparse Gaussian graphical models. Bayesian Analysis, 10(1):109– 138, March 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1422468425.

McLean:2019:VMP

[MW19] M. W. McLean and M. P. Wand. Variational message passing for elaborate response regression models. Bayesian Analysis, 14(2): 371–398, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1527213628.

McCullagh:2008:HMC

[MY08] Peter McCullagh and Jie Yang. How many clusters? Bayesian Analysis, 3(1):101–120, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue01/yang.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370563.

Mallick:2016:CAC

[MYGE16] Bani K. Mallick, Keren Yang, Nilabja Guha, and Yalchin Efendiev. Comment on article by Chkrebtii, Campbell, Calderhead, and Girolami. Bayesian Analysis, 11(4):1279–1284, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.

ba/1479805385. See [CCCG16a, CCCG16b].

Masotti:2024:GBF

[MZMK24] Maria Masotti, Lin Zhang, Gregory J. Metzger, and Joseph S. Koopmeiners. A general Bayesian functional spatial partitioning method for multiple region discovery applied to prostate cancer MRI. Bayesian Analysis, 19(2):623–647, June 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/A-General-Bayesian-Functional-Spatial-
Partitioning-Method-for-Multiple-Region/10.1214/23-BA1366.
full.

Naulet:2018:SAS

[NB18] Zacharie Naulet and Éric Barat. Some aspects of symmetric gamma process mixtures. Bayesian Analysis, 13(3):703–720, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1507341640.

Nieto-Barajas:2014:BNA

[NBCC14] Luis E. Nieto-Barajas and Alberto Contreras-Cristán. A Bayesian nonparametric approach for time series clustering. Bayesian Analysis, 9(1):147–170, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2014/vol09/issue01/barajas.
pdf;   http://projecteuclid.org/euclid.ba/1393251774.

Nishimura:2020:RIS

[ND20] Akihiko Nishimura and David Dunson. Recycling intermediate steps to improve Hamiltonian Monte Carlo. Bayesian Analysis, 15(4):1087–1108, December 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-4/Recycling-Intermediate-Steps-to-Improve-Hamiltonian-
Monte-Carlo/10.1214/19-BA1171.full.

Nott:2018:ABP

[NDME18] David J. Nott, Christopher C. Drovandi, Kerrie Mengersen, and Michael Evans. Approximation of Bayesian predictive. Bayesian Analysis, 13(1):59–83, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
13/issue-1/Approximation-of-Bayesian-Predictive-p-Values-
with-Regression-ABC/10.1214/16-BA1033.full.

Nguyen:2020:NAM

[NdVA^{+}20] Dao Nguyen, Perry de Valpine, Yves Atchade, Daniel Turek, Nicholas Michaud, and Christopher Paciorek. Nested adaptation of MCMC algorithms. Bayesian Analysis, 15(4):1323– 1343, December 2020. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-15/issue-4/Nested-
Adaptation-of-MCMC-Algorithms/10.1214/19-BA1190.full.

Neelon:2019:BZI

[Nee19] Brian Neelon. Bayesian zero-inflated negative binomial regression based on Pólya–Gamma mixtures. Bayesian Analysis, 14 (3):829–855, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1560240030.

Ning:2019:BMC

[NGT19] Bo Ning, Subhashis Ghosal, and Jewell Thomas. Bayesian method for causal inference in spatially-correlated multivariate time series. Bayesian Analysis, 14(1):1–28, March 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1522202634.

Nguyen:2010:IGC

[Ngu10] XuanLong Nguyen. Inference of global clusters from locally distributed data. Bayesian Analysis, 5(4):817–845, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2010/
vol05/issue04/nguyen.pdf;   http://projecteuclid.org/
euclid.ba/1340110856.

Nikooienejad:2021:EUM

[NJ21] Amir Nikooienejad and Valen E. Johnson. On the existence of uniformly most powerful Bayesian tests with application to non-central chi-squared tests. Bayesian Analysis, 16(1):93– 109, March 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-1/On-the-
Existence-of-Uniformly-Most-Powerful-Bayesian-Tests-
With/10.1214/19-BA1194.full.

Ni:2018:RGM

[NJM18] Yang Ni, Yuan Ji, and Peter Müller. Reciprocal graphical models for integrative gene regulatory network analysis. Bayesian Analysis, 13(4):1095–1110, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1512097377.

Nyman:2014:SGM

[NPKC14] Henrik Nyman, Johan Pensar, Timo Koski, and Jukka Corander. Stratified graphical models — context-specific independence in graphical models. Bayesian Analysis, 9(4):883–908, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1416579183.

Nemeth:2018:MMS

[NS18] Christopher Nemeth and Chris Sherlock. Merging MCMC subposteriors through Gaussian-process approximations. Bayesian Analysis, 13(2):507–530, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1502265628.

Nishimura:2023:SSS

[NS23] Akihiko Nishimura and Marc A. Suchard. Shrinkage with shrunken shoulders: Gibbs sampling shrinkage model posteriors with guaranteed convergence rates. Bayesian Analysis, 18(2):367–390, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Shrinkage-with-Shrunken-Shoulders--Gibbs-
Sampling-Shrinkage-Model-Posteriors/10.1214/22-BA1308.
full.

Nott:2021:UPE

[NSAL^{+}21] David J. Nott, Max Seah, Luai Al-Labadi, Michael Evans, Hui Khoon Ng, and Berthold-Georg Englert. Using prior expansions for prior-data conflict checking. Bayesian Analysis, 16(1):203–231, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/Using-Prior-Expansions-for-Prior-Data-Conflict-
Checking/10.1214/20-BA1204.full.

Ntzoufras:2019:PBI

[NTL19] Ioannis Ntzoufras, Claudia Tarantola, and Monia Lupparelli. Probability based independence sampler for Bayesian quantitative learning in graphical log–linear marginal models. Bayesian Analysis, 14(3):777–803, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1560240028.

Ogle:2013:FMB

[OBS13] Kiona Ogle, Jarrett Barber, and Karla Sartor. Feedback and modularization in a Bayesian meta–analysis of tree traits affecting forest dynamics. Bayesian Analysis, 8(1):133–168, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/ogle.pdf;   http://projecteuclid.org/euclid.
ba/1362406655.

Osthus:2019:DBI

[OGPD19] Dave Osthus, James Gattiker, Reid Priedhorsky, and Sara Y. Del Valle. Dynamic Bayesian influenza forecasting in the United States with hierarchical discrepancy (with discussion). Bayesian Analysis, 14(1):261–312, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1533866670.

OHagan:2006:SSS

[O’H06] Anthony O’Hagan. Science, subjectivity and software (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1 (3):445–450, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue03/ohagan.pdf;   http://
projecteuclid.org/euclid.ba/1340371043. See [Ber06a,

Gol06a].

OHagan:2013:CAM

[O’H13] Anthony O’Hagan. Comment on article by Müller and Mitra. Bayesian Analysis, 8(2):319–322, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2013/vol08/issue02/ohagan.
pdf;   http://projecteuclid.org/euclid.ba/1369407553.

See [MM13a].

Obiang:2023:BAP

[OJP23] Eunice Okome Obiang, Pascal Jézéquel, and Frédéric Proïa. A Bayesian approach for partial Gaussian graphical models with sparsity. Bayesian Analysis, 18(2):465–490, June 2023. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-2/A-Bayesian-Approach-for-Partial-Gaussian-
Graphical-Models-With-Sparsity/10.1214/22-BA1315.full.

Ohn:2022:PCF

[OK22] Ilsang Ohn and Yongdai Kim. Posterior consistency of factor dimensionality in high-dimensional sparse factor models. Bayesian Analysis, 17(2):491–514, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/Posterior-Consistency-of-Factor-Dimensionality-
in-High-Dimensional-Sparse-Factor/10.1214/21-BA1261.
full.

Overstall:2020:BDE

[OM20] Antony Overstall and James McGree. Bayesian design of experiments for intractable likelihood models using coupled auxiliary models and multivariate emulation. Bayesian Analysis, 15(1):103–131, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1550199643.

Overstall:2022:BDT

[OM22] Antony Overstall and James McGree. Bayesian decisiontheoretic design of experiments under an alternative model. Bayesian Analysis, 17(4):1021–1041, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Bayesian-Decision-Theoretic-Design-
of-Experiments-Under-an-Alternative-Model/10.1214/21-
BA1286.full.

OliveiraPrates:2019:ASC

[OMC19] Marcos Oliveira Prates, Renato Martins Assunção, and Erica Castilho Rodrigues. Alleviating spatial confounding for areal data problems by displacing the geographical centroids. Bayesian Analysis, 14(2):623–647, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1537258137.

OHara:2009:RBV

[OS09] R. B. O’Hara and M. J. Sillanpää. A review of Bayesian variable selection methods: what, how and which. Bayesian Analysis, 4(1):85–117, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue01/ohara.pdf;   http:
//projecteuclid.org/euclid.ba/1340370391.

Ohnishi:2024:BAT

[OS24] Yuki Ohnishi and Arman Sabbaghi. A Bayesian analysis of two-stage randomized experiments in the presence of interference, treatment nonadherence, and missing outcomes. Bayesian Analysis, 19(1):205–234, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/A-Bayesian-Analysis-of-Two-Stage-Randomized-
Experiments-in-the/10.1214/22-BA1347.full.

Paciorek:2006:MCP

[Pac06] Christopher J. Paciorek. Misinformation in the conjugate prior for the linear model with implications for free-knot spline modelling. Bayesian Analysis, 1(2):375–383, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue02/paciorek375-383.pdf;   http://projecteuclid.
org/euclid.ba/1340371068.

Pajor:2017:EML

[Paj17] Anna Pajor. Estimating the marginal likelihood using the arithmetic mean identity. Bayesian Analysis, 12(1):261–287, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-12/issue-1/Estimating-the-Marginal-
Likelihood-Using-the-Arithmetic-Mean-Identity/10.1214/
16-BA1001.full.

Passeggeri:2023:QID

[Pas23] Riccardo Passeggeri. On quasi-infinitely divisible random measures. Bayesian Analysis, 18(1):253–286, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-1/On-Quasi-Infinitely-Divisible-Random-
Measures/10.1214/21-BA1289.full.

Pena:2020:RTI

[PB20] Víctor Peña and James O. Berger. Restricted type II maximum likelihood priors on regression coefficients. Bayesian Analysis, 15(4):1281–1297, December 2020. ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://doi.org/10.1214/19-ba1188.

Porcu:2021:NBM

[PBT^{+}21] Emilio Porcu, Pier Giovanni Bissiri, Felipe Tagle, Rubén Soza, and Fernando A. Quintana. Nonparametric Bayesian modeling and estimation of spatial correlation functions for global data. Bayesian Analysis, 16(3):845–873, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-3/Nonparametric-Bayesian-Modeling-and-
Estimation-of-Spatial-Correlation-Functions-for/10.1214/
20-BA1228.full.

Peluso:2019:SMM

[PCM19] Stefano Peluso, Siddhartha Chib, and Antonietta Mira. Semiparametric multivariate and multiple change-point modeling. Bayesian Analysis, 14(3):727–751, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1560240025.

pour:2020:TOB

[pD20] Ali Foroughi pour and Lori A. Dalton. Theory of optimal Bayesian feature filtering. Bayesian Analysis, 15(4):1169–1197, December 2020. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-15/issue-4/Theory-
of-Optimal-Bayesian-Feature-Filtering/10.1214/19-BA1182.
full.

Peruggia:2007:BMD

[Per07] Mario Peruggia. Bayesian model diagnostics based on artificial autoregressive errors. Bayesian Analysis, 2(4):817– 841, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2007/vol02/issue04/peruggia.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370716.

Pamminger:2010:MBC

[PFS10] Christoph Pamminger and Sylvia Frühwirth-Schnatter. Modelbased clustering of categorical time series. Bayesian Analysis, 5(2):345–368, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2010/vol05/issue02/pamminger.pdf;   http:
//projecteuclid.org/euclid.ba/1340218342.

Puelz:2017:VSS

[PHC17] David Puelz, P. Richard Hahn, and Carlos M. Carvalho. Variable selection in seemingly unrelated regressions with random predictors. Bayesian Analysis, 12(4):969–989, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1488855633.

Prangle:2023:BED

[PHG23] Dennis Prangle, Sophie Harbisher, and Colin S. Gillespie. Bayesian experimental design without posterior calculations: an adversarial approach. Bayesian Analysis, 18(1):133–163, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-1/Bayesian-Experimental-
Design-Without-Posterior-Calculations-An-Adversarial-
Approach/10.1214/22-BA1306.full.

Paganin:2021:CPP

[PHOD21] Sally Paganin, Amy H. Herring, Andrew F. Olshan, and David B. Dunson. Centered partition processes: Informative priors for clustering (with discussion). Bayesian Analysis, 16(1):301–370, March 2021. ISSN 1931-6690 (print), 1931-6690 (electronic).

Pourzanjani:2021:BIS

[PJM^{+}21] Arya A. Pourzanjani, Richard M. Jiang, Brian Mitchell, Paul J. Atzberger, and Linda R. Petzold. Bayesian inference over the Stiefel manifold via the Givens representation. Bayesian Analysis, 16(2):639–666, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/Bayesian-Inference-over-the-Stiefel-Manifold-
via-the-Givens-Representation/10.1214/20-BA1202.full.

Papageorgiou:2024:PRB

[PK24] Ioannis Papageorgiou and Ioannis Kontoyiannis. Posterior representations for Bayesian context trees: Sampling, estimation and convergence. Bayesian Analysis, 19(2):501–529, June 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/Posterior-Representations-for-Bayesian-
Context-Trees--Sampling-Estimation-and/10.1214/23-BA1362.
full.

Peters:2011:BCV

[PKL^{+}11] Gareth W. Peters, Balakrishnan Kannan, Ben Lasscock, Chris Mellen, and Simon Godsill. Bayesian cointegrated vector autoregression models incorporating alpha-stable noise for interday price movements via approximate Bayesian computation. Bayesian Analysis, 6(4):755–792, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue04/peters.
pdf;   http://projecteuclid.org/euclid.ba/1339616543.

Peters:2010:MSA

[PKLM10] Gareth W. Peters, Balakrishnan Kannan, Ben Lasscock, and Chris Mellen. Model selection and adaptive Markov chain Monte Carlo for Bayesian cointegrated VAR models. Bayesian Analysis, 5(3):465–491, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue03/peters.pdf;
http://projecteuclid.org/euclid.ba/1340380537.

Pourmohamad:2016:MSP

[PL16] Tony Pourmohamad and Herbert K. H. Lee. Multivariate stochastic process models for correlated responses of mixed type. Bayesian Analysis, 11(3):797–820, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1444308210.

Plummer:2006:CAC

[Plu06] Martyn Plummer. Comment on article by Celeux et al. Bayesian Analysis, 1(4):681–686, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue04/plummer.
pdf;   http://projecteuclid.org/euclid.ba/1340370936.

See [CFRT06a].

Papamarkou:2014:ZVD

[PMG14] Theodore Papamarkou, Antonietta Mira, and Mark Girolami. Zero variance differential geometric Markov chain Monte Carlo algorithms. Bayesian Analysis, 9(1):97–128, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2014/
vol09/issue01/papamarkou.pdf;   http://projecteuclid.
org/euclid.ba/1393251772.

Paulon:2024:BNB

[PMS24] Giorgio Paulon, Peter Müller, and Victor G. Sal y Rosas. Bayesian nonparametric bivariate survival regression for current status data. Bayesian Analysis, 19(1):49–75, March 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-1/Bayesian-Nonparametric-Bivariate-Survival-
Regression-for-Current-Status-Data/10.1214/22-BA1346.
full.

Pensar:2017:MPL

[PNNC17] Johan Pensar, Henrik Nyman, Juha Niiranen, and Jukka Corander. Marginal pseudo-likelihood learning of discrete Markov network structures. Bayesian Analysis, 12(4):1195–1215, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1477918728.

Poirier:2006:GBM

[Poi06] Dale J. Poirier. The growth of Bayesian methods in statistics and economics since 1970. Bayesian Analysis, 1(4): 969–979, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue04/poirier.pdf;   http://
projecteuclid.org/euclid.ba/1340370949.

Polettini:2017:GSB

[Pol17] Silvia Polettini. A generalised semiparametric Bayesian Fay– Herriot model for small area estimation shrinking both means and variances. Bayesian Analysis, 12(3):729–752, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1473276257.

Poole:2010:CAV

[Poo10] David Poole. Comment on article by Vernon et al. Bayesian Analysis, 5(4):671–675, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2010/vol05/issue04/poole.pdf;
 http://projecteuclid.org/euclid.ba/1340110847.  See

[VGB10a].

Pajor:2024:BEP

[POWK24] Anna Pajor, Jacek Osiewalski, Justyna Wróblewska, and Lukasz Kwiatkowski. Bayesian ex post evaluation of recursive multi-step-ahead density prediction. Bayesian Analysis, 19(3):751–783, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-3/Bayesian-ex-Post-Evaluation-of-Recursive-
Multi-Step-Ahead-Density/10.1214/23-BA1367.full.

Popova:2008:BFP

[PPG08] Ivilina Popova, Elmira Popova, and Edward I. George. Bayesian forecasting of prepayment rates for individual pools of mortgages. Bayesian Analysis, 3(2):393–426, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue02/popova.pdf;
http://projecteuclid.org/euclid.ba/1340370553.

Perez:2017:SBD

[PPR17] María-Eglée Pérez, Luis Raúl Pericchi, and Isabel Cristina Ramírez. The scaled beta2 distribution as a robust prior for scales. Bayesian Analysis, 12(3):615–637, September 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1469553353.

Page:2015:PBC

[PQ15] Garritt L. Page and Fernando A. Quintana. Predictions based on the clustering of heterogeneous functions via shape and subjectspecific covariates. Bayesian Analysis, 10(2):379–410, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1422884979.

Page:2016:R

[PQ16a] Garrit L. Page and Fernando A. Quintana. Rejoinder. Bayesian Analysis, 11(1):315–323, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-11/issue-1/Rejoinder/10.1214/16-BA971REJ.full.

See [PQ16b, GPP16].

Page:2016:SPP

[PQ16b] Garritt L. Page and Fernando A. Quintana. Spatial product partition models. Bayesian Analysis, 11(1):265–298, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-11/issue-1/Spatial-Product-
Partition-Models/10.1214/15-BA971.full. See comments

[GL16, GPP16, RF16] and rejoinder [PQ16a].

Pratola:2016:EMH

[Pra16a] Matthew T. Pratola. Efficient Metropolis–Hastings proposal mechanisms for Bayesian regression tree models. Bayesian Analysis, 11(3):885–911, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1457383101. See comments

[Gra16, Han16, CLH^{+}16] and rejoinder [Pra16b].

Pratola:2016:R

[Pra16b] Matthew T. Pratola. Rejoinder. Bayesian Analysis, 11(3): 945–955, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1472829063. See [Pra16a].

Prangle:2017:AAD

[Pra17]  Dennis Prangle.  Adapting the ABC distance function.

Bayesian Analysis, 12(1):289–309, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-12/issue-1/Adapting-the-ABC-Distance-Function/
10.1214/16-BA1002.full.

Polson:2011:DAS

[PS11a] Nicholas G. Polson and Steven L. Scott. Data augmentation for support vector machines. Bayesian Analysis, 6(1):1–23, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue01/polson.pdf;   http://projecteuclid.org/

euclid.ba/1339611936. See comments [MCG11, SYvD11, Han11] and rejoinder [PS11b].

Polson:2011:RDA

[PS11b] Nicholas G. Polson and Steven L. Scott. Rejoinder: “Data augmentation for support vector machines”. Bayesian Analysis, 6(1):43–47, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue01/polson_rejoinder.
pdf;   http://projecteuclid.org/euclid.ba/1339611940.

See [PS11a].

Polson:2012:HCP

[PS12] Nicholas G. Polson and James G. Scott. On the half-Cauchy prior for a global scale parameter. Bayesian Analysis, 7(4):887–902, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue04/polson.pdf;
http://projecteuclid.org/euclid.ba/1354024466.

Peterson:2013:CAS

[PS13] Christine B. Peterson and Francesco C. Stingo. Comment on article by Scutari. Bayesian Analysis, 8(3):539– 542, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/peterson.pdf;   http:/
/projecteuclid.org/euclid.ba/1378729916. See [RS14a].

Panov:2015:FSB

[PS15] Maxim Panov and Vladimir Spokoiny. Finite sample Bernstein– von Mises theorem for semiparametric problems. Bayesian Analysis, 10(3):665–710, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884986.

Polson:2017:DLB

[PS17] Nicholas G. Polson and Vadim Sokolov. Deep learning: A Bayesian perspective. Bayesian Analysis, 12(4):1275–1304, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1510801992.

Park:2020:DSC

[PS20] Jong Hee Park and Yunkyu Sohn. Detecting structural changes in longitudinal network data. Bayesian Analysis, 15(1):133–157, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1550826221.

Picchini:2023:SGM

[PSC23] Umberto Picchini, Umberto Simola, and Jukka Corander. Sequentially guided MCMC proposals for synthetic likelihoods and correlated synthetic likelihoods. Bayesian Analysis, 18(4):1099–1129, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Sequentially-Guided-MCMC-Proposals-for-Synthetic-
Likelihoods-and-Correlated-Synthetic/10.1214/22-BA1305.
full.

Pal:2020:CPP

[PSMB20] Subhadip Pal, Subhajit Sengupta, Riten Mitra, and Arunava Banerjee. Conjugate priors and posterior inference for the matrix Langevin distribution on the Stiefel manifold. Bayesian Analysis, 15(3):871–908, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Conjugate-Priors-and-Posterior-Inference-
for-the-Matrix-Langevin-Distribution/10.1214/19-BA1176.
full.

Perala:2020:CEA

[PVC20] Tommi Perälä, Jarno Vanhatalo, and Anna Chrysafi. Calibrating expert assessments using hierarchical Gaussian process models. Bayesian Analysis, 15(4):1251–1280, December 2020. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-15/issue-4/Calibrating-Expert-Assessments-Using-
Hierarchical-Gaussian-Process-Models/10.1214/19-BA1180.
full.

Purutcuoglu:2008:BIM

[PW08] Vilda Purutçuoğlu and Ernst Wit. Bayesian inference for the MAPK/ERK pathway by considering the dependency of the kinetic parameters. Bayesian Analysis, 3(4):851–886, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue04/purutcuoglu.pdf;   http://projecteuclid.org/
euclid.ba/1340370411.

Pauger:2019:BEF

[PW19] Daniela Pauger and Helga Wagner. Bayesian effect fusion for categorical predictors. Bayesian Analysis, 14(2):341–369, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1527213627.

Paisley:2012:DIL

[PWB12] John Paisley, Chong Wang, and David M. Blei. The discrete infinite logistic normal distribution. Bayesian Analysis, 7(4): 997–1034, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue04/paisley.pdf;   http://
projecteuclid.org/euclid.ba/1354024470.

Qian:2018:BDB

[Qia18] Hang Qian. Big data Bayesian linear regression and variable selection by normal-inverse-gamma summation. Bayesian Analysis, 13(4):1011–1035, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1510110046.

Quintana:2009:CAJ

[QM09] Fernando A. Quintana and Peter Müller. Comment on article by Jensen et al. Bayesian Analysis, 4(4):665–668, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue04/quintana.pdf;   http://projecteuclid.org/euclid.
ba/1340369818. See [JMW09a].

Quintana:2008:SPB

[QMRM08] Fernando A. Quintana, Peter Müller, Gary L. Rosner, and Mark Munsell. Semi-parametric Bayesian inference for multiseason baseball data. Bayesian Analysis, 3(2):317–338, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue02/quintana.pdf;   http://projecteuclid.org/euclid.
ba/1340370550.

Quiroz:2023:GVA

[QNK23] Matias Quiroz, David J. Nott, and Robert Kohn. Gaussian variational approximations for high-dimensional state space models. Bayesian Analysis, 18(3):989–1016, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-3/Gaussian-Variational-Approximations-
for-High-dimensional-State-Space-Models/10.1214/22-BA1332.
full.

Quinlan:2024:JRP

[QPC24] José J. Quinlan, Garritt L. Page, and Luis M. Castro. Joint random partition models for multivariate change point analysis. Bayesian Analysis, 19(1):21–48, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-1/Joint-Random-Partition-Models-for-
Multivariate-Change-Point-Analysis/10.1214/22-BA1344.
full.

Quintana:2009:FUC

[QSF09] Fernando A. Quintana, Mark F. J. Steel, and José T. A. S. Ferreira. Flexible univariate continuous distributions. Bayesian Analysis, 4(3):497–521, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2009/vol04/issue03/quintana.pdf;
 http://projecteuclid.org/euclid.ba/1340369852.

Rahman:2016:BQR

[Rah16] Mohammad Arshad Rahman. Bayesian quantile regression for ordinal models. Bayesian Analysis, 11(1):1–24, March 2016. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1423083637.

Rajkowski:2019:AMP

[Raj19] Lukasz Rajkowski. Analysis of the maximal a posteriori partition in the Gaussian Dirichlet process mixture model. Bayesian Analysis, 14(2):477–494, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1532937626.

Ranjan:2010:CAV

[Ran10] Pritam Ranjan. Comment on article by Vernon et al. Bayesian Analysis, 5(4):677–681, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue04/ranjan.
pdf;   http://projecteuclid.org/euclid.ba/1340110848.

See [VGB10a].

Rodrigues:2007:BIE

[RB07] Josemar Rodrigues and Heleno Bolfarine. Bayesian inference for an extended simple regression measurement error model using skewed priors. Bayesian Analysis, 2(2):349–364, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2007/
vol02/issue02/rodrigues.pdf;   http://projecteuclid.
org/euclid.ba/1340393239.

Ruppert:2007:CAD

[RC07] David Ruppert and Raymond J. Carroll. Comment on article by Dominici et al. Bayesian Analysis, 2(1):37–42, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue01/ruppert.pdf;   http://projecteuclid.org/euclid.

ba/1340390060. See [DZP^{+}07a].

Roy:2017:STP

[RC17] Vivekananda Roy and Sounak Chakraborty. Selection of tuning parameters, solution paths and standard errors for Bayesian lassos. Bayesian Analysis, 12(3):753–778, September 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1473276258.

Ramirez-Cobo:2017:BAS

[RCLW17] P. Ramírez-Cobo, R. E. Lillo, and M. P. Wiper. Bayesian analysis of the stationary MAP_{2}. Bayesian Analysis, 12(4):1163–1194, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1477321094.

Raynal:2022:SAB

[RCMO22] Louis Raynal, Sixing Chen, Antonietta Mira, and Jukka-Pekka Onnela. Scalable approximate Bayesian computation for growing network models via extrapolated and sampled summaries. Bayesian Analysis, 17(1):165–192, ???? 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-1/Scalable-Approximate-Bayesian-Computation-
for-Growing-Network-Models-via-Extrapolated/10.1214/
20-BA1248.full.

Rodriguez:2011:NBM

[RD11] Abel Rodríguez and David B. Dunson. Nonparametric Bayesian models through probit stick-breaking processes. Bayesian Analysis, 6(1):145–177, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue01/rodriguez.pdf;   http:
//projecteuclid.org/euclid.ba/1339611944.

Rigat:2006:BMA

[RdGvP06] Fabio Rigat, Mathisca de Gunst, and Jaap van Pelt. Bayesian modelling and analysis of spatio-temporal neuronal networks. Bayesian Analysis, 1(4):733–764, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2006/vol01/issue04/rigat.pdf;
 http://projecteuclid.org/euclid.ba/1340370941.

Ryan:2016:OBE

[RDP16] Caitríona M. Ryan, Christopher C. Drovandi, and Anthony N. Pettitt. Optimal Bayesian experimental design for models with intractable likelihoods using indirect inference applied to biological process models. Bayesian Analysis, 11(3):857–883, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1444396541.

Reich:2016:CAP

[RF16] Brian J. Reich and Montserrat Fuentes. Comment on article by Page and Quintana. Bayesian Analysis, 11(1):303– 305, March 2016. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-11/issue-1/Comment-
on-Article-by-Page-and-Quintana/10.1214/15-BA971B.full.

See [PQ16b].

Roy:2020:HDS

[RGC20] Arkaprava Roy, Subhashis Ghosal, and Kingshuk Roy Choudhury. High dimensional single-index Bayesian modeling of brain atrophy. Bayesian Analysis, 15(4):1229–1249, December 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-15/issue-4/High-Dimensional-
Single-Index-Bayesian-Modeling-of-Brain-Atrophy/10.1214/
19-BA1186.full.

Roos:2011:SAB

[RH11] Małgorzata Roos and Leonhard Held. Sensitivity analysis in Bayesian generalized linear mixed models for binary data. Bayesian Analysis, 6(2):259–278, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2011/vol06/issue02/roos.pdf;
 http://projecteuclid.org/euclid.ba/1339612046.

Rigat:2010:CAM

[Rig10] Fabio Rigat. Comment on article by Manolopoulou et al. Bayesian Analysis, 5(3):451–455, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2010/vol05/issue03/rigat.pdf;
 http://projecteuclid.org/euclid.ba/1340369757.  See

[MCW10b].

Rockova:2014:IGI

[RL14] Veronika Rockova and Emmanuel Lesaffre. Incorporating grouping information in Bayesian variable selection with applications in genomics. Bayesian Analysis, 9(1):221–258, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue01/rockova.pdf;   http://projecteuclid.org/euclid.
ba/1393251777.

Robert:2008:SDP

[RM08] Christian P. Robert and Jean-Michel Marin. On some difficulties with a posterior probability approximation technique. Bayesian Analysis, 3(2):427–441, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue02/robert.pdf;
http://projecteuclid.org/euclid.ba/1340370554.

Rockova:2021:DVS

[RM21] Veronika Rockova and Kenichiro McAlinn. Dynamic variable selection with spike-and-slab process priors. Bayesian Analysis, 16(1):233–269, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/Dynamic-Variable-Selection-with-Spike-and-
Slab-Process-Priors/10.1214/20-BA1199.full.

Roos:2015:SAB

[RMHR15] Małgorzata Roos, Thiago G. Martins, Leonhard Held, and Håvard Rue. Sensitivity analysis for Bayesian hierarchical models. Bayesian Analysis, 10(2):321–349, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884977.

Rufo:2012:LLP

[RMP12] M. J. Rufo, J. Martín, and C. J. Pérez. Log-linear pool to combine prior distributions: A suggestion for a calibrationbased approach. Bayesian Analysis, 7(2):411–438, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue02/rufo.pdf;   http://projecteuclid.org/euclid.
ba/1339878894.

Robert:2007:CAJ

[Rob07] C. P. Robert. Comment on article by Jain and Neal. Bayesian Analysis, 2(3):479–482, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2007/vol02/issue03/robert.
pdf;   http://projecteuclid.org/euclid.ba/1340370722.

See [JN07b].

Robert:2010:SCC

[Rob10] Christian P. Robert. The search for certainty: a critical assessment. Bayesian Analysis, 5(2):213–222, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue02/robert.
pdf;   http://projecteuclid.org/euclid.ba/1340218334.

See comments [Was10, Gel10, Bur10].

Rossell:2022:CPM

[Ros22] David Rossell. Concentration of posterior model probabilities and normalized. Bayesian Analysis, 17(2):565–591, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-2/Concentration-of-
Posterior-Model-Probabilities-and-Normalized-L0-Criteria/
10.1214/21-BA1262.full.

Rougier:2008:CAS

[Rou08] Jonathan Rougier. Comment on article by Sansó et al. [MR2383247]. Bayesian Analysis, 3(1):45–56, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue01/rougier.pdf;   http://projecteuclid.org/euclid.
ba/1340370559. See [SFZ08a].

Rousseau:2015:CAB

[Rou15] Judith Rousseau. Comment on article by Berger, Bernardo, and Sun. Bayesian Analysis, 10(1):233–236, March 2015. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). See [BBS15a].

Rivoirard:2012:PCR

[RR12] Vincent Rivoirard and Judith Rousseau. Posterior concentration rates for infinite dimensional exponential families. Bayesian Analysis, 7(2):311–334, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2012/vol07/issue02/rousseau.pdf;
 http://projecteuclid.org/euclid.ba/1339878890.

Rabinovich:2020:FSM

[RRJW20] Maxim Rabinovich, Aaditya Ramdas, Michael I. Jordan, and Martin J. Wainwright. Function-specific mixing times and concentration away from equilibrium. Bayesian Analysis, 15(2):505– 532, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560391236.

Rubio:2013:BIU

[RS13] Francisco J. Rubio and Mark F. J. Steel. Bayesian inference for P (X < Y) using asymmetric dependent distributions. Bayesian Analysis, 8(1):43–62, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue01/rubio.pdf;   http:
//projecteuclid.org/euclid.ba/1362406651.

Rubio:2014:ITP

[RS14a] Francisco J. Rubio and Mark F. J. Steel. Inference in twopiece location-scale models with Jeffreys priors. Bayesian Analysis, 9(1):1–22, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue01/rubio.pdf;   http:
//projecteuclid.org/euclid.ba/1393251764. See comments

[Ber14, Sco14, WS14, Xu14] and rejoinder [RS14b].

Rubio:2014:R

[RS14b] Francisco J. Rubio and Mark F. J. Steel. Rejoinder. Bayesian Analysis, 9(1):45–52, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2014/vol09/issue01/rubio_rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1393251769. See [RS14a].

Ramamoorthi:2015:PCM

[RSM15] R. V. Ramamoorthi, Karthik Sriram, and Ryan Martin. On posterior concentration in misspecified models. Bayesian Analysis, 10(4):759–789, December 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1423083642.

Ruggeri:2021:NCM

[RSSSSL21] Fabrizio Ruggeri, Marta Sánchez-Sánchez, Miguel Ángel Sordo, and Alfonso Suárez-Llorens. On a new class of multivariate prior distributions: Theory and application in reliability. Bayesian Analysis, 16(1):31–60, March 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-1/On-a-New-Class-of-Multivariate-Prior-Distributions-
-Theory/10.1214/19-BA1191.full.

Ruggeri:2017:HBS

[RSST17] Fabrizio Ruggeri, Zaid Sawlan, Marco Scavino, and Raul Tempone. A hierarchical Bayesian setting for an inverse problem in linear parabolic PDEs with noisy boundary conditions. Bayesian Analysis, 12(2):407–433, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1463078272.

Ruli:2014:MPS

[RSV14] Erlis Ruli, Nicola Sartori, and Laura Ventura. Marginal posterior simulation via higher-order tail area approximations. Bayesian Analysis, 9(1):129–146, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue01/ruli.pdf;   http:/
/projecteuclid.org/euclid.ba/1393251773.

Rodriguez:2008:BDD

[RtH08] Abel Rodriguez and Enrique ter Horst. Bayesian dynamic density estimation. Bayesian Analysis, 3(2):339–365, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue02/rodriguez.pdf;   http://projecteuclid.
org/euclid.ba/1340370551.

Riihimaki:2014:LAL

[RV14] Jaakko Riihimäki and Aki Vehtari. Laplace approximation for logistic Gaussian process density estimation and regression. Bayesian Analysis, 9(2):425–448, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2014/vol09/issue02/jaakko.pdf;
http://projecteuclid.org/euclid.ba/1401148315.

Ranciati:2024:MPR

[RVWG24] Saverio Ranciati, Veronica Vinciotti, Ernst C. Wit, and Giuliano Galimberti. Mixtures of probit regression models with overlapping clusters. Bayesian Analysis, 19(3):843–867, September 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-3/Mixtures-of-Probit-
Regression-Models-with-Overlapping-Clusters/10.1214/
23-BA1372.full.

Rapley:2008:MBI

[RW08] V. E. Rapley and A. H. Welsh. Model-based inferences from adaptive cluster sampling. Bayesian Analysis, 3(4):717–736, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue04/rapley.pdf;   http://projecteuclid.org/
euclid.ba/1340370406.

Ryden:2008:EVM

[Ryd08a] Tobias Rydén. EM versus Markov chain Monte Carlo for estimation of hidden Markov models: a computational perspective. Bayesian Analysis, 3(4):659–688, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue04/ryden.pdf;   http://projecteuclid.org/euclid.

ba/1340370402. See comments [FS08, SK08] and rejoinder [Ryd08b].

Ryden:2008:R

[Ryd08b] Tobias Rydén. Rejoinder. Bayesian Analysis, 3(4):707– 715, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2008/vol03/issue04/ryden_rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1340370405. See [Ryd08a].

Salomond:2018:TSH

[Sal18] Jean-Bernard Salomond. Testing un-separated hypotheses by estimating a distance. Bayesian Analysis, 13(2):461–484, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1498204951.

Sancetta:2012:R

[San12a] Alessio Sancetta. Rejoinder. Bayesian Analysis, 7(1):47–50, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2012/
vol07/issue01/sancetta_rejoinder.pdf; http://projecteuclid.
org/euclid.ba/1339616724. See [San12b].

Sancetta:2012:UBP

[San12b] Alessio Sancetta. Universality of Bayesian predictions. Bayesian Analysis, 7(1):1–36, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue01/sancetta.pdf;   http:/
/projecteuclid.org/euclid.ba/1339616721. See comments

[Cla12, Lia12] and rejoinder [San12a].

Stramer:2011:BII

[SB11] Osnat Stramer and Matthew Bognar. Bayesian inference for irreducible diffusion processes using the pseudo-marginal approach. Bayesian Analysis, 6(2):231–258, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue02/stramer.
pdf;   http://projecteuclid.org/euclid.ba/1339612045.

Short:2006:MSC

[SC06] Margaret B. Short and Bradley P. Carlin. Multivariate spatiotemporal CDFs with random effects and measurement error. Bayesian Analysis, 1(3):595–624, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue03/shortcarlin.pdf;   http://projecteuclid.org/
euclid.ba/1340371054.

Choi:2016:EIN

[sC16] Hwan sik Choi. Expert information and nonparametric Bayesian inference of rare events. Bayesian Analysis, 11(2):421–445, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1432644837.

Sewell:2017:LSA

[SC17] Daniel K. Sewell and Yuguo Chen. Latent space approaches to community detection in dynamic networks. Bayesian Analysis, 12(2):351–377, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1461603847.

Scarpa:2012:CAK

[Sca12] Bruno Scarpa. Comment on article by Kim et al. Bayesian Analysis, 7(4):801–804, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue04/scarpa.pdf;   http://
projecteuclid.org/euclid.ba/1354024460. See [KSLP12a].

Song:2014:BAF

[SCFJ14] Xin-Yuan Song, Jing-Heng Cai, Xiang-Nan Feng, and Xue-Jun Jiang. Bayesian analysis of the functional-coefficient autoregressive heteroscedastic model. Bayesian Analysis, 9(2):371– 396, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2014/vol09/issue02/song.pdf;   http://projecteuclid.
org/euclid.ba/1401148313.

Schmidt:2009:CAC

[Sch09] Alexandra M. Schmidt. Comment on article by Craigmile et al. Bayesian Analysis, 4(1):45–53, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2009/vol04/issue01/schmidt.
pdf;   http://projecteuclid.org/euclid.ba/1340370388.

See [CCL^{+}09a].

Schmidl:2013:R

[SCHT13a] Daniel Schmidl, Claudia Czado, Sabine Hug, and Fabian J. Theis. Rejoinder. Bayesian Analysis, 8(1):33–42, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2013/vol08/
issue01/schmidl-rejoinder.pdf; http://projecteuclid.
org/euclid.ba/1362406650. See [SCHT13b].

Schmidl:2013:VCB

[SCHT13b] Daniel Schmidl, Claudia Czado, Sabine Hug, and Fabian J. Theis. A vine-copula based adaptive MCMC sampler for efficient inference of dynamical systems. Bayesian Analysis, 8(1):1–22, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue01/schmidl.pdf;   http://

projecteuclid.org/euclid.ba/1362406647. See comments

[Woo13, GM13b] and rejoinder [SCHT13a].

Simola:2021:AAB

[SCKGC21] Umberto Simola, Jessi Cisewski-Kehe, Michael U. Gutmann, and Jukka Corander. Adaptive approximate Bayesian computation tolerance selection. Bayesian Analysis, 16(2):397–423, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-16/issue-2/Adaptive-Approximate-
Bayesian-Computation-Tolerance-Selection/10.1214/20-
BA1211.full.

Shi:2022:BCR

[SCKL22] Wei Shi, Ming-Hui Chen, Lynn Kuo, and Paul O. Lewis. Bayesian concentration ratio and dissonance. Bayesian Analysis, 17(3):817–847, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-3/Bayesian-Concentration-Ratio-and-Dissonance/
10.1214/21-BA1277.full.

Scott:2011:BEI

[Sco11] James G. Scott. Bayesian estimation of intensity surfaces on the sphere via needlet shrinkage and selection. Bayesian Analysis, 6(2):307–327, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue02/scott.pdf;   http:
//projecteuclid.org/euclid.ba/1339612048.

Scott:2014:CAR

[Sco14] James G. Scott. Comment on article by Rubio and Steel. Bayesian Analysis, 9(1):25–28, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2014/vol09/issue01/scott.pdf;
 http://projecteuclid.org/euclid.ba/1393251766.  See

[RS14a].

Scricciolo:2014:ABD

[Scr14] Catia Scricciolo. Adaptive Bayesian density estimation in L^{p}- metrics with Pitman–Yor or normalized inverse-Gaussian process kernel mixtures. Bayesian Analysis, 9(2):475–520, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2014/
vol09/issue02/scricciolo.pdf;   http://projecteuclid.
org/euclid.ba/1401148317.

Scutari:2013:PPD

[Scu13a] Marco Scutari. On the prior and posterior distributions used in graphical modelling. Bayesian Analysis, 8(3):505– 532, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/scutari.pdf;   http://
projecteuclid.org/euclid.ba/1378729914. See comments

[Dob13, PS13, Wan13] and rejoinder [Scu13b].

Scutari:2013:R

[Scu13b] Marco Scutari. Rejoinder. Bayesian Analysis, 8(3):549– 552, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue03/scutari-rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1378729918. See [RS14a].

Senn:2008:CAG

[Sen08] Stephen Senn. Comment on article by Gelman. Bayesian Analysis, 3(3):459–461, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue03/senn.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370432. See [Gel08a].

Smith:2014:EBP

[SF14] Isabelle Smith and André Ferrari. Equivalence between the posterior distribution of the likelihood ratio and a p-value in an invariant frame. Bayesian Analysis, 9(4):939–962, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1416579186.

Sanso:2008:ICS

[SFZ08a] Bruno Sansó, Chris E. Forest, and Daniel Zantedeschi. Inferring climate system properties using a computer model. Bayesian Analysis, 3(1):1–37, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2008/vol03/issue01/sanso.pdf;   http:
//projecteuclid.org/euclid.ba/1340370557. See comments

[HG08, Rou08] and rejoinder [SFZ08b].

Sanso:2008:R

[SFZ08b] Bruno Sansó, Chris E. Forest, and Daniel Zantedeschi. Rejoinder. Bayesian Analysis, 3(1):57–61, March 2008. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue01/sanso_rejoinder.pdf;   http://projecteuclid.
org/euclid.ba/1340370560. See [SFZ08a].

Suarez:2016:BCF

[SG16] Adam Justin Suarez and Subhashis Ghosal. Bayesian clustering of functional data using local features. Bayesian Analysis, 11(1): 71–98, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1423083640.

Suarez:2017:BEP

[SG17] Adam J. Suarez and Subhashis Ghosal. Bayesian estimation of principal components for functional data. Bayesian Analysis, 12(2):311–333, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1461092217.

Shahbaba:2014:CAF

[Sha14a] Babak Shahbaba. Comment on article by Finegold and Drton. Bayesian Analysis, 9(3):557–560, September 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1409921104. See [FD14b].

Shalloway:2014:ECR

[Sha14b] David Shalloway. The evidentiary credible region. Bayesian Analysis, 9(4):909–922, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1416579184.

Shaochuan:2021:BMC

[Sha21] Lu Shaochuan. Bayesian multiple changepoint detection for stochastic models in continuous time. Bayesian Analysis, 16(2):521–544, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/Bayesian-Multiple-Changepoint-Detection-for-
Stochastic-Models-in-Continuous-Time/10.1214/20-BA1218.
full.

Shemyakin:2014:HDN

[She14] Arkady Shemyakin. Hellinger distance and non-informative priors. Bayesian Analysis, 9(4):923–938, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1416579185.

Short:2010:PVC

[SHG^{+}10] Margaret Short, Dave Higdon, Laura Guadagnini, Alberto Guadagnini, and Daniel M. Tartakovsky. Predicting vertical connectivity within an aquifer system. Bayesian Analysis, 5 (3):557–581, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2010/vol05/issue03/short.pdf;   http:
//projecteuclid.org/euclid.ba/1340380541.

Short:2007:EFR

[SHK07] Margaret B. Short, David M. Higdon, and Philipp P. Kronberg. Estimation of Faraday rotation measures of the near galactic sky using Gaussian process models. Bayesian Analysis, 2(4):665–680, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2007/vol02/issue04/short.pdf;   http://projecteuclid.
org/euclid.ba/1340370709.

Sona:2023:QED

[SHMM23] Hunanyan Sona, Rue Håvard, Plummer Martyn, and Roos Małgorzata. Quantification of empirical determinacy: The impact of likelihood weighting on posterior location and spread in Bayesian meta-analysis estimated with JAGS and INLA. Bayesian Analysis, 18(3):723–751, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-3/Quantification-of-Empirical-Determinacy-
-The-Impact-of-Likelihood-Weighting/10.1214/22-BA1325.
full.

Sivaganesan:2015:CAB

[Siv15] Siva Sivaganesan. Comment on article by Berger, Bernardo, and Sun. Bayesian Analysis, 10(1):223–226, March 2015. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). See [BBS15a].

Smyth:2008:CAR

[SK08] Padhraic Smyth and Sergey Kirshner. Comment on article by Rydén. Bayesian Analysis, 3(4):699–705, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue04/smyth.pdf;   http://projecteuclid.org/euclid.
ba/1340370404. See [Ryd08a].

Salimans:2013:FFV

[SK13] Tim Salimans and David A. Knowles. Fixed-form variational posterior approximation through stochastic linear regression. Bayesian Analysis, 8(4):837–882, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2013/vol08/issue04/salimans.
pdf;   http://projecteuclid.org/euclid.ba/1386166315.

Serra:2017:AEB

[SK17] Paulo Serra and Tatyana Krivobokova. Adaptive empirical Bayesian smoothing splines. Bayesian Analysis, 12(1):219–238, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1457383100.

Sparks:2015:NSC

[SKG15] Douglas K. Sparks, Kshitij Khare, and Malay Ghosh. Necessary and sufficient conditions for high-dimensional posterior consistency under g-priors. Bayesian Analysis, 10(3):627–664, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1422884985.

Skilling:2006:NSG

[Ski06] John Skilling. Nested sampling for general Bayesian computation. Bayesian Analysis, 1(4):833–859, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/skilling.pdf;   http://projecteuclid.org/euclid.
ba/1340370944.

Staerk:2024:MAS

[SKN24] Christian Staerk, Maria Kateri, and Ioannis Ntzoufras. A Metropolized adaptive subspace algorithm for high-dimensional Bayesian variable selection. Bayesian Analysis, 19(1):261–291, March 2024. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-19/issue-1/A-Metropolized-Adaptive-
Subspace-Algorithm-for-High-Dimensional-Bayesian-Variable/
10.1214/22-BA1351.full.

Santos:2013:PIS

[SLAV13] Cristiano C. Santos, Rosangela H. Loschi, and Reinaldo B. Arellano-Valle. Parameter interpretation in skewed logistic regression with random intercept. Bayesian Analysis, 8(2):381– 410, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2013/
vol08/issue02/santos.pdf;   http://projecteuclid.org/
euclid.ba/1369407557.

Siden:2021:SMP

[SLB^{+}21] Per Sidén, Finn Lindgren, David Bolin, Anders Eklund, and Mattias Villani. Spatial 3D Matérn priors for fast wholebrain fMRI analysis. Bayesian Analysis, 16(4):1251–1278, December 2021. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-16/issue-4/Spatial-
3D-Mat%c3%a9rn-Priors-for-Fast-Whole-Brain-fMRI-Analysis/
10.1214/21-BA1283.full.

Shahn:2017:LCM

[SM17] Zach Shahn and David Madigan. Latent class mixture models of treatment effect heterogeneity. Bayesian Analysis, 12 (3):831–854, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1473362569.

Soriano:2019:MMR

[SM19] Jacopo Soriano and Li Ma. Mixture modeling on related samples by ψ-stick breaking and kernel perturbation. Bayesian Analysis, 14(1):161–180, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
14/issue-1/Mixture-Modeling-on-Related-Samples-by-%cf%
88-Stick-Breaking-and/10.1214/18-BA1106.full.

Shi:2019:LIO

[SMBL19] Yushu Shi, Michael Martens, Anjishnu Banerjee, and Purushottam Laud. Low information omnibus (LIO) priors for Dirichlet process mixture models. Bayesian Analysis, 14(3):677–702, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560240023.

Sandholtz:2023:IBO

[SMBS23] Nathan Sandholtz, Yohsuke Miyamoto, Luke Bornn, and Maurice A. Smith. Inverse Bayesian optimization: Learning human acquisition functions in an exploration vs exploitation search task. Bayesian Analysis, 18(1):1–24, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-1/Inverse-Bayesian-
Optimization--Learning-Human-Acquisition-Functions-in-
an/10.1214/21-BA1303.full.

Seth:2019:MCL

[SMW19] Sohan Seth, Iain Murray, and Christopher K. I. Williams. Model criticism in latent space. Bayesian Analysis, 14(3):703–725, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560240024.

Shahbaba:2007:ICW

[SN07] Babak Shahbaba and Radford M. Neal. Improving classification when a class hierarchy is available using a hierarchy-based prior. Bayesian Analysis, 2(1):221–237, March 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue01/shahbaba.
pdf;   http://projecteuclid.org/euclid.ba/1340390069.

Shestopaloff:2018:SLS

[SN18] Alexander Y. Shestopaloff and Radford M. Neal. Sampling latent states for high-dimensional non-linear state space models with the embedded HMM method. Bayesian Analysis, 13(3):797–822, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1508551720.

Stephens:2023:CIU

[SNMS23] David A. Stephens, Widemberg S. Nobre, Erica E. M. Moodie, and Alexandra M. Schmidt. Causal inference under misspecification: Adjustment based on the propensity score (with discussion). Bayesian Analysis, 18(2):639–694, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-18/issue-2/Causal-Inference-
Under-Mis-Specification--Adjustment-Based-on-the/10.
1214/22-BA1322.full.

Solonen:2012:EMC

[SOL^{+}12] Antti Solonen, Pirkka Ollinaho, Marko Laine, Heikki Haario, Johanna Tamminen, and Heikki Järvinen. Efficient MCMC for climate model parameter estimation: Parallel adaptive chains and early rejection. Bayesian Analysis, 7(3):715– 736, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2012/vol07/issue03/solonen.pdf;   http://
projecteuclid.org/euclid.ba/1346158781.

South:2023:RZV

[SOMD23] L. F. South, C. J. Oates, A. Mira, and C. Drovandi. Regularized zero-variance control variates. Bayesian Analysis, 18(3):865–888, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Regularized-Zero-Variance-Control-Variates/
10.1214/22-BA1328.full.

South:2019:SMC

[SPD19] L. F. South, A. N. Pettitt, and C. C. Drovandi. Sequential Monte Carlo samplers with independent Markov chain Monte Carlo proposals. Bayesian Analysis, 14(3):753–776, September 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1560240027.

Si:2015:BNW

[SPG15] Yajuan Si, Natesh S. Pillai, and Andrew Gelman. Bayesian nonparametric weighted sampling inference. Bayesian Analysis, 10 (3):605–625, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1422884984.

Spitzner:2008:AVH

[Spi08] Dan J. Spitzner. An asymptotic viewpoint on high-dimensional Bayesian testing. Bayesian Analysis, 3(1):121–160, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2008/vol03/
issue01/spitzner.pdf;   http://projecteuclid.org/euclid.
ba/1340370564.

Spitzner:2011:NDC

[Spi11] Dan J. Spitzner. Neutral-data comparisons for Bayesian testing. Bayesian Analysis, 6(4):603–638, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2011/vol06/issue04/spitzner.
pdf;   http://projecteuclid.org/euclid.ba/1339616538.

Schifeling:2016:IMP

[SR16] Tracy A. Schifeling and Jerome P. Reiter. Incorporating marginal prior information in latent class models. Bayesian Analysis, 11(2):499–518, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1434649584.

Sriram:2017:CPC

[SR17] Karthik Sriram and R. V. Ramamoorthi. Correction to: “Posterior Consistency of Bayesian Quantile Regression Based on the Misspecified Asymmetric Laplace Density”. Bayesian Analysis, 12(4):1217–1219, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1505354708. See [SRG13].

Saarela:2023:BNP

[SRA23] Olli Saarela, Christian Rohrbeck, and Elja Arjas. Bayesian nonparametric ordinal regression under a monotonicity constraint. Bayesian Analysis, 18(1):193–221, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-1/Bayesian-Non-Parametric-Ordinal-Regression-
Under-a-Monotonicity-Constraint/10.1214/22-BA1310.full.

Sriram:2013:PCB

[SRG13] Karthik Sriram, R. V. Ramamoorthi, and Pulak Ghosh. Posterior consistency of Bayesian quantile regression based on the misspecified asymmetric Laplace density. Bayesian Analysis, 8(2):479–504, June 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2013/vol08/issue02/sriram.pdf;   http://
projecteuclid.org/euclid.ba/1369407561. See [SR17].

Sarkka:2008:AGT

[SS08] Simo Särkkä and Tommi Sottinen. Application of Girsanov theorem to particle filtering of discretely observed continuous-time non-linear systems. Bayesian Analysis, 3(3):555–584, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue03/sarkka.pdf;   http://projecteuclid.org/
euclid.ba/1340370438.

Skilling:2010:CAH

[SS10] John Skilling and Devinder Sivia. Comment on article by Hogg et al. Bayesian Analysis, 5(1):39–40, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2010/vol05/issue01/skilling.
pdf;   http://projecteuclid.org/euclid.ba/1340369790.

See [HKLM10a].

Shotwell:2011:BOD

[SS11] Matthew S. Shotwell and Elizabeth H. Slate. Bayesian outlier detection with Dirichlet process mixtures. Bayesian Analysis, 6 (4):665–690, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue04/shotwell.pdf;   http:/
/projecteuclid.org/euclid.ba/1339616540.

Sachs:2023:PCG

[SSLD23] Matthias Sachs, Deborshee Sen, Jianfeng Lu, and David Dunson. Posterior computation with the Gibbs zig-zag sampler. Bayesian Analysis, 18(3):909–927, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Posterior-Computation-with-the-Gibbs-Zig-
Zag-Sampler/10.1214/22-BA1319.full.

Shuler:2020:BSM

[SSML20] Kurtis Shuler, Marilou Sison-Mangus, and Juhee Lee. Bayesian sparse multivariate regression with asymmetric nonlocal priors for microbiome data analysis. Bayesian Analysis, 15(2):559– 578, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560909810.

Stanford:2012:CAK

[Sta12] Joseph B. Stanford. Comment on article by Kim et al. Bayesian Analysis, 7(4):805–808, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2012/vol07/issue04/stanford.
pdf;   http://projecteuclid.org/euclid.ba/1354024461.

See [KSLP12a].

Stern:2009:CAM

[Ste09] Hal Stern. Comment on article by Monni and Tadesse. Bayesian Analysis, 4(3):453–456, September 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2009/vol04/issue03/stern.pdf;
 http://projecteuclid.org/euclid.ba/1340369849.  See

[MT09b].

Steorts:2015:ERE

[Ste15] Rebecca C. Steorts. Entity resolution with empirically motivated priors. Bayesian Analysis, 10(4):849–875, December 2015. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1441790411.

Savitsky:2022:BDF

[SW22] Terrance D. Savitsky and Matthew R. Williams. Bayesian dependent functional mixture estimation for area and timeindexed data: an application for the prediction of monthly county employment. Bayesian Analysis, 17(3):791–815, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-3/Bayesian-Dependent-
Functional-Mixture-Estimation-for-Area-and-Time-Indexed/
10.1214/21-BA1274.full.

Sebastiani:2006:BAC

[SXR06] Paola Sebastiani, Hui Xie, and Marco F. Ramoni. Bayesian analysis of comparative microarray experiments by model averaging. Bayesian Analysis, 1(4):707–732, December 2006.

CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2006/vol01/
issue04/sebastiani.pdf;   http://projecteuclid.org/euclid.
ba/1340370940.

Shi:2017:BTS

[SY17] Haolun Shi and Guosheng Yin. Bayesian two-stage design for Phase II clinical trials with switching hypothesis tests. Bayesian Analysis, 12(1):31–51, March 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1450456405.

Shi:2019:CTE

[SY19] Haolun Shi and Guosheng Yin. Control of Type I error rates in Bayesian sequential designs. Bayesian Analysis, 14(2):399– 425, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1529719226.

Shahbaba:2011:CAP

[SYvD11] Babak Shahbaba, Yaming Yu, and David A. van Dyk. Comment on article by Polson and Scott. Bayesian Analysis, 6 (1):31–35, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue01/shahbaba.pdf;   http:/
/projecteuclid.org/euclid.ba/1339611938. See [PS11a].

Touloupou:2018:EMC

[TAN^{+}18] Panayiota Touloupou, Naif Alzahrani, Peter Neal, Simon E. F. Spencer, and Trevelyan J. McKinley. Efficient model comparison techniques for models requiring large scale data augmentation. Bayesian Analysis, 13(2):437–459, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1493431262.

Thomas:2022:LFI

[TDC^{+}22] Owen Thomas, Ritabrata Dutta, Jukka Corander, Samuel Kaski, and Michael U. Gutmann. Likelihood-free inference by ratio estimation. Bayesian Analysis, 17(1):1–31, ???? 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-1/Likelihood-Free-Inference-by-Ratio-
Estimation/10.1214/20-BA1238.full.

Turek:2017:APB

[TdVPAB17] Daniel Turek, Perry de Valpine, Christopher J. Paciorek, and Clifford Anderson-Bergman. Automated parameter blocking for efficient Markov chain Monte Carlo sampling. Bayesian Analysis, 12(2):465–490, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1464266500.

Thawornwattana:2018:DSE

[TDY18] Yuttapong Thawornwattana, Daniel Dalquen, and Ziheng Yang. Designing simple and efficient Markov Chain Monte Carlo proposal kernels. Bayesian Analysis, 13(4):1037–1063, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1510282998.

Terres:2018:BSM

[TFHP18] Maria A. Terres, Montserrat Fuentes, Dean Hesterberg, and Matthew Polizzotto. Bayesian spectral modeling for multivariate spatial distributions of elemental concentrations in soil. Bayesian Analysis, 13(1):1–28, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1478919835.

Tokdar:2011:IBA

[TGK^{+}11] Surya T. Tokdar, Iris Grossmann, Joseph B. Kadane, Anne-Sophie Charest, and Mitchell J. Small. Impact of beliefs about Atlantic tropical cyclone detection on conclusions about trends in tropical cyclone numbers. Bayesian Analysis, 6(4):547–572, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue04/tokdar.pdf;   http://projecteuclid.org/
euclid.ba/1339616536.

Tuyl:2009:PPA

[TGM09] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. Posterior predictive arguments in favor of the Bayes–Laplace prior as the consensus prior for binomial and multinomial parameters. Bayesian Analysis, 4(1):151–158, March 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2009/vol04/issue01/tuyl.pdf;
 http://projecteuclid.org/euclid.ba/1340370393.

terHorst:2014:CAW

[tHM14] Enrique ter Horst and German Molina. Comment on article by Windle and Carvalho. Bayesian Analysis, 9(4):809–818, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1416579179. See [WC14b].

Taddy:2009:MSD

[TK09] Matthew A. Taddy and Athanasios Kottas. Markov switching Dirichlet process mixture regression. Bayesian Analysis, 4(4):793–816, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2009/vol04/issue04/taddy.pdf;   http:
//projecteuclid.org/euclid.ba/1340369825.

Taddy:2012:MMM

[TK12a] Matthew A. Taddy and Athanasios Kottas. Mixture modeling for marked Poisson processes. Bayesian Analysis, 7(2):335– 362, June 2012. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2012/vol07/issue02/taddy.pdf;   http://projecteuclid.
org/euclid.ba/1339878891.

Tokdar:2012:SLQ

[TK12b] Surya T. Tokdar and Joseph B. Kadane. Simultaneous linear quantile regression: A semiparametric Bayesian approach. Bayesian Analysis, 7(1):51–72, March 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2012/vol07/issue01/tokdar2.
pdf;   http://projecteuclid.org/euclid.ba/1339616725.

Tak:2017:DDP

[TM17] Hyungsuk Tak and Carl N. Morris. Data-dependent posterior propriety of a Bayesian beta-binomial-logit model. Bayesian Analysis, 12(2):533–555, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1469021382.

Tan:2014:SVF

[TN14] Linda S. L. Tan and David J. Nott. A stochastic variational framework for fitting and diagnosing generalized linear mixed models. Bayesian Analysis, 9(4):963–1004, December 2014. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1416579187.

Tressou:2008:BNH

[Tre08] Jessica Tressou. Bayesian nonparametrics for heavy tailed distribution. Application to food risk assessment. Bayesian Analysis, 3(2):367–391, June 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2008/vol03/issue02/tressou.pdf;   http://
projecteuclid.org/euclid.ba/1340370552.

Taylor-Rodriguez:2017:JSD

[TRKS^{+}17] Daniel Taylor-Rodríguez, Kimberly Kaufeld, Erin M. Schliep, James S. Clark, and Alan E. Gelfand. Joint species distribution modeling: Dimension reduction using Dirichlet processes. Bayesian Analysis, 12(4):939–967, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1478073617.

Taylor-Rodriguez:2017:IBA

[TRWFB17] Daniel Taylor-Rodríguez, Andrew J. Womack, Claudio Fuentes, and Nikolay Bliznyuk. Intrinsic Bayesian analysis for occupancy models. Bayesian Analysis, 12(3):855–877, September 2017. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1473431536.

Trillos:2020:BUV

[TSA20] Nicolas Garcia Trillos and Daniel Sanz-Alonso. The Bayesian update: Variational formulations and gradient flows. Bayesian Analysis, 15(1):29–56, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1545296444.

Tancredi:2020:UFD

[TSL20] Andrea Tancredi, Rebecca Steorts, and Brunero Liseo. A unified framework for de-duplication and population size estimation (with discussion). Bayesian Analysis, 15(2):633–682, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1551949260.

Tokdar:2010:BDR

[TZG10] Surya T. Tokdar, Yu M. Zhu, and Jayanta K. Ghosh. Bayesian density regression with logistic Gaussian process and subspace projection. Bayesian Analysis, 5(2):319–344, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue02/tokdar.
pdf;   http://projecteuclid.org/euclid.ba/1340218341.

Underhill:2016:CDS

[US16] N. T. Underhill and J. Q. Smith. Context-dependent score based Bayesian information criteria. Bayesian Analysis, 11(4): 1005–1033, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1446124066.

vanDyk:2010:CAV

[vD10] David A. van Dyk. Comment on article by Vernon et al. Bayesian Analysis, 5(4):691–695, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2010/vol05/issue04/vandyk.
pdf;   http://projecteuclid.org/euclid.ba/1340110850.

See [VGB10a].

vandenBoom:2023:BLG

[vdBDB23] Willem van den Boom, Maria De Iorio, and Alexandros Beskos. Bayesian learning of graph substructures. Bayesian Analysis, 18(4):1311–1339, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Bayesian-Learning-of-Graph-Substructures/
10.1214/22-BA1338.full.

vanDyk:2006:DHE

[vDCE^{+}06] David A. van Dyk, Alanna Connors, David N. Esch, Peter Freeman, Hosung Kang, Margarita Karovska, Vinay Kashyap, Aneta Siemiginowska, and Andreas Zezas. Deconvolution in high-energy astrophysics: Science, instrumentation, and methods. Bayesian Analysis, 1(2):189–235, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue02/vandyk189-236.pdf;
 http://projecteuclid.org/euclid.ba/1340371059.  See

comment [LG06] and rejoinder [vDK06].

Vrbik:2012:UIL

[VDF^{+}12] Irene Vrbik, Rob Deardon, Zeny Feng, Abbie Gardner, and John Braun. Using individual-level models for infectious disease spread to model spatio-temporal combustion dynamics. Bayesian Analysis, 7(3):615–638, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2012/vol07/issue03/vrbik.pdf;   http:
//projecteuclid.org/euclid.ba/1346158778.

vanDyk:2006:R

[vDK06] David A. van Dyk and Hosung Kang. Rejoinder. Bayesian Analysis, 1(2):241–248, June 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2006/vol01/issue02/rejoinder241-248.
pdf;   http://projecteuclid.org/euclid.ba/1340371061.

See [vDCE^{+}06].

vanderLinde:2006:CAC

[vdL06] Angelika van der Linde. Comment on article by Celeux et al. Bayesian Analysis, 1(4):699–700, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2006/vol01/issue04/avdl.pdf;
 http://projecteuclid.org/euclid.ba/1340370938.  See

[CFRT06a].

vanderLinde:2007:LIP

[vdL07] Angelika van der Linde. Local influence on posterior distributions under multiplicative modes of perturbation. Bayesian Analysis, 2(2):319–332, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue02/avdl.pdf;   http:/
/projecteuclid.org/euclid.ba/1340393237.

vanderLans:2011:BEM

[vdL11a] Ralf van der Lans. Bayesian estimation of the multinomial logit model: a comment on Holmes and Held, “Bayesian auxiliary variable models for binary and multinomial regression”. Bayesian Analysis, 6(2):353–355, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2011/
vol06/issue02/vanderlans.pdf;   http://projecteuclid.
org/euclid.ba/1339612050. See [HH06, HH11].

vanderLinde:2011:RRR

[vdL11b] Angelika van der Linde. Reduced rank regression models with latent variables in Bayesian functional data analysis. Bayesian Analysis, 6(1):77–126, March 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue01/avdl.pdf;   http:/
/projecteuclid.org/euclid.ba/1339611942.

Venturini:2015:GQT

[VDP15] Sergio Venturini, Francesca Dominici, and Giovanni Parmigiani. Generalized quantile treatment effect: A flexible Bayesian approach using quantile ratio smoothing. Bayesian Analysis, 10 (3):523–552, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1422884982.

Vo:2019:BPB

[VDP19] Brenda N. Vo, Christopher C. Drovandi, and Anthony N. Pettitt. Bayesian parametric bootstrap for models with intractable likelihoods. Bayesian Analysis, 14(1):211–234, March 2019. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1524729725.

vanderPas:2017:UQH

[vdPSvdV17] Stéphanie van der Pas, Botond Szabó, and Aad van der Vaart. Uncertainty quantification for the horseshoe (with discussion). Bayesian Analysis, 12(4):1221–1274, December 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1504231319.

vanderPas:2018:BCD

[vdPvdV18] S. L. van der Pas and A. W. van der Vaart. Bayesian community detection. Bayesian Analysis, 13(3):767–796, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1508378465.

VerHoef:2006:CAG

[Ver06] Jay M. Ver Hoef. Comment on article by Gelfand et al. Bayesian Analysis, 1(1):99–101, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue01/verhoef.
pdf;   http://projecteuclid.org/euclid.ba/1340371074.

See [GSW^{+}06a].

vanErven:2021:FEB

[vES21] Tim van Erven and Botond Szabó. Fast exact Bayesian inference for sparse signals in the normal sequence model. Bayesian Analysis, 16(3):933–960, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-3/Fast-Exact-Bayesian-Inference-for-
Sparse-Signals-in-the-Normal/10.1214/20-BA1227.full.

Vernon:2023:BCM

[VG23] Ian Vernon and John Paul Gosling. A Bayesian computer model analysis of robust Bayesian analyses. Bayesian Analysis, 18(4):1367–1399, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/A-Bayesian-Computer-Model-Analysis-of-Robust-
Bayesian-Analyses/10.1214/22-BA1340.full.

Vernon:2010:GFB

[VGB10a] Ian Vernon, Michael Goldstein, and Richard G. Bower. Galaxy formation: a Bayesian uncertainty analysis. Bayesian Analysis, 5(4):619–669, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2010/vol05/issue04/vernon.pdf;
http://projecteuclid.org/euclid.ba/1340110846. See com-

ments [Poo10, Ran10, LH10, vD10] and rejoinder [VGB10b].

Vernon:2010:R

[VGB10b] Ian Vernon, Michael Goldstein, and Richard G. Bower. Rejoinder. Bayesian Analysis, 5(4):697–708, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue04/vernon_rejoinder.pdf;   http://projecteuclid.
org/euclid.ba/1340110851. See [VGB10a].

Vankov:2019:FEC

[VGE19] Emilian R. Vankov, Michele Guindani, and Katherine B. Ensor. Filtering and estimation for a class of stochastic volatility models with intractable likelihoods. Bayesian Analysis, 14(1): 29–52, March 2019. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1522202635.

Vehtari:2021:RNF

[VGS^{+}21] Aki Vehtari, Andrew Gelman, Daniel Simpson, Bob Carpenter, and Paul-Christian Bürkner. Rank-normalization, folding, and localization: an improved. Bayesian Analysis, 16(2):667–718, June 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-2/Rank-Normalization-Folding-and-Localization-
-An-Improved-R%cb%86-for/10.1214/20-BA1221.full.

Villagran:2008:CMP

[VHJS08] Alejandro Villagran, Gabriel Huerta, Charles S. Jackson, and Mrinal K. Sen. Computational methods for parameter estimation in climate models. Bayesian Analysis, 3(4):823–850, December 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2008/
vol03/issue04/villagran.pdf;   http://projecteuclid.
org/euclid.ba/1340370410.

Vanhatalo:2020:AMG

[VHV20] Jarno Vanhatalo, Marcelo Hartmann, and Lari Veneranta. Additive multivariate Gaussian processes for joint species distribution modeling with heterogeneous data. Bayesian Analysis, 15(2): 415–447, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1559548823.

Viele:2007:NEK

[Vie07] Kert Viele. Nonparametric estimation of Kullback–Leibler information illustrated by evaluating goodness of fit. Bayesian Analysis, 2(2):239–280, June 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue02/viele.pdf;   http:
//projecteuclid.org/euclid.ba/1340370800.

Viroli:2011:MBC

[Vir11] Cinzia Viroli. Model based clustering for three-way data structures. Bayesian Analysis, 6(4):573–602, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2011/vol06/issue04/viroli.
pdf;   http://projecteuclid.org/euclid.ba/1339616537.

Villa:2020:LBP

[VL20] Cristiano Villa and Jeong Eun Lee. A loss-based prior for variable selection in linear regression methods. Bayesian Analysis, 15(2): 533–558, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1560477728.

Ventura:2011:RAB

[VR11] Laura Ventura and Walter Racugno. Recent advances on Bayesian inference for P (X < Y). Bayesian Analysis, 6(3): 411–428, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue03/ventura.pdf;   http://
projecteuclid.org/euclid.ba/1339616470.

Villa:2014:OPN

[VW14] Cristiano Villa and Stephen G. Walker. Objective prior for the number of degrees of freedom of a t distribution. Bayesian Analysis, 9(1):197–220, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue01/villa.pdf;   http:
//projecteuclid.org/euclid.ba/1393251776.

Wang:2012:BGL

[Wan12] Hao Wang. Bayesian graphical Lasso models and efficient posterior computation. Bayesian Analysis, 7(4):867–886, December 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2012/vol07/
issue04/wang.pdf;   http://projecteuclid.org/euclid.
ba/1354024465.

Wang:2013:CAS

[Wan13] Hao Wang. Comment on article by Scutari. Bayesian Analysis, 8(3):543–548, September 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue03/wang.pdf;   http:/
/projecteuclid.org/euclid.ba/1378729917. See [RS14a].

Wang:2015:SIS

[Wan15] Hao Wang. Scaling it up: Stochastic search structure learning in graphical models. Bayesian Analysis, 10(2):351–377, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1422884978.

Wang:2017:MGP

[Wan17] Min Wang. Mixtures of g-priors for analysis of variance models with a diverging number of parameters. Bayesian Analysis, 12(2):511–532, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1467722664.

Wasserman:2006:FBO

[Was06] Larry Wasserman. Frequentist Bayes is objective (comment on articles by Berger and by Goldstein). Bayesian Analysis, 1 (3):451–456, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue03/wasserman.pdf;   http:
//projecteuclid.org/euclid.ba/1340371044. See [Ber06a,

Gol06a].

Wasserman:2008:CAG

[Was08] Larry Wasserman. Comment on article by Gelman. Bayesian Analysis, 3(3):463–465, September 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2008/vol03/issue03/wasserman.
pdf;   http://projecteuclid.org/euclid.ba/1340370433.

See [Gel08a].

Wasserman:2010:CAR

[Was10] Larry Wasserman. Comment on article by Robert. Bayesian Analysis, 5(2):223–228, June 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2010/vol05/issue02/wasserman.
pdf;   http://projecteuclid.org/euclid.ba/1340218335.

See [Rob10].

Wang:2018:GMR

[WB18] Chong Wang and David M. Blei. A general method for robust Bayesian modeling. Bayesian Analysis, 13(4):1163–1191, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1514970064.

Windle:2014:R

[WC14a] Jesse Windle and Carlos M. Carvalho. Rejoinder. Bayesian Analysis, 9(4):819–822, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1416579180. See [WC14b].

Windle:2014:TSS

[WC14b] Jesse Windle and Carlos M. Carvalho. A tractable state-space model for symmetric positive-definite matrices. Bayesian Analysis, 9(4):759–792, December 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid. org/euclid.ba/1416579176. See comments [Cas14, For14, tHM14] and rejoinder [WC14a].

Weng:2018:RTB

[WC18] Ruby Chiu-Hsing Weng and D. Stephen Coad. Real-time Bayesian parameter estimation for item response models. Bayesian Analysis, 13(1):115–137, March 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1482138050.

Wang:2018:NMC

[WCKL18] Yu-Bo Wang, Ming-Hui Chen, Lynn Kuo, and Paul O. Lewis. A new Monte Carlo method for estimating marginal likelihoods. Bayesian Analysis, 13(2):311–333, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1488250818.

Wang:2020:RSB

[WCO20] Junyang Wang, Jon Cockayne, and Chris. J. Oates. A role for symmetry in the Bayesian solution of differential equations. Bayesian Analysis, 15(4):1057–1085, December 2020. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-15/issue-4/A-Role-
for-Symmetry-in-the-Bayesian-Solution-of-Differential/
10.1214/19-BA1183.full.

Warr:2022:AIB

[WDML22] Richard L. Warr, David B. Dahl, Jeremy M. Meyer, and Arthur Lui. The attraction Indian buffet distribution. Bayesian Analysis, 17(3):931–967, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-3/The-Attraction-Indian-Buffet-Distribution/
10.1214/21-BA1279.full.

Weinberg:2012:CBF

[Wei12] Martin D. Weinberg. Computing the Bayes factor from a Markov chain Monte Carlo simulation of the posterior distribution. Bayesian Analysis, 7(3):737–770, September 2012. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2012/vol07/issue03/weinberg.
pdf;   http://projecteuclid.org/euclid.ba/1346158782.

Weng:2010:BEE

[Wen10] Ruby C. Weng. A Bayesian Edgeworth expansion by Stein’s identity. Bayesian Analysis, 5(4):741–763, December 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2010/vol05/
issue04/weng.pdf;   http://projecteuclid.org/euclid.
ba/1340110853.

Wyse:2011:ASF

[WFR11a] Jason Wyse, Nial Friel, and Håvard Rue. Approximate simulation-free Bayesian inference for multiple changepoint models with dependence within segments. Bayesian Analysis, 6 (4):501–528, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2011/vol06/issue04/wyse.pdf;   http:/
/projecteuclid.org/euclid.ba/1339616532. See comments

[Fea11, Koo11] and rejoinder [WFR11b].

Wyse:2011:RCA

[WFR11b] Jason Wyse, Nial Friel, and Håvard Rue. Rejoinder: “Comment on Article by Wyse et al.”. Bayesian Analysis, 6(4): 541–546, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2011/vol06/issue04/wyse_rejoinder.pdf;
http://projecteuclid.org/euclid.ba/1339616535. See [WFR11a].

Williamson:2015:PBA

[WG15] Daniel Williamson and Michael Goldstein. Posterior belief assessment: Extracting meaningful subjective judgements from Bayesian analyses with complex statistical models. Bayesian Analysis, 10(4):877–908, December 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1441011242.

Wade:2018:BCA

[WG18] Sara Wade and Zoubin Ghahramani. Bayesian cluster analysis: Point estimation and credible balls (with discussion). Bayesian Analysis, 13(2):559–626, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1508378464.

Whitaker:2017:BID

[WGBS17] Gavin A. Whitaker, Andrew Golightly, Richard J. Boys, and Chris Sherlock. Bayesian inference for diffusion-driven mixedeffects models. Bayesian Analysis, 12(2):435–463, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1464035697.

Williams:2006:CED

[WHG^{+}06] Brian Williams, Dave Higdon, Jim Gattiker, Leslie Moore, Michael McKay, and Sallie Keller-McNulty. Combining experimental data and computer simulations, with an application to flyer plate experiments. Bayesian Analysis, 1(4): 765–792, December 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.cmu.

edu/journal/2006/vol01/issue04/williams.pdf;   http:/
/projecteuclid.org/euclid.ba/1340370942.

Whiteley:2010:CAM

[Whi10] Nick Whiteley. Comment on article by Manolopoulou et al. Bayesian Analysis, 5(3):457–460, September 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2010/vol05/issue03/nick.pdf;
 http://projecteuclid.org/euclid.ba/1340380535.  See

[MCW10b].

Wilson:2018:SIP

[Wil18] Kevin James Wilson. Specification of informative prior distributions for multinomial models using vine copulas. Bayesian Analysis, 13(3):749–766, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1507536526.

Wu:2023:CLR

[WM23] Pei-Shien Wu and Ryan Martin. A comparison of learning rate selection methods in generalized Bayesian inference.

Bayesian Analysis, 18(1):105–132, March 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-1/A-Comparison-of-Learning-Rate-Selection-Methods-
in-Generalized-Bayesian/10.1214/21-BA1302.full.

Wade:2011:ECP

[WMP11] Sara Wade, Silvia Mongelluzzo, and Sonia Petrone. An enriched conjugate prior for Bayesian nonparametric inference. Bayesian Analysis, 6(3):359–385, September 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2011/vol06/issue03/wade.pdf;
 http://projecteuclid.org/euclid.ba/1339616468.

Wu:2021:BMQ

[WN21] Teng Wu and Naveen N. Narisetty. Bayesian multiple quantile regression for linear models using a score likelihood. Bayesian Analysis, 16(3):875–903, September 2021. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
16/issue-3/Bayesian-Multiple-Quantile-Regression-for-
Linear-Models-Using-a-Score/10.1214/20-BA1217.full.

Wang:2022:EEA

[WOJL22] Jin Wang, Yunbo Ouyang, Yuan Ji, and Feng Liang. An ensemble EM algorithm for Bayesian variable selection. Bayesian Analysis, 17(3):879–900, September 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/journals/bayesian-analysis/volume-17/issue-3/An-
Ensemble-EM-Algorithm-for-Bayesian-Variable-Selection/
10.1214/21-BA1275.full.

Woodard:2013:CAS

[Woo13] Dawn B. Woodard. Comment on article by Schmidl et al. Bayesian Analysis, 8(1):23–26, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2013/vol08/issue01/woodard.
pdf;   http://projecteuclid.org/euclid.ba/1362406648.

See [SCHT13b].

Wooff:2014:BLS

[Woo14] David A. Wooff. Bayes linear sufficiency in non-exchangeable multivariate multiple regressions. Bayesian Analysis, 9(1):77– 96, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://ba.stat.cmu.edu/journal/
2014/vol09/issue01/wooff.pdf;   http://projecteuclid.
org/euclid.ba/1393251771.

Wand:2011:MFV

[WOPF11] Matthew P. Wand, John T. Ormerod, Simone A. Padoan, and Rudolf Frührwirth. Mean field variational Bayes for elaborate distributions. Bayesian Analysis, 6(4):847–900, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue04/wand.pdf;   http://projecteuclid.org/euclid.
ba/1339616546.

Wade:2022:CWL

[WPCAV22] Sara Wade, Raffaella Piccarreta, Andrea Cremaschi, and Isadora Antoniano-Villalobos. Colombian women’s life patterns: a multivariate density regression approach. Bayesian Analysis, 17(2):405–433, June 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-2/Colombian-Womens-Life-Patterns-A-Multivariate-
Density-Regression-Approach/10.1214/20-BA1256.full.

Wang:2023:DAB

[WPS23] Yuexi Wang, Nicholas Polson, and Vadim O. Sokolov. Data augmentation for Bayesian deep learning. Bayesian Analysis, 18(4):1041–1069, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Data-Augmentation-for-Bayesian-Deep-Learning/
10.1214/22-BA1331.full.

Wang:2011:DFI

[WRC11] Hao Wang, Craig Reeson, and Carlos M. Carvalho. Dynamic financial index models: Modeling conditional dependencies via graphs. Bayesian Analysis, 6(4):639–664, December 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/
issue04/wang.pdf;   http://projecteuclid.org/euclid.
ba/1339616539.

Weiss:2014:CAR

[WS14] Robert E. Weiss and Marc A. Suchard. Comment on article by Rubio and Steel. Bayesian Analysis, 9(1):29–38, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2014/vol09/
issue01/weiss.pdf;   http://projecteuclid.org/euclid.
ba/1393251767. See [RS14a].

Williams:2020:BEU

[WS20] Matthew R. Williams and Terrance D. Savitsky. Bayesian estimation under informative sampling with unattenuated dependence. Bayesian Analysis, 15(1):57–77, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1546570987.

Warne:2022:VOA

[WSD22] David J. Warne, Scott A. Sisson, and Christopher Drovandi. Vector operations for accelerating expensive Bayesian computations — a tutorial guide. Bayesian Analysis, 17(2):593– 622, June 2022. CODEN ???? ISSN 1931-6690 (print),

1931-6690 (electronic). URL https://projecteuclid.org/
journals/bayesian-analysis/volume-17/issue-2/Vector-
Operations-for-Accelerating-Expensive-Bayesian-Computations-
-A-Tutorial/10.1214/21-BA1265.full.

Wang:2013:STM

[WSDC13] Eric Wang, Esther Salazar, David Dunson, and Lawrence Carin. Spatio-temporal modeling of legislation and votes. Bayesian Analysis, 8(1):233–268, March 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue01/carin.pdf;   http:
//projecteuclid.org/euclid.ba/1362406659.

Wang:2006:CPG

[WT06] Bo Wang and D. M. Titterington. Convergence properties of a general algorithm for calculating variational Bayesian estimates for a normal mixture model. Bayesian Analysis, 1(3):625–650, September 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2006/vol01/issue03/tiwang.pdf;
http://projecteuclid.org/euclid.ba/1340371055.

Wang:2020:BQR

[WT20] Zhi-Qiang Wang and Nian-Sheng Tang. Bayesian quantile regression with mixed discrete and nonignorable missing covariates. Bayesian Analysis, 15(2):579–604, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/euclid.ba/1560909811.

Weaver:2016:CEB

[WWACH16] Brian P. Weaver, Brian J. Williams, Christine M. Anderson-Cook, and David M. Higdon. Computational enhancements to Bayesian design of experiments using Gaussian processes. Bayesian Analysis, 11(1):191–213, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1425492493.

Xie:2022:BSS

[XCPX22] Fangzheng Xie, Joshua Cape, Carey E. Priebe, and Yanxun Xu. Bayesian sparse spiked covariance model with a continuous matrix shrinkage prior. Bayesian Analysis, 17(4):1193–1217, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/journals/
bayesian-analysis/volume-17/issue-4/Bayesian-Sparse-
Spiked-Covariance-Model-with-a-Continuous-Matrix-Shrinkage/
10.1214/21-BA1292.full.

Xiong:2016:RLS

[XJC16] Jie Xiong, Väinö Jääskinen, and Jukka Corander. Recursive learning for sparse Markov models. Bayesian Analysis, 11 (1):247–263, March 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1429105670.

Xi:2016:BQR

[XLH16] Ruibin Xi, Yunxiao Li, and Yiming Hu. Bayesian quantile regression based on the empirical likelihood with spike and slab priors. Bayesian Analysis, 11(3):821–855, September 2016. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

http://projecteuclid.org/euclid.ba/1444396540.

Xu:2013:NBB

[XLY^{+}13] Yanxun Xu, Juhee Lee, Yuan Yuan, Riten Mitra, Shoudan Liang, Peter Müller, and Yuan Ji. Nonparametric Bayesian bi-clustering for next generation sequencing count data. Bayesian Analysis, 8(4):759–780, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue04/xu.pdf;   http://
projecteuclid.org/euclid.ba/1386166312.

Xing:2007:HMD

[XS07] Eric P. Xing and Kyung-Ah Sohn. Hidden Markov Dirichlet process: Modeling genetic inference in open ancestral space. Bayesian Analysis, 2(3):501–527, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:/

/ba.stat.cmu.edu/journal/2007/vol02/issue03/xing.pdf;
 http://projecteuclid.org/euclid.ba/1340370725.

Xu:2017:DTC

[XTMR17] Yanxun Xu, Peter F. Thall, Peter Müller, and Mehran J. Reza. A decision-theoretic comparison of treatments to resolve air leaks after lung surgery based on nonparametric modeling. Bayesian Analysis, 12(3):639–652, September 2017. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1469553354.

Xu:2014:CAR

[Xu14] Xinyi Xu. Comment on article by Rubio and Steel. Bayesian Analysis, 9(1):39–44, March 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2014/vol09/issue01/xu.pdf;   http://
projecteuclid.org/euclid.ba/1393251768. See [RS14a].

Xie:2020:ABN

[XX20] Fangzheng Xie and Yanxun Xu. Adaptive Bayesian nonparametric regression using a kernel mixture of polynomials with application to partial linear models. Bayesian Analysis, 15(1):159– 186, March 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1550826222.

Yau:2011:HBN

[YH11] Christopher Yau and Chris Holmes. Hierarchical Bayesian nonparametric mixture models for clustering with variable relevance determination. Bayesian Analysis, 6(2):329–351, June 2011. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2011/vol06/

issue02/yau.pdf;   http://projecteuclid.org/euclid.ba/
1339612049.

Yang:2016:BLF

[YHW16] Wen-Hsi Yang, Scott H. Holan, and Christopher K. Wikle. Bayesian lattice filters for time-varying autoregression and timefrequency analysis. Bayesian Analysis, 11(4):977–1003, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1445263834.

Yin:2009:BGM

[Yin09a] Guosheng Yin. Bayesian generalized method of moments. Bayesian Analysis, 4(2):191–207, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2009/vol04/issue02/yin.pdf;
http://projecteuclid.org/euclid.ba/1340370272. See com-

ments [CK09, Cra09] and rejoinder [Yin09b].

Yin:2009:R

[Yin09b] Guosheng Yin. Rejoinder. Bayesian Analysis, 4(2):217–222, June 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue02/yin_rejoinder.pdf;   http://projecteuclid.org/
euclid.ba/1340370275. See [Yin09a].

Yu:2013:CBM

[YMP13] Qingzhao Yu, Steven N. MacEachern, and Mario Peruggia. Clustered Bayesian model averaging. Bayesian Analysis, 8 (4):883–908, December 2013. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2013/vol08/issue04/yu.pdf;   http://
projecteuclid.org/euclid.ba/1386166316.

Yuchi:2023:BUQ

[YMX23] Henry Shaowu Yuchi, Simon Mak, and Yao Xie. Bayesian uncertainty quantification for low-rank matrix completion. Bayesian Analysis, 18(2):491–518, June 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-2/Bayesian-Uncertainty-Quantification-for-Low-
Rank-Matrix-Completion/10.1214/22-BA1317.full.

Yang:2020:CGS

[YN20] Xinming Yang and Naveen N. Narisetty. Consistent group selection with Bayesian high dimensional modeling. Bayesian Analysis, 15(3):909–935, September 2020. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
15/issue-3/Consistent-Group-Selection-with-Bayesian-
High-Dimensional-Modeling/10.1214/19-BA1178.full.

Yao:2022:BHS

[YPVG22] Yuling Yao, Gregor Pirs, Aki Vehtari, and Andrew Gelman. Bayesian hierarchical stacking: Some models are (somewhere) useful. Bayesian Analysis, 17(4):1043–1071, December 2022. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-17/issue-4/Bayesian-Hierarchical-Stacking-Some-
Models-Are-Somewhere-Useful/10.1214/21-BA1287.full.

Yan:2007:BDT

[YS07] Guofen Yan and J. Sedransk. Bayesian diagnostic techniques for detecting hierarchical structure. Bayesian Analysis, 2 (4):735–760, December 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.stat.

cmu.edu/journal/2007/vol02/issue04/yan.pdf;   http://
projecteuclid.org/euclid.ba/1340370713.

Yin:2022:FME

[YSB22] Fan Yin, Weining Shen, and Carter T. Butts. Finite mixtures of ERGMs for modeling ensembles of networks. Bayesian Analysis, 17(4):1153–1191, December 2022. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
17/issue-4/Finite-Mixtures-of-ERGMs-for-Modeling-Ensembles-
of-Networks/10.1214/21-BA1298.full.

Yang:2018:SMC

[YSH18] Biao Yang, Jonathan R. Stroud, and Gabriel Huerta. Sequential Monte Carlo smoothing with parameter estimation. Bayesian Analysis, 13(4):1137–1161, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/euclid.ba/1514516432.

Yue:2014:BAS

[YSLR14] Yu Ryan Yue, Daniel Simpson, Finn Lindgren, and Håvard Rue. Bayesian adaptive smoothing splines using stochastic differential equations. Bayesian Analysis, 9(2):397–424, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2014/vol09/issue02/yue.pdf;
 http://projecteuclid.org/euclid.ba/1401148314.

Yao:2018:USA

[YVSG18] Yuling Yao, Aki Vehtari, Daniel Simpson, and Andrew Gelman. Using stacking to average Bayesian predictive distributions (with discussion). Bayesian Analysis, 13(3):917–1007, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1516093227.

Yang:2016:SMC

[YZCC16] Jingjing Yang, Hongxiao Zhu, Taeryon Choi, and Dennis D. Cox. Smoothing and mean-covariance estimation of functional data with a Bayesian hierarchical model. Bayesian Analysis, 11 (3):649–670, September 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1440594947.

Zhang:2018:VSP

[ZB18] Yan Zhang and Howard D. Bondell. Variable selection via penalized credible regions with Dirichlet–Laplace global-local shrinkage priors. Bayesian Analysis, 13(3):823–844, September 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1508551721.

Zhang:2020:BIS

[ZC20] Bohai Zhang and Noel Cressie. Bayesian inference of spatiotemporal changes of Arctic sea ice. Bayesian Analysis, 15(2): 605–631, June 2020. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL https://projecteuclid.org/euclid.
ba/1589421852.

Zhu:2017:BFD

[ZD17] Bin Zhu and David B. Dunson. Bayesian functional data modeling for heterogeneous volatility. Bayesian Analysis, 12(2):335– 350, June 2017. CODEN ???? ISSN 1931-6690 (print), 1931-

6690 (electronic). URL http://projecteuclid.org/euclid.
ba/1461603846.

Zhou:2019:FMF

[ZG19] Quan Zhou and Yongtao Guan. Fast model-fitting of Bayesian variable selection regression using the iterative complex factorization algorithm. Bayesian Analysis, 14(2):573–594, June 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL https://projecteuclid.org/euclid.ba/1535508388.

Zhang:2023:CDU

[ZHA23] Hongmei Zhang, Xianzheng Huang, and Hasan Arshad. Comparing dependent undirected Gaussian networks. Bayesian Analysis, 18(4):1341–1366, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-4/Comparing-Dependent-Undirected-Gaussian-Networks/
10.1214/22-BA1337.full.

Zhang:2016:TCP

[ZHG^{+}16] Hongmei Zhang, Xianzheng Huang, Jianjun Gan, Wilfried Karmaus, and Tara Sabo-Attwood. A two-component G-prior for variable selection. Bayesian Analysis, 11(2):353–380, June 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic).

URL http://projecteuclid.org/euclid.ba/1430830144.

Zhou:2018:NBN

[Zho18] Mingyuan Zhou. Nonparametric Bayesian negative binomial factor analysis. Bayesian Analysis, 13(4):1065–1093, December 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1510801993.

Zidek:2015:CAF

[Zid15] James V. Zidek. Comment on article by Ferreira and Gamerman. Bayesian Analysis, 10(3):749–752, September 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//projecteuclid.org/euclid.ba/1430318842. See [dG15].

Zhang:2010:BIA

[ZJLC10] Xiaoxi Zhang, Timothy D. Johnson, Roderick J. A. Little, and Yue Cao. A Bayesian image analysis of radiation induced changes in tumor vascular permeability. Bayesian Analysis, 5(1):189–212, March 2010. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL http://ba.stat.cmu.edu/journal/2010/
vol05/issue01/zhang2.pdf;   http://projecteuclid.org/
euclid.ba/1340369798.

Zareifard:2018:MSS

[ZKRVA18] Hamid Zareifard, Majid Jafari Khaledi, Firoozeh Rivaz, and Mohammad Q. Vahidi-Asl. Modeling skewed spatial data using a convolution of Gaussian and log-Gaussian processes. Bayesian Analysis, 13(2):531–557, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/euclid.ba/1502762659.

Zheng:2023:NNM

[ZKS23] Xiaotian Zheng, Athanasios Kottas, and Bruno Sansó. Nearestneighbor mixture models for non-Gaussian spatial processes. Bayesian Analysis, 18(4):1191–1222, December 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-18/issue-4/Nearest-Neighbor-Mixture-Models-for-
Non-Gaussian-Spatial-Processes/10.1214/23-BA1405.full.

Zhang:2015:CPP

[ZL15] Zhihua Zhang and Jin Li. Compound Poisson processes, latent shrinkage priors and Bayesian nonconvex penalization. Bayesian Analysis, 10(2):247–274, June 2015. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

projecteuclid.org/euclid.ba/1422884974.

Zhang:2024:BAE

[ZL24] Qian Zhang and Faming Liang. Bayesian analysis of exponential random graph models using stochastic gradient Markov chain Monte Carlo. Bayesian Analysis, 19(2):595–621, June 2024. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-19/issue-2/Bayesian-Analysis-of-Exponential-Random-
Graph-Models-Using-Stochastic-Gradient/10.1214/23-BA1364.
full.

Zeng:2024:BIS

[ZLV24] Zijian Zeng, Meng Li, and Marina Vannucci. Bayesian image-on-scalar regression with a spatial global-local spike-and-slab prior. Bayesian Analysis, 19(1):235–260, March 2024. CODEN ????

ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
19/issue-1/Bayesian-Image-on-Scalar-Regression-with-
a-Spatial-Global-Local/10.1214/22-BA1352.full.

Zemplenyi:2023:BOE

[ZM23] Michele Zemplenyi and Jeffrey W. Miller. Bayesian optimal experimental design for inferring causal structure. Bayesian Analysis, 18(3):929–956, September 2023. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https:/

/projecteuclid.org/journals/bayesian-analysis/volume-
18/issue-3/Bayesian-Optimal-Experimental-Design-for-
Inferring-Causal-Structure/10.1214/22-BA1335.full.

Zanella:2021:MLM

[ZR21] Giacomo Zanella and Gareth Roberts. Multilevel linear models, Gibbs samplers and multigrid decompositions (with discussion). Bayesian Analysis, 16(4):1309–1391, December 2021. CO-DEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL

https://projecteuclid.org/journals/bayesian-analysis/
volume-16/issue-4/Multilevel-Linear-Models-Gibbs-Samplers-
and-Multigrid-Decompositions-with-Discussion/10.1214/
20-BA1242.full.

Zhang:2009:SSC

[ZS09] Hongmei Zhang and Hal Stern. Sample size calculation for finding unseen species. Bayesian Analysis, 4(4):763–792, December 2009. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2009/vol04/
issue04/zhang.pdf;   http://projecteuclid.org/euclid.
ba/1340369824.

Zhang:2007:SAB

[ZSM07] Song Zhang, Ya-Chen Tina Shih, and Peter Müller. A spatiallyadjusted Bayesian additive regression tree model to merge two datasets. Bayesian Analysis, 2(3):611–633, September 2007. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (elec-

tronic). URL http://ba.stat.cmu.edu/journal/2007/vol02/
issue03/zhang.pdf;   http://projecteuclid.org/euclid.
ba/1340370729.

Zhang:2018:VHM

[ZSZ18] Cheng Zhang, Babak Shahbaba, and Hongkai Zhao. Variational Hamiltonian Monte Carlo via score matching. Bayesian Analysis, 13(2):485–506, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://projecteuclid.

org/euclid.ba/1500948232.

Zhao:2016:BAC

[ZWC^{+}16] Tingting Zhao, Ziyu Wang, Alexander Cumberworth, Joerg Gsponer, Nando de Freitas, and Alexandre Bouchard-Côté. Bayesian analysis of continuous time Markov chains with application to phylogenetic modelling. Bayesian Analysis, 11(4): 1203–1237, December 2016. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://projecteuclid.

org/euclid.ba/1448899900.

Zhang:2014:MVD

[ZWDJ14] Zhihua Zhang, Dakan Wang, Guang Dai, and Michael I. Jordan. Matrix-variate Dirichlet process priors with applications. Bayesian Analysis, 9(2):259–286, June 2014. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://

ba.stat.cmu.edu/journal/2014/vol09/issue02/zhang.pdf;
 http://projecteuclid.org/euclid.ba/1401148309.

Zhao:2018:BAR

[ZWF^{+}18] Lili Zhao, Weisheng Wu, Dai Feng, Hui Jiang, and XuanLong Nguyen. Bayesian analysis of RNA-Seq data using a family of negative binomial models. Bayesian Analysis, 13(2):411–436, June 2018. CODEN ???? ISSN 1931-6690 (print), 1931-6690

(electronic). URL https://projecteuclid.org/euclid.ba/
1491616976.